### v2

# Hypergraph Representations of scRNA-seq Data for Improved Clustering with Random Walks

Abstract
Analysis of single-cell RNA sequencing data is often conducted through considering their network projections such as co-expression networks, primarily due to the abundant availability of network analysis tools for downstream tasks. However, this approach has several limitations. First, projecting scRNA-seq data into networks can result in the loss of higher-order information since it summarizes the gene expression profiles across cells into pairwise cell or gene similarities. Secondly, scRNA-seq data is often very sparse, but the constructed co-expression network projections are fully connected without sparsification, resulting in dense and inefficient data representation. Further, the sparsity of scRNA-seq data leads to co-occurrence of zero expression in a large proportion of the genes. Such zero-expression co-occurence will be accounted as similarity in gene expression among cells by correlation methods such as Spearman’s rank and Pearson’s correlation coefficient and lead to an overestimation of co-expression.
To address these limitations, we propose conceptualizing scRNA-seq expression data as hypergraphs, which are generalized graphs in which the hyperedges can connect more than two vertices. This flexibility makes hypergraphs particularly well-suited for representing scRNA-seq data without requiring further transformation such as co-expression network generation. In the context of scRNA-seq data, the hypergraph nodes represent cells and the edges represent genes. Specifically, each hyperedge connects to all the cells where the gene is actively expressed and records the expression of the gene across different cells. The hypergraph conceptualization enables us to explore multi-way relationships beyond the pairwise interactions in co-expression networks for scRNA-seq data analysis while utilizing network analysis tools.
We propose two novel hypergraph random walk-based cell clustering methods: (1) the Dual-Importance Preference Hypergraph Walk (DIPHW) and (2) the co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk (CoMem-DIPHW). These methods outperform widely used and established methods in terms of clustering accuracy on both simulated and real scRNA-seq datasets. The improvement is especially significant when data modularity is weak. By conceptualizing the scRNA-seq data as hypergraphs, our methods provide a more nuanced representation of gene expression data, allowing the incorporation of network perspectives without the information loss caused by enforcing a unipartite co-expression network projection. Furthermore, CoMem-DIPHW incorporates the gene co-expression network, cell co-expression network, and the cell-gene expression hypergraph from the single-cell abundance counts data altogether for embedding computation. This approach accounts for both the local information from single-cell level gene expression and the global information from pairwise similarities in the two co-expression networks.
I Introduction
Network analysis has become a popular tool for studying complex systems in biology, as a reflection of the combinatorial interaction mechanism of biomolecules. Its applications include functional analysis  [1, 2, 3, 4] and interaction prediction in protein-protein interaction (PPI) networks [5, 6, 7], regulator identification and pathway analysis in gene regulatory networks [8, 9, 10], disease modeling, prediction, and intervention in epidemiology [11, 12, 13], and cell type identification in co-expression networks [14, 15, 16, 17].
While these methods employed unipartite networks to aid analysis, it is often overlooked that unipartite networks are often not the most natural or information-preserving representation of the complex system we are interested in studying since many complex systems involve interactions from more than one type of entities. Examples include protein-protein interaction (PPI) networks where the nodes are proteins and edges could be physical contacts among proteins or functional interactions in biological pathways, gene regulatory networks (GRNs) where nodes represent genes and a directed edge exists if one gene regulates the expression of another gene, co-expression networks where the nodes are genes, cells, or tissues and the weighted edges indicate similarity in expression profile between node pairs.
An example of a network that captures higher-order interactions between different entity types is the bipartite network, where nodes are partitioned by two separate groups, and edges only connect nodes from different groups. Bipartite networks are particularly suitable for modeling systems where two types of entities interact, such as authors and papers in a collaboration network [18], recommendation systems [19] where nodes represent users and the recommending items, and specifically in biology, plant-pollinators and predator-prey interaction networks [20] in community ecology studies, microbe-host interaction networks [21] in microbiology, drug–protein interaction networks [22] , and transcription factor-gene networks in genomics.
However, instead of modelling these systems using network models that could capture higher order interactions, most often, a unipartite network projection is chosen such that the system could be analyzed with the abundant network analysis methods developed based on unipartite networks [23, 14]. For example, in the bipartite representation of a drug interaction network, the two classes of nodes represent drugs and protein targets respectively. A drug node and a protein node are connected if the drug targets the protein. In its unipartite transformed network, a drug-drug interaction network could be constructed based on shared targets. Similarly, in the bipartite protein complex network, proteins and biological functions are two types of nodes that are connected if the proteins are involved in the biological function. Yet, in the unipartite representation of the protein complex network, the clique graph of the original bipartite graph is constructed where the only class of nodes are the proteins and the protein nodes are connected if they are in the same complex that performs a biological function. Unipartite representations fail to capture the biological functions associated with groups of proteins. While unipartite projections can reveal relationships between nodes of the same type in a bipartite network, they often obscure the specificity of these interactions.
Single-cell RNA sequencing (scRNA-seq) has enabled the profiling of gene expression at the individual cell level [24, 25, 26, 27], whereas conventional bulk tissue RNA sequencing measures expression at the tissue level, averaging gene expression over an ensemble of cells. Cell type identification, as one of the most important downstream tasks in single-cell RNA-seq (scRNA-seq) data analysis  [28], has applications across biology and medicine, including tracing the trajectories of distinct cell lineages in development for cell differentiation studies  [29], tissue heterogeneity analysis for cancer research  [30, 31, 32], immune cell profiling for therapy development  [33, 34, 35, 36], and biomarker discovery for diagnosis.
[…]
Given the issues with unipartite projections, it is crucial to explore alternative network representations that can accurately reflect the scRNA-seq data. Hypergraphs offer a neat solution by representing multi-way relationships in scRNA-seq directly, without the need for further data projection. Moreover, each bipartite graph can also be represented as a hypergraph, by modelling one set of entities as nodes and the other as hyperedges, offering a generalized representation of high-order interactions. In this work, we propose a hypergraph-based approach to analyze scRNA-seq data to overcome the limitations of unipartite projections.
Figure 1: Network Representations of the scRNA-seq Data. The scRNA-seq cell-gene expression data can be represented by various networks, including unipartite network projections, such as cell and gene co-expression networks. In a cell co-expression network, the nodes are cells and the edge weights represent similarity in cell-pair expression. Similarly, in gene co-expression networks, nodes represent genes. Additionally, hypergraph and bipartite network representations preserve the exact expression level of a gene in a cell. In the hypergraph representation, nodes represent cells and hyperedges represent genes. In the bipartite representation, one set of nodes represents cells, and the other set represents genes.
Figure 2: CoMem Random Walk Illustration. In the hypergraph representation of the scRNA-seq data, nodes represent cells and edges represent genes. The expression level of a gene in a cell is captured by the weight of a node in an edge. Random walks are performed on the hypergraph by alternately considering the probability of the walker choosing a node given an edge, and choosing an edge given the current node. For DIPHW, the input is the hypergraph representation of the scRNA-seq data. For CoMem-DIPHW, the input includes the hypergraph, cell co-expression network, and gene co-expression network, which serve as node and edge similarity matrices to aid memory incorporation during the walk. A small neural network (Word2Vec) is applied to the sampled node paths from the hypergraph random walk process to compute cell embeddings. Finally, cell clustering is achieved by applying K-means to the obtained cell embeddings.
II Methods
In scRNA-seq analysis, clustering cells based on co-expression networks projected from count data is a common approach for cell type identification. However, as discussed in III.3, the correlation inflation problem associated with co-expression networks caused by the induced sparsity in scRNA-seq data, as well as information loss from projecting a hypergraph to a unipartite simple graph, motivated us to look beyond co-expression networks for scRNA-seq analysis.
We conceptualize scRNA-seq data as a hypergraph $\mathcal{H}$. In this framework, we propose a hypergraph random walk-based clustering method that models the relationships between cells and genes in the cell-gene expression abundance count data. Specifically, each cell is represented as a node, and each gene is represented as a hyperedge. Each hyperedge captures the expression profile of a gene across all cells, allowing us to apply hypergraph based analysis tools without enduring the loss of information caused by projecting the scRNA-seq data into a simple unipartite graph, e.g. the co-expression networks. We design random walk algorithms that capture cell-cell interactions by sampling the cell-to-gene and gene-to-cell paths in the conceptualized hypergraph. The random walk sampling probability distribution is designed to account for the specific cell-gene expressions via the edge-dependent vertex selection probability and the vertex-dependent edge selection probability. The sampled paths embed information regarding the amplified expression level of a gene in a cell, both in comparison to the expression levels of other genes in that cell and to the expression levels of the same gene in other cells. As a step further, CoMem-DIPHW (the co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk), the second hypergraph-based random walk algorithm proposed for clustering, incorporates a memory mechanism that takes into account both the conceptualized hypergraph from the scRNA-seq expression data, and the two projected unipartite co-expression networks, i.e. the cell and gene co-expression network, for cell-cell interaction sampling. The sampled cell-cell interaction paths would then be fed to a small language model (Word2Vec) to generate efficient numerical representations (embeddings) of the cells. These embeddings would then be used to cluster the cells by applying a clustering algorithm such as UMAP or t-SNE to group the cells based on their similarity in the embedded space.
We emphasize that this hypergraph conceptualization of the scRNA-seq data does not require further transformation on the data and consequently would not result in any information loss or computation cost from converting the scRNA-seq into co-expression networks. We remark that a matrix is a 2-dimensional array structured for organizing information related to two variables. Thus a matrix could naturally be conceptualized as a hypergraph by considering the column variables as nodes and the row variables as edges and vice versa.
II.1 Hypergraph representation
The single-cell RNA sequencing (scRNA-seq) data can be conceptualized as a hypergraph $\mathcal{H}=(V,E)$, where:
  • $V$ represents the vertex set corresponding to the cells profiled in the experiment.
  • $E$ denotes the hyperedge set representing genes, with each hyperedge $e\in E$ connecting to all cells $\in V$ in which the gene $e$ is actively expressed.
The incidence matrix $\mathbf{I}_{\mathcal{H}}$ of the hypergraph $\mathcal{H}$ is constructed such that $\mathbf{I}_{\mathcal{H}}(v,e)$ is given by the expression level, i.e. abundance counts, of gene $e$ in cell $v$.
II.2 Bipartite representation
We chose the hypergraph conceptualization over bipartite representation for consistency since the DIPHW algorithm we introduced is an extension of the edge-dependent vertex weight hypergraph random walk (EDVW) [41]. DIPHW extended EDVW hypergraph random walk [41] with a vertex-dependent edge selection probability, as well as a preference exponent for clustering acceleration.
A bipartite conceptualization could be adopted for our proposed methods DIPHW, CoMem, and CoMem-DIPHW by considering the incidence matrix $\mathbf{I}_{\mathcal{H}}$ of the hypergraph $\mathcal{H}$ as a bipartite graph. The two sets of nodes in the bipartite representation are the set of cells V and the set of genes E. Similarly, each gene is connected to all cells in which it is actively expressed, with weights corresponding to the expression level. This alternative representation allows for the use of bipartite graph algorithms and conceptualization for analyses. For example, we used Barber’s bipartite modularity  [42] to assess the how the clustering performance of different algorithms vary with modularity of the underlying scRNA-seq dataset.
Notation Description
$\mathcal{H}$ Hypergraph
$\mathbf{I}_{\mathcal{H}}$ Incidence matrix of the hypergraph $\mathcal{H}$
$V$ Set of all nodes (cells)
$E$ Set of all hyperedges (genes) in the hypergraph
$G_{V}$ Cell co-expression network/ Node similarity matrix
$G_{E}$ Gene co-expression network/ Edge similarity matrix
$\Gamma_{e}(v)$ Weight of node $v$ in hyperedge $e$, i.e., expression of gene $e$ in cell $v$
$P(V=v\mid E=e)$ Probability of selecting node $v$ given hyperedge $e$
$P(E=e\mid V=v)$ Probability of selecting hyperedge $e$ given node $v$
$P(E_{t+1}=e^{\prime}\mid E_{t}=e,V_{t}=v)$ Probability of transitioning to hyperedge $e^{\prime}$ at time $t+1$, given current hyperedge $e$ and node $v$ at time $t$
$P(V_{t+1}=w\mid E_{t+1}=e,V_{t}=v)$ Probability of selecting node $w$ at time $t+1$, given hyperedge $e$ and previous node $v$
$P(V_{t+1}=w\mid V_{t}=v)$ Unipartite node-to-node transition probability
Table 1: Hypergraph Random Walk Notation Table.
II.3 Method 1: Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW)
We introduced a novel hypergraph random walk approach, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW), highlighting the consideration of both the relative importance of edges to nodes and nodes to edges, or importance of genes relative to cells and cell to genes, along with a preference exponent for faster clustering. In a hypergraph random walk process, the walker typically alternates between transitioning from a hyperedge to a node and from a node to a hyperedge, following the node-to-edge transition probabilities $P_{E|V}(e,v)$ and edge-to-node transition probabilities $P_{V|E}(v,e)$. In the hypergraph random walk with edge-dependent vertex weight (EDVW) by Chitra and Raphael (2019) [41], the edge-to-node transition probability is edge-dependent, but the node-to-edge transition probability does not consider the previously visited node, i.e., $P_{V|E}(v,e)=P_{E}(e)$ for all $v\in V$. As a natural extension to EDVW, we defined the node-to-edge transition probability to be vertex-dependent in DIPHW, and used a preference exponent to accelerate the clustering process. The modified random walk process is described as follows: at time $t$, assuming the random walker is at vertex $v_{t}$, the walker proceeds by iteratively selecting the next hyperedge and the subsequent vertex within that hyperedge to transition to, based on the following probabilities:
  1. 1.
    Node-to-Hyperedge Transition Probability:
    The probability of selecting hyperedge $e_{t+1}$ given that the walker is currently at vertex $v_{t}$ is defined as:
    $P_{E|V}(v_{t}\rightarrow e_{t+1})=\frac{\omega(e_{t+1})\gamma_{e_{t+1}}(v_{t})}{\sum_{e^{\prime}\in E(v_{t})}\omega(e^{\prime})\gamma_{e^{\prime}}(v_{t})}$
    In the context of scRNA-seq cell clustering, this cell-dependent gene selection probability depends on the relative expression level of a gene in a cell compared to the expression of other genes in the same cell.
  2. 2.
    Hyperedge-to-Node Transition Probability:
    Upon the walker’s arrival at hyperedge $e_{t+1}$, the probability of moving to vertex $v_{t+1}\in e_{t+1}$ is defined as:
    $P_{V|E}(e_{t+1}\rightarrow v_{t+1})=\frac{\gamma_{e_{t+1}}^{\alpha}(v_{t+1})}{\sum_{v^{\prime}\in e_{t+1}}\gamma_{e_{t+1}}^{\alpha}(v^{\prime})}$
    In the context of scRNA-seq cell clustering, the transition from a gene to a cell is based on the relative expression level of a gene in a cell, amplified by the preference exponent for faster clustering convergence, and normalized by the same gene’s expression in other cells.
  3. 3.
    Node-to-Node Transition Probability: Marginalizing over all possible hyperedges the walker can traverse gives the unipartite node-to-node transition probability of DIPHW. This probability is used for sampling the random walk node paths which capture the structural and relational information of the original hypergraph and are subsequently used as inputs to a neural network for cell embedding computation.
    $P(v_{t}\rightarrow v_{t+1})=\sum_{e\in E}P_{E|V}(e|v_{t})P_{V|E}(v_{t+1}|e).$
    Substituting the expressions for $P_{E|V}(e_{t+1}|v_{t})$ and $P_{V|E}(v_{t+1}|e_{t+1})$, we get:
    $P(v_{t}\rightarrow v_{t+1})=\sum_{e_{t+1}\in E}\frac{\omega(e_{t+1})\gamma_{e_{t+1}}(v_{t})}{\sum_{e^{\prime}\in E}\omega(e^{\prime})\gamma_{e^{\prime}}(v_{t})}\frac{\gamma_{e_{t+1}}^{\alpha}(v_{t+1})}{\sum_{v^{\prime}\in e_{t+1}}\gamma_{e_{t+1}}^{\alpha}(v^{\prime})}$
    The expression of node-to-node transition probability could be simplified by removing reference to any particular time points as,
    $P(u\rightarrow v)=\sum_{e\in E(u)}\frac{\omega(e)\gamma_{e}(u)}{\sum_{e^{\prime}\in E}\omega(e^{\prime})\gamma_{e^{\prime}}(u)}\frac{\gamma_{e}^{\alpha}(v)}{\sum_{u^{\prime}\in e}\gamma_{e}^{\alpha}(u^{\prime})}$
  4. 4.
    Node-to-Node Transition Probability in Vectorized Format:
    For efficient computation, the $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ could be written in the vectorized form as,
    $P=D_{E|V}^{-1}W_{E|V}D_{V|E}^{-1}W_{V|E}$
    The vectorization circumvents the need for explicit iteration and reduces the computation time. Here, $W_{E|V}=\mathbf{I}_{\mathcal{H}}^{T}W_{E}$ is the $|V|\times|E|$ node-to-edge transition weight matrix where column $e$ of $W_{E|V}$ is the vertex weights in hyperedge $e$ multiplied by the weight of hyperedge $e$. Similarly, $W_{V|E}$ represents the edge-to-node transition weight matrix before normalization, with dimensions $|E|\times|V|$. Each entry in this matrix is obtained by raising the corresponding edge-dependent vertex weight to the power determined by the preference exponent $\alpha$, expressed as $W_{V|E}=(\mathbf{I}_{\mathcal{H}}^{\alpha})$, where $w_{ve}=[\mathbf{I}_{\mathcal{H}}(e,v)]^{\alpha}$. Finally, $D_{E|V}$ and $D_{V|E}$ are diagonal normalizing matrices used to normalize the distribution of the transition from vertex-to-edge and edge-to-vertex respective so that the probability distribution sums to 1.
II.4 Method 2: co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk (CoMem-DIPHW)
In this section, we introduce the co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk (CoMem-DIPHW) designed to integrate scRNA-seq expression data and its co-expression networks for improved cell embedding computation. CoMem-DIPHW builds upon the Dual-Importance Preference Hypergraph Walk (DIPHW) and incorporates a memory mechanism to consider previously visited nodes and edges during transitions.
Hypergraph Representation and the unipartite network projections of the scRNA-seq data
We remark that a matrix can be directly interpreted as the incidence matrix of a hypergraph to represent the relationship between two variables, without the need for any additional computation. Moreover, we can construct two correlation networks as derivatives of the matrix. In each correlation network, the nodes correspond to one of the two variables, and the edges represent the correlations between the node pairs of the corresponding variable. These two unipartite correlation network projections embed global level information that summarizes the association between node pairs through their correlation.
The Node and Edge Similarity Networks: Cell and Gene co-expression Networks and their construction
Using scRNA-seq data, we construct two co-expression networks: the gene-co-expression network (GCN) $G_{E}$ and the cell-co-expression network (CCN) $G_{V}$. In the GCN, genes are represented as nodes, and edges correspond to correlation coefficients between gene pairs, indicating the similarity between their expression profiles. Similarly, in the CCN, cells are represented as nodes, and edges correspond to correlation coefficients between cell pairs. These two co-expression networks are unipartite projection networks of the original scRNA-seq expression data, with edge weights providing a summary statistic of pairwise similarity in expression profiles. Spearman’s rank coefficient was chosen to measure pairwise similarity of expression profiles as it is a non-parametric measure that captures non-linearity compared to Pearson’s correlation coefficient. The cell and gene co-expression networks respectively are the node and edge similarity networks derived from the hypergraph. These two unipartite correlation network projections embed global level information that summarizes the association between cell and gene pairs through their correlation.
Memory-Integrated Random Walk
Our proposed random walk process CoMem-DIPHW integrates a memory component by considering the similarity between consecutively visited nodes and edges. This memory mechanism adjusts the transition probabilities such that the node-to-edge transition probability depends on the previous edge visited, and the edge-to-node transition probability depends on the previous node visited. The node-node and edge-edge dependencies are incorporated through the node-node and edge-edge similarity networks, which are the two correlation networks derived from the original matrix. The memory incorporation allows both the local and global structures in the hypergraph to be captured by the walk. Moreover, the memory mechanism constrains the hypergraph walker to focus on network(cell-gene) interactions among similar nodes(cells) and edges(genes) to better uncover modularity within the data.
Capturing Global and Local Geometry
In our proposed CoMem-DIP-HW random walk, we leverage global information from unipartite projections—the gene-co-expression network $G_{E}$ and the cell-co-expression network $G_{V}$—along with local information from the hypergraph structure in the cell-gene scRNA-seq expression data to enhance clustering performance. The co-expression networks serve as node and edge similarity matrices, facilitating the memory-incorporated hypergraph random walk. For cell clustering and cell type identification in scRNA-seq data, cells are modeled as nodes and genes as hyperedges. The hypergraph represents the expression profiles of cells across genes, where each hyperedge (gene) connects to nodes (cells) in which it is expressed. The weight of connection in a hyperedge to each node (cell) is determined by the gene’s expression level in that cell. The cell co-expression network $G_{V}$ functions as the node similarity matrix, with $G_{V_{i,j}}$ representing the similarity between the expression profiles of cell pair $i$ and $j$. Similarly, the gene co-expression network $G_{E}$ serves as the edge similarity matrix, representing the similarity in expression profile between pairs of genes across different cells. Each entry in these matrices aggregates information from the expression profiles of pairs of cells or genes into pairwise similarities, providing a global perspective through unipartite network projections. However, specific expression levels of genes in individual cells, representing local information, are lost in these projected networks. Therefore, our proposed hypergraph random walk methods aim to integrate global information while preserving local information, to achieve more accurate cell type identification.
Formulation
The random walk process in CoMem-DIPHW is formulated as follows:
Let $G_{V}$ be the vertex similarity matrix, where $G_{V}(v,w)$ quantifies the similarity between vertices $v$ and $w$. Similarly, let $G_{E}$ be the edge similarity matrix, where each element $G_{E}(e,f)$ represents the similarity between edge $e$ and $f$.
Method 2.0: CoMem
  1. 1.
    The probability of selecting vertex $v$ in edge $e$ is:
    $P_{V|E}(e\rightarrow v)=\frac{\Gamma_{e}(v)}{\sum_{u\in e}\Gamma_{e}(u)}$
  2. 2.
    The probability of selecting edge $e$ given vertex $v$ is:
    $P_{E|V}(v\rightarrow e)=\frac{\Gamma_{e}(v)}{\sum_{e\in E}\Gamma_{e}(v)}$
  3. 3.
    The probability of transitioning from node $v$ in edge $e_{1}$ to edge $e_{2}$, considering edge similarity, is given by:
    $P_{E_{t+1}|V,E_{t}}(e_{2}|v,e_{1})=\frac{G_{E}(e_{1},e_{2})P(e_{2}|v)}{\sum_{e^{\prime}\in E}G_{E}(e_{1},e^{\prime})P(e^{\prime}|v)}$
  4. 4.
    Similarly, the probability of selecting node $w$ from edge $e$ after transitioning from node $v$, incorporating node similarity and the node-edge connection, is formulated as:
    $P_{V_{t+1}|E,V_{t}}(w|e,v)=\frac{G_{V}(v,w)P(w|e)}{\sum_{w^{\prime}\in e}G_{V}(v,w^{\prime})P(w^{\prime}|e)}$
Finally, the unipartite node-to-node transition probability of the multilayer network random walk process is:
$P_{V_{t}\rightarrow V_{t+1}(v,w)}$
$\propto$ $\sum_{e_{1},e_{2}\in E}P_{V|E}(v|e_{1})P_{E_{2}|V,E_{1}}(e_{2}|v,e_{1})P_{V_{t+1}|E,V_{t}}(w|e_{2},v)$
To encourage exploration and ensure non-lazy walks, we modify the node and edge similarity graphs such that the diagonal elements are set to zero. This ensures that during the random walk, the random walker does not remain at the same edge or node in consecutive steps.
Method 2.1: CoMem based on DIPHW
Further, for CoMem-DIPHW, we incorporate DIPHW proposed in section II.3 to CoMem walk II.4. We define
  • The probability of selecting vertex $v$ in edge $e$ is:
    $P_{V|E}(e\rightarrow v)=\frac{\gamma_{e}(v)^{\alpha}}{\sum_{v^{\prime}\in e}\gamma_{e}(v^{\prime})^{\alpha}}$
  • The probability of selecting edge $e$ given vertex $v$ is:
    $P_{E|V}(v\rightarrow e)=\frac{\omega(e)\gamma_{e}(v)}{\sum_{e^{\prime}\in E(v)}\omega(e^{\prime})\gamma_{e^{\prime}}(v)}$
    The other probability distributions are the same as defined in section II.4
II.5 Complexity Analysis
DIPHW
The computation cost of each element $P(u\rightarrow v)$ in the $|V|\times|V|$ node-to-node transition probability matrix is $O(|E|)$ when the intermediate matrices for the node-to-edge and edge-to-node transition probabilities $P_{E|V}(e|u)$ and $P_{V|E}(v|e)$ are precomputed. Constructing the entire $|V|\times|V|$ node-to-node transition probability matrix thus involves computing these probabilities for all node pairs, resulting in an overall time complexity of $O(|V|^{2}\cdot|E|)$.
For efficient computation, the $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ in vectorized form is given by:
$P_{V|V}=D_{E|V}^{-1}W_{E|V}D_{V|E}^{-1}W_{V|E}$
While the theoretical complexity remains $O(|E|^{2}\cdot|V|+|E|\cdot|V|^{2})$ for the vectorized approach, it utilizes parallel computation and is significantly faster and hence more suitable for large datasets. The memory complexity is $O(|E|\cdot|V|)$ for the node-to-edge and edge-to-node transition probability matrices.
CoMem-DIPHW
The time complexity of CoMem-DIPHW is $O(|E|^{2}\cdot|V|^{2})$, as for each node pair, the memory-incorporated transition traverses all possible edges that could have arrived at the first node and all possible edges that could then connect the fisrt node with the second node. The memory complexity is $O(|E|^{2}\cdot|V|+|E|\cdot|V|^{2})$ for the memory incorporated node-to-edge and edge-to-node transition probability matrices.
III Experiments and Results
III.1 Baseline Methods
To evaluate the performance of DIPHW and CoMem-DIPHW, we compared them to several widely used and established cell clustering methods, including five embedding-based methods (PCA [43], Node2Vec [44], EDVW [41], UMAP, and t-SNE) and four unipartite-graph-based community detection methods (Louvain, Greedy Modularity, Infomap, and Leiden). These methods are adopted by popular scRNA-seq analysis packages such as Seurat [15], WGCNA [14], and Scanpy [17]. We provide a summary of these methods in Table 2. For a fair comparison under consistent conditions, we focus on the core clustering algorithms rather than specific implementations in packages such as Seurat and Scanpy. This allows us to evaluate the clustering capabilities of each method independently of any additional preprocessing or optimization steps in the package implementations. The hyperparameters used for each experiment are detailed in Table S1 in the supplementary materials Supplementary Materials.
Method Type Data Input Packages
Greedy
Modularity [45]
Community Detection Unipartite network
(Cell co-expression network)
igraph [46]
Louvain [47] Community Detection Unipartite network
(Cell co-expression network)
Seurat [15], Scanpy [17], WGCNA [14]
Infomap [48] Community Detection Unipartite network
(Cell co-expression network)
igraph [46]
Leiden [49] Community Detection Unipartite network
(Cell co-expression network)
igraph [46]
PCA [43] Embedding Matrix (Genes by cells) Seurat [15], Scanpy [17], SC3 [50], Cell Ranger [26]
t-SNE [51] Embedding Matrix (Genes by cells) Seurat [15], Scanpy [17]
Node2Vec [44] Embedding Unipartite network
(Cell co-expression network)
Multiple implementations available [52, 53] .
UMAP [54] Embedding Matrix (Genes by cells) Seurat [15], Scanpy [17]
EDVW [41] +
word2vec [55]
Embedding Hypergraph incidence matrix
(Genes by cells)
Implemented based on Chitra and Raphael [41]; available in our GitHub repository [56]. Not in standard scRNA-seq packages.
DIPHW +
word2vec [55]
Embedding Hypergraph incidence matrix
(Genes by cells)
Our proposed method and implementation is available on GitHub [56]
CoMem +
word2vec [55]
Embedding Hypergraph incidence matrix
(Genes by cells);
Gene co-expression network;
Cell co-expression network
Our proposed method and implementation is available on GitHub [56]
Table 2: Comparison of Clustering Methods Evaluated in Experiments. For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
III.2 Data and Preprocessing
The data used in this study consists of both simulated and publicly available scRNA-seq datasets. The simulated data were generated using our proposed simulation algorithm described in detail in the Methods section, with code available on GitHub [56]. We have provided a detailed README file to guide users through the process of running the code and reproducing our results. For preprocessing, the raw single-cell RNA-seq data were first normalized to equalize total counts per cell to account for differences in sequencing depth among cells. After normalization, genes and cells with zero expression were removed. For empirical analysis, we retained only the top n most variable genes to focus on the most relevant signals. For empirical analysis, we used four publicly available scRNA-seq datasets: human brain [57], human pancreas [58], mouse brain [59], and mouse pancreas [60], and a curated benchmarking dataset with ground-truth cell line labels [61].
III.3 Sparsity-Induced Correlation Inflation in scRNA-seq Co-expression Networks
The zero inflation problem in single-cell RNA sequencing (scRNA-seq) data analysis refers to the excessive number of zero counts (or "dropouts") observed in the data. These zeros could either be a result of technical limitation in detecting low abundance transcripts (false negatives), or true absence biological zeros (true negatives) [37, 39, 40]. In scRNA-seq data, when individually profiled cell expression are merged to create a unified dataset, all genes are included in the expression profile of every cell. However, since cells typically express different sets of genes, the inclusion of all genes introduces significant sparsity into the merged dataset. Consequentially, when the scRNA-seq data are analyzed through co-expression networks, these induced zeros can inflate correlation coefficients between cells or genes. This occurs because during co-expression network construction, edge weights represent correlation between expression profiles and are often computed by Pearson’s correlation coefficient or Spearman’s rank correlation coefficient. In these commonly adopted correlation measures, zeros are treated as expression magnitudes in the same way as non-zeros. As a result, if two expression profiles share a large number of common zeros, their correlation will be inflated since the common zeros will be considered as evidence of their expression homogeneity, which can lead to an overestimation of the similarity between expression profiles. This issue is particularly pronounced at the single-cell level compared to bulk tissue RNAseq data, as the proportion of genes expressed at the single-cell level is much smaller than at the tissue level, resulting in a higher degree of induced data sparsity.
In the experiment shown in Fig. 3, we investigate the impact of induced sparsity in scRNA-seq data on correlation coefficients and its implications for interpreting gene and cell connections in co-expression networks. We demonstrate that induced zeros lead to inflated correlation coefficients between cells or genes. To simulate this effect, we generated gene expression data with a fixed base size and appended varying proportions of zeros to mimic different degrees of sparsity induced when merging cell expression profiles. We show that as the proportion of zeros appended to the data increases, the correlation coefficient also increases, leading to an overestimation of the similarity between expression profiles. Our results show that as the level of induced sparsity increases, the correlation coefficient also increases, leading to an overestimation of similarity between expression profiles. These findings highlight the importance of considering zero inflation in scRNA-seq data analysis and the need for accurate representation of the underlying data structures.
[…]
III.4 Simulation Analysis Results
In our study, we developed an algorithm to simulate data that mimics the characteristics of scRNA-seq expression data. In addition to the underlying assumption of within-cell-type homogeneous expression that governs most, if not all, cell clustering methods, we also incorporated between-cell-type crosstalk [62, 63, 64, 65, 66] by integrating configurable intermodular covariance, density, and signal strength into the simulation model. The algorithm generates a sparse matrix representing the simulated scRNA-seq expression data, with numerous user-configurable parameters, including number of modules, module density, shape, background signal strength, modular signal strength, intermodular signal strength, intermodular covariance, and noise level. This method is implemented as a function in our codebase and is available for use by other researchers.
Our simulation model is designed with high flexibility to accommodate different types of scRNA-seq data. For instance, the model can simulate both deeply sequenced datasets with high density (up to 50%) and shallowly sequenced datasets with low density (as low as 1%) [67]. The proportion of non-housekeeping genes can be adjusted by changing the number of differentially expressed genes in each module. Moreover, the variance in module sizes and the number of co-expressed genes in each cell type can be configured by setting different variances for the Poisson distribution used in the embedded module simulation. The modules embedded in the simulated data can be configured with higher density and average signal strength to model the expression profiles of biomarker genes that are highly expressed in certain cell types. This allows us to mimic the sparsity and modularity of real scRNA-seq data, where a small number of genes are highly expressed in specific cell types. Alternatively, the embedding modules can also be user-specified without using the embedded module simulation.
III.4.1 Evaluation Metrics
To evaluate the performance of the clustering methods on datasets with known ground-truth, we employed two widely used metrics for measuring coherence between the identified clusters and the ground-truth clusters: Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI). Both metrics are invariant to permutations, with a score of 0 representing random labeling and a score of 1 indicating identical clusters. For embedding-based methods that do not directly assign cluster membership, we used K-means to partition the embedded data into clusters.
[…]
III.4.2 Clustering Performance vs. Bipartite Modularity
We further evaluated the performance of these clustering methods under varying data modularity conditions. The modularity variation was controlled by two key parameters: the average number of co-expressed genes per cell type and the number of embedded cell modules. Additionally, we implemented Barber’s bipartite modularity [42], as shown by the barplots in Fig.4, to help quantify modularity of the underlying data for clustering and analyze the association between modularity and clustering performance. As shown in Fig.4, clustering performance generally improves with increased data modularity. The modularity bar plots in Fig.4 and heatmap visualization in Fig.S6 demonstrate that the simulated scRNA-seq data displays stronger modularity when the average number of co-expressed genes per cell type is greater or when the number of modules is smaller. Our proposed methods—DIPHW, CoMem, and CoMem-DIPHW—consistently outperformed all other clustering methods that use unipartite network projections, including Node2Vec, and community detection-based methods such as Louvain and Greedy Modularity. Notably, CoMem-DIPHW consistently demonstrated superior performance across almost all parameter settings, with its advantage being particularly pronounced under weak data modularity conditions.
[…]
III.4.3 Clustering Performance Comparison under strong and weak data modularity parameter regimes
To evaluate clustering performance of the proposed methods across modularity conditions, we considered two parameters that determine the modularity of the simulated scRNA-seq data: (i) the average number of co-expressed genes per module and (ii) the number of embedded modules. The number of modules here represents the number of cell types in the simulated scRNA-seq data. Weak modularity corresponds to a small number of co-expressed genes per module or a large number of embedded modules, making data structures more challenging to detect. Conversely, strong modularity corresponds to a large number of co-expressed genes per module or a small number of embedded modules. For clarity of presentation, we present the most contrasting cases from our analysis: (i) an average of 10 and 100 co-expressed genes per module in Fig. 5 and (ii) 10 and 100 embedded modules in Fig. 6. In Fig. 5 and Fig. 6, we compare the performance of our proposed hypergraph-based clustering methods (highlighted in red and yellow) against five embedding-based methods and four community detection methods under both strong and weak modularity conditions. All embedding-based methods are coupled with K-means to produce clustering results. In weak modularity regimes (where the average number of co-expressed genes = 10 or when number of modules = 100), CoMem-DIPHW and DIPHW exhibited a significant performance advantage—for example, when the average number of co-expressed genes = 10, CoMem-DIPHW had an ARI of 0.29 compared to 0.05 for the best-performing baseline method (PCA), representing a performance improvement of approximately 450%. Under strong modularity conditions (when the average number of genes per module = 100 or when number of modules = 10), several methods—including PCA, CoMem-DIPHW, CoMem, Node2Vec, and DIPHW—performed well, achieving average ARI scores above 0.9.
Additional results from intermediate parameter settings (average number of co-expressed genes per module: 20, 30, 40, 70, 80, 90; number of embedded modules: 20, 30, 40, 70, 80, 90) and by NMI support the same observation. Out of the 16 tested scenarios, CoMem-DIPHW or DIPHW consistently ranked first by ARI and in 10 scenarios by NMI. Our hypergraph-based methods showed significant advantages in weak modularity regimes, where data structures are harder to detect. Under comparatively strong modularity conditions, several methods (PCA, CoMem-DIPHW, CoMem, Node2Vec, DIPHW, and Louvain) performed well, with average NMI scores above 0.9. Notably, PCA slightly outperformed CoMem-DIPHW by NMI under strong modularity, with differences in NMI scores ranging from 0.01 to 0.02. These additional results are presented in Supplementary Figs. S2, Fig. S3, Fig. S4, Fig. S5.
III.4.4 Evaluating the Impact of Correlation Methods on Network-Based Clustering Performance
There are two primary strategies for handling shared zeros in cell expression profiles of scRNA-seq data during correlation computation. The first strategy, exemplified by cosine similarity, is to ignore the shared zeros. This means that inactive expression (zeros) does not contribute to the similarity between cells. However, our experiments indicate that this strategy results in significantly worse clustering performance. Understandably, the absence of expression in the same genes between cell pairs provides information about cellular similarity and should not be disregarded entirely.
In Table 3, we evaluated how co-expression networks generated by different correlation computation methods lead to varying performance across various network-based clustering methods. The correlation methods considered were Pearson’s, Spearman’s, Kendall’s Tau, and cosine similarity. Our results show that cosine similarity often led to the worst clustering results among the correlation computation methods tested.
Method ARI NMI
Pearson Cosine Spearman Kendall Pearson Cosine Spearman Kendall
GreedyModularity 0.205 0.135 0.194 0.197 0.534 0.429 0.496 0.502
Louvain 0.493 0.322 0.445 0.473 0.784 0.666 0.768 0.781
Infomap 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
Leiden -0.001 -0.001 -$0$0.000 0.062 0.040 0.082 0.073
Multilevel 0.032 0.034 0.034 0.035 0.305 0.289 0.325 0.319
Eigenvector 0.002 0.000 0.003 0.005 0.097 0.060 0.082 0.091
Table 3: Impact of Correlation Methods on Network-Based Clustering Performance. The table evaluates the impact of correlation computation methods—Pearson’s, Spearman’s, Kendall’s Tau, and cosine similarity—on the performance of various network-based clustering methods. The correlation computation method resulting in the worst clustering performance is marked in bold. Cosine similarity, which ignores shared zeros in cell expression profiles, often resulted in the worst clustering performance. These results show that, although using correlation computation methods that treat zeros the same as non-zero signals could inflate correlations, ignoring these zeros, as in cosine similarity, can lead to worse performance.
The primary implication of this experiment is that both strategies have inherent drawbacks when dealing with the sparsity of scRNA-seq data. Ignoring zeros misses out on significant information, while treating them as homogeneous signals inflates correlations and introduces noise. Ignoring the shared zeros in cell expression profiles, as cosine similarity does, is not an effective strategy and results in the worst performance. Conversely, methods that consider all the shared zeros in the expression profiles as an indication of cell homogeneity —such as Pearson’s, Spearman’s, and Kendall’s Tau—neglect the fact that many shared zeros are merely a result of the high dimensionality of scRNA-seq data. When sequencing results from multiple cells are merged, many zero entries are introduced since different cells often express different sets of genes. Consequently, cell correlations are inflated by these methods when data is sparse. The clustering outcomes from these methods are suboptimal when compared with approaches that do not require correlation computation, such as CoMem, DIPHW, and PCA, potentially due to misleading signals in the inflated correlation networks S1 4 5 6. Our results suggest that alternative approaches, such as hypergraph-based methods, provide a better solution by circumventing the need for unipartite network projections and avoiding the intricacy associated with handling zeros in correlation computations.
III.5 scMixology Benchmark Results
A major challenge in scRNA-seq analysis is the lack of reliable ground-truth for cell type annotation, which complicates the evaluation of clustering algorithms. In most scRNA-seq datasets, cell type labels are typically assigned using a combination of clustering results, differential expression (DE) analysis, and canonical marker genes. First, cells are grouped into clusters based on transcriptomic similarity using clustering algorithms (e.g., Leiden, Louvain, or k-means). Then, DE analysis is performed to identify genes that are differentially expressed across clusters. Finally, these clusters are annotated based on known canonical marker genes associated with specific cell types. Because this annotation process relies on clustering results, using these labels to evaluate clustering performance introduces circular logic and bias toward the method originally used for annotation. To address this issue, we utilized the mixture control benchmarking dataset [61], where ground truth is established based on cell line identity.
We evaluated the performance of our proposed method, CoMem-DIPHW, using two benchmarking datasets from scMixology [61]. These datasets provide ground-truth labels for cell type annotation based on human lung adenocarcinoma cell lines. As shown in Table 4, CoMem-DIPHW outperformed other methods, particularly on the 5-class dataset, achieving the highest NMI and ARI scores. In comparison, Leiden showed strong performance on the 3-class dataset but underperformed significantly on the 5-class dataset. Among the methods evaluated, PCA, CoMem, Node2Vec and CoMem-DIPHW consistently demonstrated robust performance across both datasets.
Dataset Greedy Modularity Louvain Infomap Leiden PCA t-SNE Node2Vec UMAP EDVW DIPHW CoMem CoMem-DIPHW
scMixology ARI 0.5836 0.1976 0.2327 0.9933 0.9900 0.9898 0.9900 0.9898 0.6619 0.9868 0.9933 0.9900
scMixology NMI 0.7193 0.3818 0.3843 0.9864 0.9797 0.9810 0.9797 0.9810 0.6609 0.9743 0.9864 0.9797
scMixology 5 class ARI 0.6645 0.3374 0.2998 0.6746 0.9873 0.9867 0.9461 0.9838 0.6873 0.9373 0.9880 0.9923
scMixology 5 class NMI 0.7852 0.4498 0.4355 0.7917 0.9747 0.9748 0.9148 0.9707 0.7740 0.9001 0.9770 0.9849
Table 4: Comparison of ARI and NMI across Clustering Methods for the ScMixology [61] Benchmark Datasets. The table is divided into 5-class (top) and 3-class (bottom) results. Best values are in bold, second-best are underlined. While PCA also performs well on these simple clustering tasks, CoMem and CoMem-DIPHW show robust performance across both datasets and are expected to outperform on more complex, weakly modular data, based on the previous simulation results.
III.6 Human Pancreas Results
For datasets without ground-truth annotations—human pancreas, mouse pancreas, human brain, and mouse brain—we conducted differential expression (DE) analysis based on clustering results. We present our differential expression analysis results from the human pancreas dataset by CoMem-DIPHW and PCA in detail and include a brief overview of results from the other datasets in the Supplementary Materials Section  Supplementary Materials.
We report the top 10 differentially expressed genes (DEGs) from each human pancreas cell cluster identified by CoMem-DIPHW and PCA in Fig. 7a and Fig. 8a respectively. To evaluate the clustering performance of CoMem-DIPHW applied to human pancreas scRNA-seq data, we assessed the across-cluster differentiation in the average expression of the cluster-specific DEGs, as shown in Fig. 7b. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. To compare clustering performance between PCA and CoMem-DIPHW, we examined the expression patterns of cluster-specific DEGs identified by CoMem-DIPHW (Fig. 7b) and PCA (Fig. 8b). In these figures, each subplot corresponds to one cluster’s DEGs, with violin plots showing their average expression levels across all clusters. The DEGs identified by CoMem-DIPHW exhibit more cluster-specific expression patterns compared to DEGs found by PCA. For example, comparing the PCA clustering results in Fig. 8b, the average expression of the DEGs from PCA cluster 1 is higher in clusters 2, 5, and 6. Similarly, the DEGs from cluster 7 show higher average expression in cluster 6. In general, the average expression of the DEGs from the PCA clusters is less cluster-specific than the DEGs from the CoMem-DIPHW clusters.
Moreover, we used the canonical gene markers reported in the original study of the human pancreas scRNA-seq data [58] to tentatively label the cell clusters found by CoMem-DIPHW. The markers include GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells). Clusters with high expression of corresponding gene markers are assigned to their respective cell types. As shown in Fig. 7d, each cell-type specific marker was distinctly expressed in a cohesive and well-defined cluster in the embedding generated by CoMem-DIPHW. The results show that CoMem-DIPHW embeddings effectively distinguish different cell types. In contrast, as shown in Fig. 8d, PCA embeddings lacked clear separation between different cell types.
As shown in Fig. 7d, Cluster 0 was identified as acinar cells by the high expression of PRSS1. In Table 7a, the top 10 DEGs for cluster 0 include 5 genes involved in digestive enzyme production, a primary function of pancreatic acinar cells: CPA1, CPA2 [68], PLA2G1B [69], CTRB1 [70], and PRSS1 [71]. Clusters 2 and 7 were identified as alpha cells by the high expression of GCG. Alpha cells are responsible for glucagon secretion and blood glucose regulation [72, 73]. Markers for alpha cells among the DEGs in cluster 2 found in the PanglaoDB database include GLS, CRYBA2, FAP, LOXL4, MAFB, and RGS4. In cluster 7, alpha cell marker DEGs include LOXL4, PAX6, RGS4, and CRYBA2. Cluster 3 was classified as ductal cells by the high expression of KRT19. Six out of the 10 DEGs identified by CoMem-DIPHW are markers for ductal cells in the pancreas, including SLC4A4, ANXA4, KRT19, CFTR, CLDN1, and ONECUT2. Cluster 5 was classified as delta cells by the high expression of SST. Established markers strongly associated with delta cells in DEGs identified by CoMem-DIPHW include GAD2, PCSK1, LEPR, and CASR. Cluster 6 was classified as beta cells by the high expression of INS. This is also strongly supported by the DEGs MAFA, ERO1LB, PDX1, IAPP, SYT13, HADH and ADCYAP1. Both COL1A1 and ESAM are highly expressed in cluster 4, suggesting mesenchymal or endothelial cells. These align well with 9 out of the top 10 DEGs found by CoMem-DIPHW, including COL4A2, COL4A1, SPARC, EMP1, IGFBP4, HTRA1, PXDN, LAMA4 and NID1.
To further validate the tentative cell type annotations and clustering results, we cross-referenced the cluster-specific DEGs (see Fig.7a) with markers from the PanglaoDB cell type marker database (version 27 Mar 2020)[74]. The assignment of cell types was determined based on the overlap between the DEGs of each cluster and the cell type-specific markers in the PanglaoDB marker database, with a match score computed based on the proportion of matched markers. Cell types with a score above a threshold of 0.05 were considered potential matches. Table 7c presents the top matches for each cluster. These results align well with our previous analysis based on marker gene distribution (Figure 7d), providing additional support for our cell type assignments. The only discrepancy between the results based on PanglaoDB and the marker gene distribution experiments lies in cluster 4. Matching the DEGs with PanglaoDB cell type markers suggests Pancreatic stellate cells (PSCs) or Granulosa cells, while the COL1A1 and ESAM expression plots indicate mesenchymal or endothelial cells. However, this apparent discrepancy can be explained by the fact that PSCs can differentiate into cells with endothelial-like properties [75], and granulosa cells can exhibit mesenchymal-like behaviors [76]. Both cell types share characteristics with mesenchymal and endothelial cells, explaining the overlap in gene expression patterns observed in cluster 4. Hence, the cell clusters determined by CoMem-DIPHW and annotated by the canonical markers selected by the original study [58] are also supported by the strong overlap between each cluster’s top DEGs and cell type-specific markers from the PanglaoDB database.
On the other hand, the differential analysis based on the PCA clustering results on the human pancreas dataset (Fig.8) shows limited cluster separation. While some of the PCA clusters could be consistently annotated with cell types based on evidence from both PanglaoDB markers and canonical markers, considerable overlap remains among PCA clusters. For example, cluster 8 is identified as beta cells by both the PanglaoDB markers and the canonical marker INS. However, cluster 8 is not well separated from cluster 5, which potentially represents delta cells based on its high expression of SST. Similarly, cluster 7 is identified as ductal cells by both the PanglaoDB markers and the canonical marker KRT19, while the endothelial cells marked by the high expression of ESAM are embedded very close to cluster 7 in the PCA embedding.
Figure 7: CoMem-DIPHW Clustering Performance Assessment on the Humen Pancreas Dataset and Cell Type Annotation using Differentially Expressed Genes (DEGs) and Canonical Markers. (a) Top 10 DEGs identified for each cluster by CoMem-DIPHW. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers. (d) Expression of human pancreatic canonical markers in CoMem-DIPHW embeddings. UMAP visualizations show CoMem-DIPHW cell embeddings, with cluster IDs assigned by K-means clustering. Each subplot displays the expression of a canonical marker gene for a specific human pancreatic cell type. The cell-type specific markers considered include: GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells). Color intensity represents gene expression levels, with redder hues indicating higher expression. Clusters show high expression of cell-type-specific canonical markers are assigned to the corresponding cell types. Overall, the clustering quality is demonstrated through: high within-cluster expression of DEGs shown in the violin plots in (b), distinct separation of cell types in (d), and consistency between canonical marker-based annotation in (d) and PanglaoDB-based annotation in (c).
Figure 8: PCA Performance Assessment on the Humen Pancreas Dataset and Cluster Cell Type Annotation using DEGs and Canonical Markers. (a) Top 10 DEGs identified for each cluster by PCA. (b) Across-cluster average expression of cluster-specific DEGs. (c) Cell type annotation using the PanglaoDB marker database. (d) Expression of human pancreatic canonical markers in PCA embeddings. The cell-type specific markers considered include: GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells). UMAP visualizations show PCA cell embeddings, with cluster IDs assigned by K-means clustering. The overall cluster or cell type separation is limited compared to CoMem-DIPHW clustering, as evidenced by the overlapping beta cells, delta cells and PP cells in (d) and the lack of distinct high expression patterns of the DEGs in (b).
Following discussion of the results from human pancreas dataset, we provide a brief overview of the results from the mouse pancreas, human brain, and mouse brain datasets. For these datasets, we have included visualizations in the appendix that show the average expression of the top DEGs across clusters identified by different methods, which shows our proposed algorithm’s ability in effectively distinguishing cell clusters by their expression profiles. The supplements also contain tables listing the top DEGs for clusters identified by CoMem-DIPHW and PCA in each dataset for reference.
IV Conclusion
In this study, we developed and analyzed two hypergraph-based clustering algorithms, DIPHW and CoMem-DIPHW for improved cell clustering performance on scRNA-seq data. Firstly, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW) we developed extends the edge-dependent vertex weight (EDVW) hypergraph random walk proposed by [41], and consistently outperformed it in our experiments. DIPHW’s key idea lies in its symmetrized account of importance: it considers both the relative importance of edges to nodes and nodes to edges, or importance of genes relative to cells and cell to genes, along with a preference exponent for faster clustering.
Secondly, the CoMem and CoMem-DIPHW clustering methods we developed incorporate a Memory mechanism that considers both the abundance counts data (conceptualized as hypergraphs) and the two co-expression networks (unipartite network projections of the scRNA-seq abundance counts: gene co-expression networks and cell co-expression networks). This approach allows us to integrate both local and global information when computing embeddings for clustering.
Thirdly, we identified and showed with experiments 3 the issue of inflated signals in co-expression networks brought by sparsity of scRNA-seq data, which thus provided additional motivation for conceptualizing the scRNA-seq data as hypergraphs for embedding computation. The sparsity-induced correlation inflation leads to misleading cell connections in cell co-expression networks. Conventional correlation measures (Pearson’s, Spearman’s, Kendall’s Tau) either treat the common zeros in cell expression profiles as a sign of homogeneity or ignore them during correlation computation (as in cosine similarity). Neither strategy handles the intricacy of the data sparsity well enough to avoid suboptimal results. We showed in Tab. 3 that unipartite network-based clustering approaches perform poorly on correlation networks constructed using cosine similarity compared to using other correlation measures. Furthermore, unipartite network-based clustering methods generally underperform compared to hypergraph-based methods or PCA, which do not require transforming scRNA-seq count data into unipartite networks.
Fourthly, we compared our proposed methods with well-established clustering methods commonly used in scRNA-seq analysis packages. These include five embedding-based methods (PCA, Node2Vec, EDVW [41], UMAP, and t-SNE) and four community detection methods (Louvain, Greedy Modularity, Infomap, and Leiden). Our clustering performance comparison experiments span comprehensive parameter regimes, with CoMem-DIPHW and DIPHW consistently delivering superior performance, especially when data modularity is weak. In these weakly modular regimes, where clusters are hardly detectable, all the established methods we used for comparison show a much more pronounced decay in performance compared to CoMem-DIPHW. For empirical analysis, we tested our algorithms on 4 real scRNA-seq datasets and 2 benchmark datasets. In these experiments, CoMem-DIPHW consistently demonstrated superior performance in differentiating and classifying distinct cell types compared to the baseline methods.
Additionally, we designed a versatile simulation algorithm for scRNA-seq data that is highly customizable through its many user-specifiable parameters. This algorithm is designed to model a wide range of expression scenarios and modular structures. Notably, it incorporates intermodular co-expressions, which mimic cell-type cross-talk observed in real biological systems.
For future work, we plan to optimize the implementation of CoMem-DIPHW to reduce its memory complexity, enhancing its scalability for larger datasets. We are also interested in exploring alternative ways to incorporate the memory mechanism into the hypergraph random walk. Currently, CoMem employs a simple product approach, moving beyond the simple product used in CoMem could potentially uncover even more effective clustering strategies.
Supplementary Materials
Figure/Table Description Value
Clustering results on simulated scRNA-seq data using embedding-based and community detection Methods (Figure S1) Number of cells 3000
Number of genes 1000
Number of embedded modules 30
Average number of genes per module 30
Average number of cells per module 38
Target density of the gene expression matrix 0.03
Density within modules 0.3
Density between modules 0.1
Probability of inter-module connections 0.6
Average background expression 10
Average within-module expression 20
Average inter-module expression 10
Workers 30
Walk length 30
Embedding dimensions 30
Preference exponent 50
Performance comparison of clustering methods on simulated scRNA-seq data modularity (Figure 4 for average genes per cell type plots) Average number of genes per module {10, 20, 30, 40, 70, 80, 90, 100}
Other parameters same as Figure S1
Performance comparison of clustering methods on simulated scRNA-seq data modularity (Figure 4 for number of modules plots) Number of embedded modules {10, 20, 30, 40, 70, 80, 90, 100}
Average number of genes per module 50
Average number of cells per module ( # of genes / # of modules) +3
Other parameters same as Figure S1
Clustering performance comparison by ARI across varying module sizes (Figure 5) (Same as in Figure 4 for average genes per cell type plots)
Clustering performance comparison by ARI across varying module counts (Figure 6) (Same as in Figure 4 for number of modules)
Comparison of ARI and NMI across clustering methods for the ScMixology benchmark datasets (Table 4) Number of highly expressed genes used 500
Preference exponent 10
Workers 50
Walk length 10
Embedding dimensions 10
Human pancreas (Section III.6) Number of highly expressed genes used 1000
Number of clusters 9
Preference exponent 10
Workers 100
Walk length 50
Embedding dimensions 10
Mouse pancreas (Section Additional Results for Mouse Pancreas and Mouse Brain Datasets) Number of clusters 13
Other parameters same as for human pancreas analysis (Section III.6)
Human brain (Section Additional Results for Human Brain Dataset) Number of clusters 7
Other parameters same as for human pancreas analysis (Section III.6)
Mouse brain (Section Additional Results for Mouse Pancreas and Mouse Brain Datasets) Number of clusters 9
Other parameters same as for human pancreas analysis (Section III.6)
Table S1: Hyperparameters Used in Experiments.
.1 Embedding Visualizations
We present the visualizations and clustering performance of embeddings generated by DIPHW, CoMem, and CoMem-DIPHW in Fig. S1. We compare these against five other embedding-based methods (PCA, Node2Vec, EDVW [41], UMAP, and t-SNE) and four community detection methods (Louvain, Greedy Modularity, Infomap, and Leiden) applied to simulated single-cell RNA sequencing (scRNA-seq) data. Embedding-based methods compute representations of the cells in a lower-dimensional space while preserving their global geometry [77, 78, 79]. For these methods, we used UMAP for embedding visualization and K-means for clustering. In contrast, community detection methods identify clusters (communities) based on the structures of the unipartite networks projected from the scRNA-seq data. Although t-SNE and UMAP are typically not used as a standalone embedding algorithms, we have included them in the comparison for reference, as they are adopted by popular packages in single-cell analysis, such as Seurat [15, 16], Scanpy [17], SC3 [50], and Cell Ranger [26]. There has been ongoing discussions about using them as the main clustering methods in scRNA-seq analysis [80, 81]. We note that the strength of t-SNE and UMAP lies in preserving local structures, such as the nearest neighbors [80, 54], while UMAP can also preserve global distance [54]. However, they are not designed to group data into clusters with high internal cohesion. Therefore, they are generally used in conjunction with other clustering methods as visualizing tools that map computed embeddings to a 2D space for evaluation.
The results are compared with the ground-truth cell-type membership in the embedding visualization and in the NMI and ARI bar plots. For visualization, UMAP was applied to the embeddings produced by the dimensionality reduction algorithms. Each dot represents a cell, with its color indicating the ground-truth cell cluster membership and its position representing the embedding. As illustrated, DIPHW and CoMem-DIPHW demonstrated the highest degree of internal cohesion within clusters and the best cluster separation, as evidenced by both the embedding visualizations and the ARI and NMI scores. We also evaluate CoMem as a baseline method, which shares the same memory mechanism as CoMem-DIPHW but without the dual importance (edge-dependent vertex selection probability and vertex-dependent edge selection probability) or preference exponent. This comparison shows the improvements brought by integrating the hypergraph-based random walk DIPHW with the memory mechanism in CoMem.
The parameters used to simulate the scRNA-seq data in this experiment were chosen to capture the main characteristics and underlying assumptions of scRNA-seq analysis. Although these parameters may not perfectly match the values observed in real scRNA-seq data, they provide a challenging enough clustering task to differentiate the performance of the evaluated clustering algorithms. The inter-module density is set at 0.1, and inter-module connection probability at 0.6 to represent cell-type cross-talk [63, 66, 65, 64, 82], resulting in a weaker expression signal density than within cell-type expression. To distinguish cell-type-specific signals from background expression, we set the baseline gene expression level at 10 and the within-module average expression at 20.
Figure S1: Clustering Results on Simulated ScRNA-seq Data Using Embedding-based and Community Detection Methods. (a) UMAP visualization of the embedding-based clustering results, with colors indicating ground-truth cell cluster membership. (b) NMI and ARI bar plots comparing the performance of the embedding and community detection methods. Higher NMI and ARI values suggest greater consistency with ground-truth clusters, indicating better clustering performance. DIPHW and CoMem-DIPHW demonstrated the highest degree of internal cohesion within clusters and the best cluster separation, as evidenced by both the embedding visualizations and the ARI and NMI scores. For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
Additional Results for Clustering Performance Across Modularity
Figure S2: Clustering Performance Comparison by ARI Across Varying Module Sizes. Simulated scRNA-seq data were used for this evaluation. Here, we present the complementary results across intermediate parameter settings, which show consistent patterns: our proposed methods (highlighted in red and yellow) consistently ranked first by ARI in all scenarios, with a stronger advantage when data modularity is weak, i.e., when the average number of co-expressed genes per module is small. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest ARI values or best ARI rankings. For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
Figure S3: Clustering Performance Comparison by ARI Across Varying Module Counts. Simulated scRNA-seq data were used for this evaluation. Here, we present the complementary results across intermediate parameter settings, which show consistent patterns: our proposed methods (highlighted in red and yellow) consistently ranked first by ARI in all scenarios, with a stronger advantage when data modularity is weak, i.e., when the number of embedded modules in the simulated scRNA-seq data is greater. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest ARI values or best ARI rankings. For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
Figure S4: Clustering Performance Comparison by NMI Across Varying Module Sizes. Simulated scRNA-seq data were used for this evaluation. The NMI results support the same conclusion: when modularity is weak (i.e., when the average number of co-expressed genes per module is small), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest NMI values or best NMI rankings. For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
Figure S5: Clustering Performance Comparison by NMI Across Varying Module Counts. Simulated scRNA-seq data were used for this evaluation. The results by NMI support the same conclusion: when modularity is weak (i.e., when the number of modules is greater), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest NMI values or best NMI rankings. For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
Figure S6: Visualization of Modularity in Simulated scRNA-seq Data. Modularity increases with a higher average number of co-expressed genes per cell type and fewer modules. The number of modules models the number of cell types in the simulated data.
Additional Results for Human Brain Dataset
Figure S7: CoMem-DIPHW Clustering Performance Assessment on the Human Brain Dataset [57] and Cell Type Annotation using DEGs and the PanglaoDB Marker Database. (a) Top 10 DEGs identified for each cluster by CoMem-DIPHW. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
Figure S8: PCA Clustering Performance Assessment on the Human Brain Dataset [57] and Cell Type Annotation using DEGs and the PanglaoDB Marker Database. (a) Top 10 DEGs identified for each cluster by PCA. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
Additional Results for Mouse Pancreas and Mouse Brain Datasets
Figure S9: CoMem-DIPHW Clustering Performance Assessment on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and the PanglaoDB Marker Database. (a) Top 10 DEGs identified for each cluster by CoMem-DIPHW. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
Figure S10: PCA Clustering Performance Assessment on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and the PanglaoDB Marker Database. (a) Top 10 DEGs identified for each cluster by PCA. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters. (c) Cell type annotation using the PanglaoDB marker database. Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
Figure S11: CoMem-DIPHW Clustering Performance Assessment on the Mouse Brain Dataset [59] and Cell Type Annotation using DEGs and the PanglaoDB Marker Database. (a) Top 10 DEGs identified for each cluster by CoMem-DIPHW. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. (c) Cell type annotation using the PanglaoDB marker database.
Figure S12: PCA Clustering Performance Assessment on the Mouse Brain Dataset [59] and Cell Type Annotation using DEGs and the PanglaoDB Marker Database. (a) Top 10 DEGs identified for each cluster by PCA. (b) Across-cluster average expression of cluster-specific DEGs. Violin plots show the distribution of average expression levels of these DEGs across all clusters. (c) Cell type annotation using the PanglaoDB marker database.

### v3

@@
 Its applications include functional analysis  [1, 2, 3, 4] and interaction prediction in protein-protein interaction (PPI) networks [5, 6, 7], regulator identification and pathway analysis in gene regulatory networks [8, 9, 10], disease modeling, prediction, and intervention in epidemiology [11, 12, 13], and cell type identification in co-expression networks [14, 15, 16, 17].
-While these methods employed unipartite networks to aid analysis, it is often overlooked that unipartite networks are often not the most natural or information-preserving representation of the complex system we are interested in studying since many complex systems involve interactions from more than one type of entities.
-Examples include protein-protein interaction (PPI) networks where the nodes are proteins and edges could be physical contacts among proteins or functional interactions in biological pathways, gene regulatory networks (GRNs) where nodes represent genes and a directed edge exists if one gene regulates the expression of another gene, co-expression networks where the nodes are genes, cells, or tissues and the weighted edges indicate similarity in expression profile between node pairs.
-An example of a network that captures higher-order interactions between different entity types is the bipartite network, where nodes are partitioned by two separate groups, and edges only connect nodes from different groups.
-Bipartite networks are particularly suitable for modeling systems where two types of entities interact, such as authors and papers in a collaboration network [18], recommendation systems [19] where nodes represent users and the recommending items, and specifically in biology, plant-pollinators and predator-prey interaction networks [20] in community ecology studies, microbe-host interaction networks [21] in microbiology, drug–protein interaction networks [22] , and transcription factor-gene networks in genomics.
-However, instead of modelling these systems using network models that could capture higher order interactions, most often, a unipartite network projection is chosen such that the system could be analyzed with the abundant network analysis methods developed based on unipartite networks [23, 14].
-For example, in the bipartite representation of a drug interaction network, the two classes of nodes represent drugs and protein targets respectively.
-A drug node and a protein node are connected if the drug targets the protein.
-In its unipartite transformed network, a drug-drug interaction network could be constructed based on shared targets.
-Similarly, in the bipartite protein complex network, proteins and biological functions are two types of nodes that are connected if the proteins are involved in the biological function.
-Yet, in the unipartite representation of the protein complex network, the clique graph of the original bipartite graph is constructed where the only class of nodes are the proteins and the protein nodes are connected if they are in the same complex that performs a biological function.
-Unipartite representations fail to capture the biological functions associated with groups of proteins.
-While unipartite projections can reveal relationships between nodes of the same type in a bipartite network, they often obscure the specificity of these interactions.
+However, traditional methods represent the biology systems using unipartite networks, which only captures relationships (edges) among one type of entity (nodes), are often not the most natural or information-preserving representation of the complex systems since many involve interactions between multiple types of entities.
+Instead of modelling these systems using network models that could capture higher order interactions, most often, a unipartite network projection is chosen such that the system could be analyzed with the abundant network analysis methods developed based on unipartite networks [23, 14].
+For example, in protein-protein interaction (PPI) networks, unipartite representations capture relationships between proteins but fail to capture the biological functions associated with groups of proteins.
+Similarly, unipartite projections of gene regulatory networks connect genes based on regulatory relationships but do not capture relationships among different entities such as transcription factors or their binding sites.
 Single-cell RNA sequencing (scRNA-seq) has enabled the profiling of gene expression at the individual cell level [24, 25, 26, 27], whereas conventional bulk tissue RNA sequencing measures expression at the tissue level, averaging gene expression over an ensemble of cells.
@@
 Given the issues with unipartite projections, it is crucial to explore alternative network representations that can accurately reflect the scRNA-seq data.
-Hypergraphs offer a neat solution by representing multi-way relationships in scRNA-seq directly, without the need for further data projection.
-Moreover, each bipartite graph can also be represented as a hypergraph, by modelling one set of entities as nodes and the other as hyperedges, offering a generalized representation of high-order interactions.
-In this work, we propose a hypergraph-based approach to analyze scRNA-seq data to overcome the limitations of unipartite projections.
+Hypergraphs offer a neat solution by representing multiway relationships in scRNA-seq directly, without the need for further data projection.
 Figure 1: Network Representations of the scRNA-seq Data.
@@
 II Methods
-In scRNA-seq analysis, clustering cells based on co-expression networks projected from count data is a common approach for cell type identification.
-However, as discussed in III.3, the correlation inflation problem associated with co-expression networks caused by the induced sparsity in scRNA-seq data, as well as information loss from projecting a hypergraph to a unipartite simple graph, motivated us to look beyond co-expression networks for scRNA-seq analysis.
-We conceptualize scRNA-seq data as a hypergraph $\mathcal{H}$.
-In this framework, we propose a hypergraph random walk-based clustering method that models the relationships between cells and genes in the cell-gene expression abundance count data.
-Specifically, each cell is represented as a node, and each gene is represented as a hyperedge.
-Each hyperedge captures the expression profile of a gene across all cells, allowing us to apply hypergraph based analysis tools without enduring the loss of information caused by projecting the scRNA-seq data into a simple unipartite graph, e.g.
-the co-expression networks.
-We design random walk algorithms that capture cell-cell interactions by sampling the cell-to-gene and gene-to-cell paths in the conceptualized hypergraph.
-The random walk sampling probability distribution is designed to account for the specific cell-gene expressions via the edge-dependent vertex selection probability and the vertex-dependent edge selection probability.
-The sampled paths embed information regarding the amplified expression level of a gene in a cell, both in comparison to the expression levels of other genes in that cell and to the expression levels of the same gene in other cells.
-As a step further, CoMem-DIPHW (the co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk), the second hypergraph-based random walk algorithm proposed for clustering, incorporates a memory mechanism that takes into account both the conceptualized hypergraph from the scRNA-seq expression data, and the two projected unipartite co-expression networks, i.e.
-the cell and gene co-expression network, for cell-cell interaction sampling.
-The sampled cell-cell interaction paths would then be fed to a small language model (Word2Vec) to generate efficient numerical representations (embeddings) of the cells.
-These embeddings would then be used to cluster the cells by applying a clustering algorithm such as UMAP or t-SNE to group the cells based on their similarity in the embedded space.
-We emphasize that this hypergraph conceptualization of the scRNA-seq data does not require further transformation on the data and consequently would not result in any information loss or computation cost from converting the scRNA-seq into co-expression networks.
-We remark that a matrix is a 2-dimensional array structured for organizing information related to two variables.
-Thus a matrix could naturally be conceptualized as a hypergraph by considering the column variables as nodes and the row variables as edges and vice versa.
+We conceptualize scRNA-seq data as a hypergraph $\mathcal{H}$ and design two hypergraph random walk algorithms that capture cell-cell relationships by sampling the cell-to-gene and gene-to-cell interactions in the conceptualized hypergraph.
+We emphasize that this hypergraph representation of the scRNA-seq data does not require further transformation of the data and consequently would not result in any information loss or additional computational cost.
 II.1 Hypergraph representation
@@
   • $E$ denotes the hyperedge set representing genes, with each hyperedge $e\in E$ connecting to all cells $\in V$ in which the gene $e$ is actively expressed.
+The weight of connection of a hyperedge (gene) to each node (cell) is determined by the gene’s expression level in that cell.
+Thus, each hyperedge captures the expression profile of a gene across all cells, preventing any loss of information.
 The incidence matrix $\mathbf{I}_{\mathcal{H}}$ of the hypergraph $\mathcal{H}$ is constructed such that $\mathbf{I}_{\mathcal{H}}(v,e)$ is given by the expression level, i.e.
 abundance counts, of gene $e$ in cell $v$.
-II.2 Bipartite representation
-We chose the hypergraph conceptualization over bipartite representation for consistency since the DIPHW algorithm we introduced is an extension of the edge-dependent vertex weight hypergraph random walk (EDVW) [41].
-DIPHW extended EDVW hypergraph random walk [41] with a vertex-dependent edge selection probability, as well as a preference exponent for clustering acceleration.
-A bipartite conceptualization could be adopted for our proposed methods DIPHW, CoMem, and CoMem-DIPHW by considering the incidence matrix $\mathbf{I}_{\mathcal{H}}$ of the hypergraph $\mathcal{H}$ as a bipartite graph.
-The two sets of nodes in the bipartite representation are the set of cells V and the set of genes E.
-Similarly, each gene is connected to all cells in which it is actively expressed, with weights corresponding to the expression level.
-This alternative representation allows for the use of bipartite graph algorithms and conceptualization for analyses.
-For example, we used Barber’s bipartite modularity  [42] to assess the how the clustering performance of different algorithms vary with modularity of the underlying scRNA-seq dataset.
 Notation Description
@@
 Table 1: Hypergraph Random Walk Notation Table.
-II.3 Method 1: Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW)
-We introduced a novel hypergraph random walk approach, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW), highlighting the consideration of both the relative importance of edges to nodes and nodes to edges, or importance of genes relative to cells and cell to genes, along with a preference exponent for faster clustering.
-In a hypergraph random walk process, the walker typically alternates between transitioning from a hyperedge to a node and from a node to a hyperedge, following the node-to-edge transition probabilities $P_{E|V}(e,v)$ and edge-to-node transition probabilities $P_{V|E}(v,e)$.
-In the hypergraph random walk with edge-dependent vertex weight (EDVW) by Chitra and Raphael (2019) [41], the edge-to-node transition probability is edge-dependent, but the node-to-edge transition probability does not consider the previously visited node, i.e., $P_{V|E}(v,e)=P_{E}(e)$ for all $v\in V$.
+II.2 Method 1: Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW)
+Building on this hypergraph representation, we introduced a novel hypergraph random walk approach for cell embedding computation, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW), which accounts for both the relative importance of edges to nodes and nodes to edges, or importance of genes relative to cells and cell to genes, along with a preference exponent for faster clustering.
+In a hypergraph random walk process, the walker alternates between transitioning from a hyperedge to a node and from a node to a hyperedge, following the node-to-edge transition probabilities $P_{E|V}(e,v)$ and edge-to-node transition probabilities $P_{V|E}(v,e)$.
+In the hypergraph random walk with edge-dependent vertex weight (EDVW) by Chitra and Raphael (2019) [41], the edge-to-node transition probability is edge-dependent, but the node-to-edge transition probability does not consider the importance of the edge to the node.
 As a natural extension to EDVW, we defined the node-to-edge transition probability to be vertex-dependent in DIPHW, and used a preference exponent to accelerate the clustering process.
-The modified random walk process is described as follows: at time $t$, assuming the random walker is at vertex $v_{t}$, the walker proceeds by iteratively selecting the next hyperedge and the subsequent vertex within that hyperedge to transition to, based on the following probabilities:
+The modified random walk process is described as follows:
   1. 1.
@@
     $P_{E|V}(v_{t}\rightarrow e_{t+1})=\frac{\omega(e_{t+1})\gamma_{e_{t+1}}(v_{t})}{\sum_{e^{\prime}\in E(v_{t})}\omega(e^{\prime})\gamma_{e^{\prime}}(v_{t})}$
-    In the context of scRNA-seq cell clustering, this cell-dependent gene selection probability depends on the relative expression level of a gene in a cell compared to the expression of other genes in the same cell.
+    In scRNA-seq cell clustering, this represents the probability of selecting a gene based on its expression level relative to other genes in the same cell.
   2. 2.
     Hyperedge-to-Node Transition Probability:
-    Upon the walker’s arrival at hyperedge $e_{t+1}$, the probability of moving to vertex $v_{t+1}\in e_{t+1}$ is defined as:
+    Upon the walker’s arrival at hyperedge $e_{t+1}$, the probability of moving to vertex $v_{t+1}\in e_{t+1}$ is given by:
     $P_{V|E}(e_{t+1}\rightarrow v_{t+1})=\frac{\gamma_{e_{t+1}}^{\alpha}(v_{t+1})}{\sum_{v^{\prime}\in e_{t+1}}\gamma_{e_{t+1}}^{\alpha}(v^{\prime})}$
-    In the context of scRNA-seq cell clustering, the transition from a gene to a cell is based on the relative expression level of a gene in a cell, amplified by the preference exponent for faster clustering convergence, and normalized by the same gene’s expression in other cells.
+    In scRNA-seq cell clustering, the transition from a gene to a cell is based on the relative expression level of a gene in a cell, amplified by a preference exponent to improve clustering convergence.
   3. 3.
-    Node-to-Node Transition Probability: Marginalizing over all possible hyperedges the walker can traverse gives the unipartite node-to-node transition probability of DIPHW.
-This probability is used for sampling the random walk node paths which capture the structural and relational information of the original hypergraph and are subsequently used as inputs to a neural network for cell embedding computation.
+    Node-to-Node Transition Probability:
+    The unipartite node-to-node transition probability of DIPHW is obtained by summing the product of the node-to-edge and edge-to-node transition probabilities over all hyperedges.
+This probability is used to sample random walk node paths, which serve as inputs to a neural network for cell embedding computation.
     $P(v_{t}\rightarrow v_{t+1})=\sum_{e\in E}P_{E|V}(e|v_{t})P_{V|E}(v_{t+1}|e).$
@@
     Node-to-Node Transition Probability in Vectorized Format:
-    For efficient computation, the $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ could be written in the vectorized form as,
+    The $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ could be written in the vectorized form as,
     $P=D_{E|V}^{-1}W_{E|V}D_{V|E}^{-1}W_{V|E}$
-    The vectorization circumvents the need for explicit iteration and reduces the computation time.
-Here, $W_{E|V}=\mathbf{I}_{\mathcal{H}}^{T}W_{E}$ is the $|V|\times|E|$ node-to-edge transition weight matrix where column $e$ of $W_{E|V}$ is the vertex weights in hyperedge $e$ multiplied by the weight of hyperedge $e$.
-Similarly, $W_{V|E}$ represents the edge-to-node transition weight matrix before normalization, with dimensions $|E|\times|V|$.
-Each entry in this matrix is obtained by raising the corresponding edge-dependent vertex weight to the power determined by the preference exponent $\alpha$, expressed as $W_{V|E}=(\mathbf{I}_{\mathcal{H}}^{\alpha})$, where $w_{ve}=[\mathbf{I}_{\mathcal{H}}(e,v)]^{\alpha}$.
-Finally, $D_{E|V}$ and $D_{V|E}$ are diagonal normalizing matrices used to normalize the distribution of the transition from vertex-to-edge and edge-to-vertex respective so that the probability distribution sums to 1.
-II.4 Method 2: co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk (CoMem-DIPHW)
-In this section, we introduce the co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk (CoMem-DIPHW) designed to integrate scRNA-seq expression data and its co-expression networks for improved cell embedding computation.
-CoMem-DIPHW builds upon the Dual-Importance Preference Hypergraph Walk (DIPHW) and incorporates a memory mechanism to consider previously visited nodes and edges during transitions.
-Hypergraph Representation and the unipartite network projections of the scRNA-seq data
-We remark that a matrix can be directly interpreted as the incidence matrix of a hypergraph to represent the relationship between two variables, without the need for any additional computation.
-Moreover, we can construct two correlation networks as derivatives of the matrix.
-In each correlation network, the nodes correspond to one of the two variables, and the edges represent the correlations between the node pairs of the corresponding variable.
-These two unipartite correlation network projections embed global level information that summarizes the association between node pairs through their correlation.
+    Vectorization eliminates explicit iteration and reduces the computation time.
+Here, $W_{E|V}=\mathbf{I}_{\mathcal{H}}^{T}W_{E}$ is the $|V|\times|E|$ node-to-edge transition weight matrix where column $e$ of $W_{E|V}$ contains the vertex weights in hyperedge $e$ multiplied by the weight of hyperedge $e$.
+Similarly, $W_{V|E}$ represents the edge-to-node transition weight matrix before normalization with dimensions $|E|\times|V|$, given by $W_{V|E}=\mathbf{I}_{\mathcal{H}}^{\alpha},\quad w_{ve}=\left[\mathbf{I}_{\mathcal{H}}(e,v)\right]^{\alpha}$.
+Finally, $D_{E|V}$ and $D_{V|E}$ are diagonal matrices that normalize vertex-to-edge and edge-to-vertex transitions to ensure the probability distribution sums to 1.
+II.3 Method 2: co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk (CoMem-DIPHW)
+We now introduce CoMem-DIPHW, an extension of DIPHW that incorporates co-expression networks and single-cell transcriptomic profiles.
+The memory mechanism of CoMem-DIPHW leverages cell and gene co-expression networks to account for previously visited nodes and edges during transitions to capture both local expression relationships and global co-expression patterns.
+Hypergraph Representation of the scRNA-seq data
+We remark that a matrix can be directly interpreted as the incidence matrix of a hypergraph to represent the relationship between two variables, without additional computation.
 The Node and Edge Similarity Networks: Cell and Gene co-expression Networks and their construction
-Using scRNA-seq data, we construct two co-expression networks: the gene-co-expression network (GCN) $G_{E}$ and the cell-co-expression network (CCN) $G_{V}$.
-In the GCN, genes are represented as nodes, and edges correspond to correlation coefficients between gene pairs, indicating the similarity between their expression profiles.
-Similarly, in the CCN, cells are represented as nodes, and edges correspond to correlation coefficients between cell pairs.
-These two co-expression networks are unipartite projection networks of the original scRNA-seq expression data, with edge weights providing a summary statistic of pairwise similarity in expression profiles.
-Spearman’s rank coefficient was chosen to measure pairwise similarity of expression profiles as it is a non-parametric measure that captures non-linearity compared to Pearson’s correlation coefficient.
-The cell and gene co-expression networks respectively are the node and edge similarity networks derived from the hypergraph.
-These two unipartite correlation network projections embed global level information that summarizes the association between cell and gene pairs through their correlation.
+Using scRNA-seq data, we construct two co-expression networks: the cell-co-expression network (CCN) $G_{V}$ and the gene-co-expression network (GCN) $G_{E}$.
+In the CCN, cells are nodes, and edges correspond to correlation coefficients between cell pairs, indicating the similarity between their expression profiles.
+Similarly, in the GCN, genes are nodes and edges correspond to correlation coefficients between gene pairs.
+Spearman’s rank coefficient was chosen to measure pairwise similarity of expression profiles as it captures non-linearity better than Pearson’s correlation.
+These unipartite networks serve as node and edge similarity matrices, embedding global-level information that summarizes associations between cell and gene pairs.
 Memory-Integrated Random Walk
-Our proposed random walk process CoMem-DIPHW integrates a memory component by considering the similarity between consecutively visited nodes and edges.
+CoMem-DIPHW integrates a memory component by considering the similarity between consecutively visited nodes and edges.
 This memory mechanism adjusts the transition probabilities such that the node-to-edge transition probability depends on the previous edge visited, and the edge-to-node transition probability depends on the previous node visited.
-The node-node and edge-edge dependencies are incorporated through the node-node and edge-edge similarity networks, which are the two correlation networks derived from the original matrix.
-The memory incorporation allows both the local and global structures in the hypergraph to be captured by the walk.
-Moreover, the memory mechanism constrains the hypergraph walker to focus on network(cell-gene) interactions among similar nodes(cells) and edges(genes) to better uncover modularity within the data.
-Capturing Global and Local Geometry
-In our proposed CoMem-DIP-HW random walk, we leverage global information from unipartite projections—the gene-co-expression network $G_{E}$ and the cell-co-expression network $G_{V}$—along with local information from the hypergraph structure in the cell-gene scRNA-seq expression data to enhance clustering performance.
-The co-expression networks serve as node and edge similarity matrices, facilitating the memory-incorporated hypergraph random walk.
-For cell clustering and cell type identification in scRNA-seq data, cells are modeled as nodes and genes as hyperedges.
-The hypergraph represents the expression profiles of cells across genes, where each hyperedge (gene) connects to nodes (cells) in which it is expressed.
-The weight of connection in a hyperedge to each node (cell) is determined by the gene’s expression level in that cell.
-The cell co-expression network $G_{V}$ functions as the node similarity matrix, with $G_{V_{i,j}}$ representing the similarity between the expression profiles of cell pair $i$ and $j$.
-Similarly, the gene co-expression network $G_{E}$ serves as the edge similarity matrix, representing the similarity in expression profile between pairs of genes across different cells.
-Each entry in these matrices aggregates information from the expression profiles of pairs of cells or genes into pairwise similarities, providing a global perspective through unipartite network projections.
-However, specific expression levels of genes in individual cells, representing local information, are lost in these projected networks.
-Therefore, our proposed hypergraph random walk methods aim to integrate global information while preserving local information, to achieve more accurate cell type identification.
+The node-node and edge-edge dependencies are incorporated through the node and edge similarity networks, which are the two correlation networks derived from the transcriptomic expression data.
+The memory mechanism constrains the hypergraph walker to focus on cell-gene interactions among similar nodes(cells) and edges(genes) to better uncover modularity within the data.
+While the co-expression networks provide global perspectives through unipartite projections, specific expression levels of genes in individual cells (local information) are preserved in the hypergraph structure.
+CoMem-DIPHW integrates both global and local information for more accurate cell type identification.
 Formulation
@@
     $P_{V_{t+1}|E,V_{t}}(w|e,v)=\frac{G_{V}(v,w)P(w|e)}{\sum_{w^{\prime}\in e}G_{V}(v,w^{\prime})P(w^{\prime}|e)}$
-Finally, the unipartite node-to-node transition probability of the multilayer network random walk process is:
-$P_{V_{t}\rightarrow V_{t+1}(v,w)}$
-$\propto$ $\sum_{e_{1},e_{2}\in E}P_{V|E}(v|e_{1})P_{E_{2}|V,E_{1}}(e_{2}|v,e_{1})P_{V_{t+1}|E,V_{t}}(w|e_{2},v)$
+Finally, the unipartite node-to-node transition probability of the hypergraph random walk process is:
+$P_{V_{t}\rightarrow V_{t+1}(v_{1},v_{2})}$
+$\propto$ $\sum_{e_{1},e_{2}\in E}P_{V|E}(v_{1}|e_{1})P_{E_{2}|V,E_{1}}(e_{2}|v_{1},e_{1})P_{V_{t+1}|E,V_{t}}(v_{2}|e_{2},v_{1})$
 To encourage exploration and ensure non-lazy walks, we modify the node and edge similarity graphs such that the diagonal elements are set to zero.
@@
 Method 2.1: CoMem based on DIPHW
-Further, for CoMem-DIPHW, we incorporate DIPHW proposed in section II.3 to CoMem walk II.4. We define
-  • The probability of selecting vertex $v$ in edge $e$ is:
-    $P_{V|E}(e\rightarrow v)=\frac{\gamma_{e}(v)^{\alpha}}{\sum_{v^{\prime}\in e}\gamma_{e}(v^{\prime})^{\alpha}}$
-  • The probability of selecting edge $e$ given vertex $v$ is:
-    $P_{E|V}(v\rightarrow e)=\frac{\omega(e)\gamma_{e}(v)}{\sum_{e^{\prime}\in E(v)}\omega(e^{\prime})\gamma_{e^{\prime}}(v)}$
-    The other probability distributions are the same as defined in section II.4
-II.5 Complexity Analysis
-DIPHW
-The computation cost of each element $P(u\rightarrow v)$ in the $|V|\times|V|$ node-to-node transition probability matrix is $O(|E|)$ when the intermediate matrices for the node-to-edge and edge-to-node transition probabilities $P_{E|V}(e|u)$ and $P_{V|E}(v|e)$ are precomputed.
-Constructing the entire $|V|\times|V|$ node-to-node transition probability matrix thus involves computing these probabilities for all node pairs, resulting in an overall time complexity of $O(|V|^{2}\cdot|E|)$.
-For efficient computation, the $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ in vectorized form is given by:
-$P_{V|V}=D_{E|V}^{-1}W_{E|V}D_{V|E}^{-1}W_{V|E}$
-While the theoretical complexity remains $O(|E|^{2}\cdot|V|+|E|\cdot|V|^{2})$ for the vectorized approach, it utilizes parallel computation and is significantly faster and hence more suitable for large datasets.
-The memory complexity is $O(|E|\cdot|V|)$ for the node-to-edge and edge-to-node transition probability matrices.
-CoMem-DIPHW
-The time complexity of CoMem-DIPHW is $O(|E|^{2}\cdot|V|^{2})$, as for each node pair, the memory-incorporated transition traverses all possible edges that could have arrived at the first node and all possible edges that could then connect the fisrt node with the second node.
-The memory complexity is $O(|E|^{2}\cdot|V|+|E|\cdot|V|^{2})$ for the memory incorporated node-to-edge and edge-to-node transition probability matrices.
+For CoMem-DIPHW, we integrate DIPHW into the CoMem walk framework from section II.3.
+Specifically, the vertex selection probability $P_{V|E}(e\rightarrow v)$ and the edge selection probability $P_{E|V}(v\rightarrow e)$ are both defined as in DIPHW (section II.2), while all other probability distributions remain as specified in section II.3.
 III Experiments and Results
@@
 For a fair comparison under consistent conditions, we focus on the core clustering algorithms rather than specific implementations in packages such as Seurat and Scanpy.
-This allows us to evaluate the clustering capabilities of each method independently of any additional preprocessing or optimization steps in the package implementations.
+This allows us to evaluate the clustering capabilities of each method independently of any additional preprocessing or optimization steps.
 The hyperparameters used for each experiment are detailed in Table S1 in the supplementary materials Supplementary Materials.
@@
 The simulated data were generated using our proposed simulation algorithm described in detail in the Methods section, with code available on GitHub [56].
-We have provided a detailed README file to guide users through the process of running the code and reproducing our results.
-For preprocessing, the raw single-cell RNA-seq data were first normalized to equalize total counts per cell to account for differences in sequencing depth among cells.
-After normalization, genes and cells with zero expression were removed.
-For empirical analysis, we retained only the top n most variable genes to focus on the most relevant signals.
+For preprocessing, we first normalized the raw single-cell RNA-seq data to account for differences in sequencing depth.
+Genes and cells with zero expression were then removed, and only the top n most variable genes were retained for analysis.
 For empirical analysis, we used four publicly available scRNA-seq datasets: human brain [57], human pancreas [58], mouse brain [59], and mouse pancreas [60], and a curated benchmarking dataset with ground-truth cell line labels [61].
@@
 The zero inflation problem in single-cell RNA sequencing (scRNA-seq) data analysis refers to the excessive number of zero counts (or "dropouts") observed in the data.
-These zeros could either be a result of technical limitation in detecting low abundance transcripts (false negatives), or true absence biological zeros (true negatives) [37, 39, 40].
-In scRNA-seq data, when individually profiled cell expression are merged to create a unified dataset, all genes are included in the expression profile of every cell.
-However, since cells typically express different sets of genes, the inclusion of all genes introduces significant sparsity into the merged dataset.
-Consequentially, when the scRNA-seq data are analyzed through co-expression networks, these induced zeros can inflate correlation coefficients between cells or genes.
-This occurs because during co-expression network construction, edge weights represent correlation between expression profiles and are often computed by Pearson’s correlation coefficient or Spearman’s rank correlation coefficient.
-In these commonly adopted correlation measures, zeros are treated as expression magnitudes in the same way as non-zeros.
-As a result, if two expression profiles share a large number of common zeros, their correlation will be inflated since the common zeros will be considered as evidence of their expression homogeneity, which can lead to an overestimation of the similarity between expression profiles.
-This issue is particularly pronounced at the single-cell level compared to bulk tissue RNAseq data, as the proportion of genes expressed at the single-cell level is much smaller than at the tissue level, resulting in a higher degree of induced data sparsity.
+These zeros could either be a result of technical limitation in detecting low abundance transcripts (false negatives), or true absence of biological zeros (true negatives) [37, 39, 40].
+The process of integrating individual cell expression profiles into a unified scRNA-seq dataset requires constructing a matrix that accounts for every gene detected in any cell, which artificially introduces zero entries wherever genes are not expressed in those cells.
+Since cells typically express different sets of genes, this inclusion introduces significant sparsity into the merged dataset.
+In co-expression network construction, edge weights are inflated due to these artificially introduced common zeros since the common zeros will be considered as evidence of their expression homogeneity.
+This issue is particularly pronounced at the single-cell level compared to bulk tissue RNAseq data, as the proportion of expressed genes at the single-cell level is much smaller, resulting in higher induced data sparsity.
 In the experiment shown in Fig.
-3, we investigate the impact of induced sparsity in scRNA-seq data on correlation coefficients and its implications for interpreting gene and cell connections in co-expression networks.
-We demonstrate that induced zeros lead to inflated correlation coefficients between cells or genes.
-To simulate this effect, we generated gene expression data with a fixed base size and appended varying proportions of zeros to mimic different degrees of sparsity induced when merging cell expression profiles.
-We show that as the proportion of zeros appended to the data increases, the correlation coefficient also increases, leading to an overestimation of the similarity between expression profiles.
-Our results show that as the level of induced sparsity increases, the correlation coefficient also increases, leading to an overestimation of similarity between expression profiles.
-These findings highlight the importance of considering zero inflation in scRNA-seq data analysis and the need for accurate representation of the underlying data structures.
+3, we investigate how induced sparsity affects correlation coefficients and the interpretation of gene and cell connections in co-expression networks.
+To simulate this effect, we generated gene expression data with a fixed base size and introduced varying proportions of zeros to mimic different sparsity levels.
+Our results show that as the level of induced sparsity increases, correlation coefficients between expression profiles become artificially inflated, leading to an overestimation of similarity between cell expression profiles.
+These findings highlight the importance of addressing zero inflation in scRNA-seq data analysis and the need for accurate representation of the underlying biological relationships.
+While induced sparsity inflates correlations, our supplementary experiments (.3) show that ignoring shared zeros using cosine similarity leads to worse clustering performance, as these inactive genes still hold biological significance.
 […]
@@
 […]
-III.4.2 Clustering Performance vs. Bipartite Modularity
-We further evaluated the performance of these clustering methods under varying data modularity conditions.
+III.4.2 Clustering Performance vs. Bipartite Modularity and Method Comparison
+We evaluated the performance of these clustering methods under varying data modularity conditions.
 The modularity variation was controlled by two key parameters: the average number of co-expressed genes per cell type and the number of embedded cell modules.
-Additionally, we implemented Barber’s bipartite modularity [42], as shown by the barplots in Fig.4, to help quantify modularity of the underlying data for clustering and analyze the association between modularity and clustering performance.
-As shown in Fig.4, clustering performance generally improves with increased data modularity.
-The modularity bar plots in Fig.4 and heatmap visualization in Fig.S6 demonstrate that the simulated scRNA-seq data displays stronger modularity when the average number of co-expressed genes per cell type is greater or when the number of modules is smaller.
+The number of modules here represents the number of cell types in the simulated scRNA-seq data.
+Weak modularity corresponds to a small number of co-expressed genes per module or a large number of embedded modules, making data structures more challenging to detect.
+The modularity bar plots in Fig.
+4 and heatmap visualization in Fig.
+S6 demonstrate that the simulated scRNA-seq data displays stronger modularity when the average number of co-expressed genes per cell type is greater or when the number of modules is smaller.
+We implemented Barber’s bipartite modularity [42], as shown by the barplots in Fig.
+4, to help quantify modularity of the underlying data for clustering and analyze the association between modularity and clustering performance.
+Clustering performance generally improves with increased data modularity.
 Our proposed methods—DIPHW, CoMem, and CoMem-DIPHW—consistently outperformed all other clustering methods that use unipartite network projections, including Node2Vec, and community detection-based methods such as Louvain and Greedy Modularity.
 Notably, CoMem-DIPHW consistently demonstrated superior performance across almost all parameter settings, with its advantage being particularly pronounced under weak data modularity conditions.
-[…]
-[…]
-III.4.3 Clustering Performance Comparison under strong and weak data modularity parameter regimes
-To evaluate clustering performance of the proposed methods across modularity conditions, we considered two parameters that determine the modularity of the simulated scRNA-seq data: (i) the average number of co-expressed genes per module and (ii) the number of embedded modules.
-The number of modules here represents the number of cell types in the simulated scRNA-seq data.
-Weak modularity corresponds to a small number of co-expressed genes per module or a large number of embedded modules, making data structures more challenging to detect.
-Conversely, strong modularity corresponds to a large number of co-expressed genes per module or a small number of embedded modules.
 For clarity of presentation, we present the most contrasting cases from our analysis: (i) an average of 10 and 100 co-expressed genes per module in Fig.
@@
 6.
+[…]
+[…]
 In Fig.
@@
 S5.
-III.4.4 Evaluating the Impact of Correlation Methods on Network-Based Clustering Performance
-There are two primary strategies for handling shared zeros in cell expression profiles of scRNA-seq data during correlation computation.
-The first strategy, exemplified by cosine similarity, is to ignore the shared zeros.
-This means that inactive expression (zeros) does not contribute to the similarity between cells.
-However, our experiments indicate that this strategy results in significantly worse clustering performance.
-Understandably, the absence of expression in the same genes between cell pairs provides information about cellular similarity and should not be disregarded entirely.
-In Table 3, we evaluated how co-expression networks generated by different correlation computation methods lead to varying performance across various network-based clustering methods.
-The correlation methods considered were Pearson’s, Spearman’s, Kendall’s Tau, and cosine similarity.
-Our results show that cosine similarity often led to the worst clustering results among the correlation computation methods tested.
-Method ARI NMI
-Pearson Cosine Spearman Kendall Pearson Cosine Spearman Kendall
-GreedyModularity 0.205 0.135 0.194 0.197 0.534 0.429 0.496 0.502
-Louvain 0.493 0.322 0.445 0.473 0.784 0.666 0.768 0.781
-Infomap 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
-Leiden -0.001 -0.001 -$0$0.000 0.062 0.040 0.082 0.073
-Multilevel 0.032 0.034 0.034 0.035 0.305 0.289 0.325 0.319
-Eigenvector 0.002 0.000 0.003 0.005 0.097 0.060 0.082 0.091
-Table 3: Impact of Correlation Methods on Network-Based Clustering Performance.
-The table evaluates the impact of correlation computation methods—Pearson’s, Spearman’s, Kendall’s Tau, and cosine similarity—on the performance of various network-based clustering methods.
-The correlation computation method resulting in the worst clustering performance is marked in bold.
-Cosine similarity, which ignores shared zeros in cell expression profiles, often resulted in the worst clustering performance.
-These results show that, although using correlation computation methods that treat zeros the same as non-zero signals could inflate correlations, ignoring these zeros, as in cosine similarity, can lead to worse performance.
-The primary implication of this experiment is that both strategies have inherent drawbacks when dealing with the sparsity of scRNA-seq data.
-Ignoring zeros misses out on significant information, while treating them as homogeneous signals inflates correlations and introduces noise.
-Ignoring the shared zeros in cell expression profiles, as cosine similarity does, is not an effective strategy and results in the worst performance.
-Conversely, methods that consider all the shared zeros in the expression profiles as an indication of cell homogeneity —such as Pearson’s, Spearman’s, and Kendall’s Tau—neglect the fact that many shared zeros are merely a result of the high dimensionality of scRNA-seq data.
-When sequencing results from multiple cells are merged, many zero entries are introduced since different cells often express different sets of genes.
-Consequently, cell correlations are inflated by these methods when data is sparse.
-The clustering outcomes from these methods are suboptimal when compared with approaches that do not require correlation computation, such as CoMem, DIPHW, and PCA, potentially due to misleading signals in the inflated correlation networks S1 4 5 6.
-Our results suggest that alternative approaches, such as hypergraph-based methods, provide a better solution by circumventing the need for unipartite network projections and avoiding the intricacy associated with handling zeros in correlation computations.
 III.5 scMixology Benchmark Results
@@
 Finally, these clusters are annotated based on known canonical marker genes associated with specific cell types.
-Because this annotation process relies on clustering results, using these labels to evaluate clustering performance introduces circular logic and bias toward the method originally used for annotation.
+Since this annotation process relies on clustering results, using these labels to evaluate clustering performance introduces circular logic and bias toward the method originally used for annotation.
 To address this issue, we utilized the mixture control benchmarking dataset [61], where ground truth is established based on cell line identity.
@@
 8a respectively.
-To evaluate the clustering performance of CoMem-DIPHW applied to human pancreas scRNA-seq data, we assessed the across-cluster differentiation in the average expression of the cluster-specific DEGs, as shown in Fig.
-7b.
-Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
-To compare clustering performance between PCA and CoMem-DIPHW, we examined the expression patterns of cluster-specific DEGs identified by CoMem-DIPHW (Fig.
+To evaluate clustering performance, we examined the expression of these DEGs across clusters (Fig.
 7b) and PCA (Fig.
 8b).
-In these figures, each subplot corresponds to one cluster’s DEGs, with violin plots showing their average expression levels across all clusters.
-The DEGs identified by CoMem-DIPHW exhibit more cluster-specific expression patterns compared to DEGs found by PCA.
-For example, comparing the PCA clustering results in Fig.
-8b, the average expression of the DEGs from PCA cluster 1 is higher in clusters 2, 5, and 6.
+Violin plots show that the DEGs identified by CoMem-DIPHW exhibit more cluster-specific expression patterns than PCA-derived DEGs, which tend to show higher expression in multiple clusters.
+For example, the average expression of the DEGs from PCA cluster 1 is higher in clusters 2, 5, and 6.
 Similarly, the DEGs from cluster 7 show higher average expression in cluster 6.
 In general, the average expression of the DEGs from the PCA clusters is less cluster-specific than the DEGs from the CoMem-DIPHW clusters.
-Moreover, we used the canonical gene markers reported in the original study of the human pancreas scRNA-seq data [58] to tentatively label the cell clusters found by CoMem-DIPHW.
-The markers include GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells).
-Clusters with high expression of corresponding gene markers are assigned to their respective cell types.
-As shown in Fig.
-7d, each cell-type specific marker was distinctly expressed in a cohesive and well-defined cluster in the embedding generated by CoMem-DIPHW.
-The results show that CoMem-DIPHW embeddings effectively distinguish different cell types.
-In contrast, as shown in Fig.
-8d, PCA embeddings lacked clear separation between different cell types.
-As shown in Fig.
+To tentatively label the cell clusters, we used canonical gene markers reported in the original human pancreas scRNA-seq study [58].
+We assigned clusters based on the expression of these canonical markers including GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells).
+Each marker is distinctly expressed in a well-defined cluster in the embedding generated by CoMem-DIPHW, as shown in Fig.
+7d, confirming that CoMem-DIPHW embeddings effectively distinguish different cell types.
+In contrast, PCA embeddings lacked clear separation between different cell types (Fig.
+8d).
+For example, in Fig.
 7d, cluster 0 was identified as acinar cells by the high expression of PRSS1.
-In Table 7a, the top 10 DEGs for cluster 0 include 5 genes involved in digestive enzyme production, a primary function of pancreatic acinar cells: CPA1, CPA2 [68], PLA2G1B [69], CTRB1 [70], and PRSS1 [71].
 Clusters 2 and 7 were identified as alpha cells by the high expression of GCG.
-Alpha cells are responsible for glucagon secretion and blood glucose regulation [72, 73].
-Markers for alpha cells among the DEGs in cluster 2 found in the PanglaoDB database include GLS, CRYBA2, FAP, LOXL4, MAFB, and RGS4.
-In cluster 7, alpha cell marker DEGs include LOXL4, PAX6, RGS4, and CRYBA2.
 Cluster 3 was classified as ductal cells by the high expression of KRT19.
-Six out of the 10 DEGs identified by CoMem-DIPHW are markers for ductal cells in the pancreas, including SLC4A4, ANXA4, KRT19, CFTR, CLDN1, and ONECUT2.
-Cluster 5 was classified as delta cells by the high expression of SST.
-Established markers strongly associated with delta cells in DEGs identified by CoMem-DIPHW include GAD2, PCSK1, LEPR, and CASR.
-Cluster 6 was classified as beta cells by the high expression of INS.
-This is also strongly supported by the DEGs MAFA, ERO1LB, PDX1, IAPP, SYT13, HADH and ADCYAP1.
 Both COL1A1 and ESAM are highly expressed in cluster 4, suggesting mesenchymal or endothelial cells.
-These align well with 9 out of the top 10 DEGs found by CoMem-DIPHW, including COL4A2, COL4A1, SPARC, EMP1, IGFBP4, HTRA1, PXDN, LAMA4 and NID1.
-To further validate the tentative cell type annotations and clustering results, we cross-referenced the cluster-specific DEGs (see Fig.7a) with markers from the PanglaoDB cell type marker database (version 27 Mar 2020)[74].
-The assignment of cell types was determined based on the overlap between the DEGs of each cluster and the cell type-specific markers in the PanglaoDB marker database, with a match score computed based on the proportion of matched markers.
-Cell types with a score above a threshold of 0.05 were considered potential matches.
-Table 7c presents the top matches for each cluster.
-These results align well with our previous analysis based on marker gene distribution (Figure 7d), providing additional support for our cell type assignments.
-The only discrepancy between the results based on PanglaoDB and the marker gene distribution experiments lies in cluster 4.
+Cluster 5 was classified as delta cells by the high expression of SST, while cluster 6 was classified as beta cells by the high expression of INS.
+The specific DEGs supporting each classification are detailed in Fig.
+7a.
+To further validate the cell type assignments and clustering results, we cross-referenced the cluster-specific DEGs (see Fig.7a) with markers from the PanglaoDB cell type marker database (version 27 Mar 2020)[74].
+The assignment of cell types was determined based on the overlap between the DEGs of each cluster and the cell type-specific markers in the PanglaoDB marker database, with a match score computed based on the proportion of matched markers (Table 7c).
+These results align well with our previous analysis based on marker gene distribution (Figure 7d).
+The only discrepancy lies in cluster 4.
 Matching the DEGs with PanglaoDB cell type markers suggests Pancreatic stellate cells (PSCs) or Granulosa cells, while the COL1A1 and ESAM expression plots indicate mesenchymal or endothelial cells.
 However, this apparent discrepancy can be explained by the fact that PSCs can differentiate into cells with endothelial-like properties [75], and granulosa cells can exhibit mesenchymal-like behaviors [76].
-Both cell types share characteristics with mesenchymal and endothelial cells, explaining the overlap in gene expression patterns observed in cluster 4.
-Hence, the cell clusters determined by CoMem-DIPHW and annotated by the canonical markers selected by the original study [58] are also supported by the strong overlap between each cluster’s top DEGs and cell type-specific markers from the PanglaoDB database.
-On the other hand, the differential analysis based on the PCA clustering results on the human pancreas dataset (Fig.8) shows limited cluster separation.
-While some of the PCA clusters could be consistently annotated with cell types based on evidence from both PanglaoDB markers and canonical markers, considerable overlap remains among PCA clusters.
-For example, cluster 8 is identified as beta cells by both the PanglaoDB markers and the canonical marker INS.
-However, cluster 8 is not well separated from cluster 5, which potentially represents delta cells based on its high expression of SST.
-Similarly, cluster 7 is identified as ductal cells by both the PanglaoDB markers and the canonical marker KRT19, while the endothelial cells marked by the high expression of ESAM are embedded very close to cluster 7 in the PCA embedding.
-Figure 7: CoMem-DIPHW Clustering Performance Assessment on the Humen Pancreas Dataset and Cell Type Annotation using Differentially Expressed Genes (DEGs) and Canonical Markers.
-(a) Top 10 DEGs identified for each cluster by CoMem-DIPHW.
+In contrast, PCA clustering results (Fig.8) showed limited cluster separation.
+While some PCA clusters could be annotated consistently using both PanglaoDB markers and canonical markers, significant overlap remained among clusters.
+For example, PCA cluster 8, identified as beta cells by both the PanglaoDB markers and the canonical marker INS, is not well separated from cluster 5 (potentially delta cells, marked by SST expression).
+Similarly, cluster 7, identified as ductal cells by both the PanglaoDB markers and the canonical marker KRT19, is closely embedded with endothelial cells (ESAM).
+These results highlight CoMem-DIPHW’s improved ability to distinguish pancreatic cell types compared to PCA.
+Figure 7: Clustering Performance of CoMem-DIPHW on the Human Pancreas Dataset and Cell Type Annotation using Differentially Expressed Genes (DEGs) and Canonical Markers.
+(a) Top 10 DEGs per cluster identified by CoMem-DIPHW.
+For example, Cluster 0 (acinar cells) includes genes associated with digestive enzyme production, a primary function of pancreatic acinar cells, including CPA1, CPA2 [68], PLA2G1B [69], CTRB1 [70], and PRSS1 [71].
+Clusters 2 and 7 (alpha cells) include markers GLS, CRYBA2, FAP, LOXL4, MAFB, and RGS4 (cluster 2) and LOXL4, PAX6, RGS4, and CRYBA2 (cluster 7).
+Cluster 3 (ductal cells) includes SLC4A4, ANXA4, KRT19, CFTR, CLDN1, and ONECUT2.
+Cluster 4, likely mesenchymal or endothelial cells, aligns with 9 of the top 10 DEGs found by CoMem-DIPHW, including COL4A2, COL4A1, SPARC, EMP1, IGFBP4, HTRA1, PXDN, LAMA4, and NID1.
+Cluster 5 (delta cells) includes GAD2, PCSK1, LEPR, and CASR, while Cluster 6 (beta cells) includes MAFA, ERO1LB, PDX1, IAPP, SYT13, HADH, and ADCYAP1.
+Cell type assignments are based on canonical marker expression patterns shown in (d).
 (b) Across-cluster average expression of cluster-specific DEGs.
 Violin plots show the distribution of average expression levels of these DEGs across all clusters.
-Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
+Strong clustering performance is indicated by high within-cluster expression of DEGs and low expression in other clusters.
 (c) Cell type annotation using the PanglaoDB marker database.
-Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
+Cell types are determined by the overlap between each cluster’s DEGs and PanglaoDB cell type-specific markers, with match scores computed based on the proportion of matched markers.
 (d) Expression of human pancreatic canonical markers in CoMem-DIPHW embeddings.
@@
 Each subplot displays the expression of a canonical marker gene for a specific human pancreatic cell type.
-The cell-type specific markers considered include: GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells).
-Color intensity represents gene expression levels, with redder hues indicating higher expression.
+The cell-type-specific markers considered include: GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells).
+Color intensity represents gene expression levels, with red hues indicating higher expression.
 Clusters show high expression of cell-type-specific canonical markers are assigned to the corresponding cell types.
-Overall, the clustering quality is demonstrated through: high within-cluster expression of DEGs shown in the violin plots in (b), distinct separation of cell types in (d), and consistency between canonical marker-based annotation in (d) and PanglaoDB-based annotation in (c).
-Figure 8: PCA Performance Assessment on the Humen Pancreas Dataset and Cluster Cell Type Annotation using DEGs and Canonical Markers.
-(a) Top 10 DEGs identified for each cluster by PCA.
+Overall, clustering quality is demonstrated through (b) high within-cluster DEG expression, (c) consistency between PanglaoDB-based annotation and clustering, and (d) distinct separation of cell types.
+Figure 8: Clustering Performance of PCA on the Human Pancreas Dataset and Cell Type Annotation using Differentially Expressed Genes (DEGs) and Canonical Markers.
+(a) Top 10 DEGs per cluster identified by PCA.
 (b) Across-cluster average expression of cluster-specific DEGs.
+Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
 (c) Cell type annotation using the PanglaoDB marker database.
 (d) Expression of human pancreatic canonical markers in PCA embeddings.
-The cell-type specific markers considered include: GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells).
+The cell-type-specific markers considered include: GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells).
 UMAP visualizations show PCA cell embeddings, with cluster IDs assigned by K-means clustering.
-The overall cluster or cell type separation is limited compared to CoMem-DIPHW clustering, as evidenced by the overlapping beta cells, delta cells and PP cells in (d) and the lack of distinct high expression patterns of the DEGs in (b).
-Following discussion of the results from human pancreas dataset, we provide a brief overview of the results from the mouse pancreas, human brain, and mouse brain datasets.
+The overall clustering of cell types is limited compared to CoMem-DIPHW clustering, as evidenced by the overlapping beta cells, delta cells and PP cells in (d) and the lack of distinct high expression patterns of the DEGs in (b).
+In addition to the human pancreas results presented here, we also analyzed mouse pancreas, human brain, and mouse brain datasets.
 For these datasets, we have included visualizations in the appendix that show the average expression of the top DEGs across clusters identified by different methods, which shows our proposed algorithm’s ability in effectively distinguishing cell clusters by their expression profiles.
@@
 DIPHW’s key idea lies in its symmetrized account of importance: it considers both the relative importance of edges to nodes and nodes to edges, or importance of genes relative to cells and cell to genes, along with a preference exponent for faster clustering.
-Secondly, the CoMem and CoMem-DIPHW clustering methods we developed incorporate a Memory mechanism that considers both the abundance counts data (conceptualized as hypergraphs) and the two co-expression networks (unipartite network projections of the scRNA-seq abundance counts: gene co-expression networks and cell co-expression networks).
-This approach allows us to integrate both local and global information when computing embeddings for clustering.
-Thirdly, we identified and showed with experiments 3 the issue of inflated signals in co-expression networks brought by sparsity of scRNA-seq data, which thus provided additional motivation for conceptualizing the scRNA-seq data as hypergraphs for embedding computation.
-The sparsity-induced correlation inflation leads to misleading cell connections in cell co-expression networks.
-Conventional correlation measures (Pearson’s, Spearman’s, Kendall’s Tau) either treat the common zeros in cell expression profiles as a sign of homogeneity or ignore them during correlation computation (as in cosine similarity).
-Neither strategy handles the intricacy of the data sparsity well enough to avoid suboptimal results.
-We showed in Tab.
-3 that unipartite network-based clustering approaches perform poorly on correlation networks constructed using cosine similarity compared to using other correlation measures.
-Furthermore, unipartite network-based clustering methods generally underperform compared to hypergraph-based methods or PCA, which do not require transforming scRNA-seq count data into unipartite networks.
+Secondly, the CoMem and CoMem-DIPHW clustering methods we developed incorporate a memory mechanism that considers both the abundance counts data (conceptualized as hypergraphs) and the two co-expression networks to integrate both local and global information when computing embeddings for clustering.
+Thirdly, we identified and showed with experiments 3 the issue of inflated signals in co-expression networks brought by sparsity of scRNA-seq data, which provides additional motivation for the use of hypergraph representations.
 Fourthly, we compared our proposed methods with well-established clustering methods commonly used in scRNA-seq analysis packages.
-These include five embedding-based methods (PCA, Node2Vec, EDVW [41], UMAP, and t-SNE) and four community detection methods (Louvain, Greedy Modularity, Infomap, and Leiden).
-Our clustering performance comparison experiments span comprehensive parameter regimes, with CoMem-DIPHW and DIPHW consistently delivering superior performance, especially when data modularity is weak.
-In these weakly modular regimes, where clusters are hardly detectable, all the established methods we used for comparison show a much more pronounced decay in performance compared to CoMem-DIPHW.
+Our clustering performance evaluation spanned a broad range of modularity conditions, where CoMem-DIPHW and DIPHW consistently exhibited superior performance, particularly in weak modularity regimes.
+When clusters were harder to detect, baseline methods experienced a much more pronounced decay in performance compared to CoMem-DIPHW.
 For empirical analysis, we tested our algorithms on 4 real scRNA-seq datasets and 2 benchmark datasets.
@@
 Table S1: Hyperparameters Used in Experiments.
-.1 Embedding Visualizations
+.1 Complexity Analysis
+DIPHW
+The computation cost of each element $P(u\rightarrow v)$ in the $|V|\times|V|$ node-to-node transition probability matrix is $O(|E|)$ when the intermediate matrices for the node-to-edge and edge-to-node transition probabilities $P_{E|V}(e|u)$ and $P_{V|E}(v|e)$ are precomputed.
+Constructing the entire $|V|\times|V|$ node-to-node transition probability matrix thus involves computing these probabilities for all node pairs, resulting in an overall time complexity of $O(|V|^{2}\cdot|E|)$.
+For efficient computation, the $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ in vectorized form is given by:
+$P_{V|V}=D_{E|V}^{-1}W_{E|V}D_{V|E}^{-1}W_{V|E}$
+While the theoretical complexity remains $O(|E|^{2}\cdot|V|+|E|\cdot|V|^{2})$ for the vectorized approach, it utilizes parallel computation and is significantly faster and hence more suitable for large datasets.
+The memory complexity is $O(|E|\cdot|V|)$ for the node-to-edge and edge-to-node transition probability matrices.
+CoMem-DIPHW
+The time complexity of CoMem-DIPHW is $O(|E|^{2}\cdot|V|^{2})$, as for each node pair, the memory-incorporated transition traverses all possible edges that could have arrived at the first node and all possible edges that could then connect the fisrt node with the second node.
+The memory complexity is $O(|E|^{2}\cdot|V|+|E|\cdot|V|^{2})$ for the memory incorporated node-to-edge and edge-to-node transition probability matrices.
+.2 Bipartite representation
+We chose the hypergraph conceptualization over bipartite representation for consistency since the DIPHW algorithm we introduced is an extension of the edge-dependent vertex weight hypergraph random walk (EDVW) [41].
+DIPHW extended EDVW hypergraph random walk [41] with a vertex-dependent edge selection probability, as well as a preference exponent for clustering acceleration.
+A bipartite conceptualization could be adopted for our proposed methods DIPHW, CoMem, and CoMem-DIPHW by considering the incidence matrix $\mathbf{I}_{\mathcal{H}}$ of the hypergraph $\mathcal{H}$ as a bipartite graph.
+The two sets of nodes in the bipartite representation are the set of cells V and the set of genes E.
+Similarly, each gene is connected to all cells in which it is actively expressed, with weights corresponding to the expression level.
+This alternative representation allows for the use of bipartite graph algorithms and conceptualization for analyses.
+For example, we used Barber’s bipartite modularity  [42] to assess the how the clustering performance of different algorithms vary with modularity of the underlying scRNA-seq dataset.
+.3 Evaluating the Impact of Correlation Methods on Network-Based Clustering Performance
+There are two primary strategies for handling shared zeros in cell expression profiles of scRNA-seq data during correlation computation.
+The first strategy, exemplified by cosine similarity, is to ignore the shared zeros.
+This means that inactive expression (zeros) does not contribute to the similarity between cells.
+However, our experiments indicate that this strategy results in significantly worse clustering performance.
+Understandably, the absence of expression in the same genes between cell pairs provides information about cellular similarity and should not be disregarded entirely.
+In Table S2, we evaluated how co-expression networks generated by different correlation computation methods lead to varying performance across various network-based clustering methods.
+The correlation methods considered were Pearson’s, Spearman’s, Kendall’s Tau, and cosine similarity.
+Our results show that cosine similarity often led to the worst clustering results among the correlation computation methods tested.
+Method ARI NMI
+Pearson Cosine Spearman Kendall Pearson Cosine Spearman Kendall
+GreedyModularity 0.205 0.135 0.194 0.197 0.534 0.429 0.496 0.502
+Louvain 0.493 0.322 0.445 0.473 0.784 0.666 0.768 0.781
+Infomap 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+Leiden -0.001 -0.001 -$0$0.000 0.062 0.040 0.082 0.073
+Multilevel 0.032 0.034 0.034 0.035 0.305 0.289 0.325 0.319
+Eigenvector 0.002 0.000 0.003 0.005 0.097 0.060 0.082 0.091
+Table S2: Impact of Correlation Methods on Network-Based Clustering Performance.
+The table evaluates the impact of correlation computation methods—Pearson’s, Spearman’s, Kendall’s Tau, and cosine similarity—on the performance of various network-based clustering methods.
+The correlation computation method resulting in the worst clustering performance is marked in bold.
+Cosine similarity, which ignores shared zeros in cell expression profiles, often resulted in the worst clustering performance.
+These results show that, although using correlation computation methods that treat zeros the same as non-zero signals could inflate correlations, ignoring these zeros, as in cosine similarity, can lead to worse performance.
+The primary implication of this experiment is that both strategies have inherent drawbacks when dealing with the sparsity of scRNA-seq data.
+Ignoring zeros misses out on significant information, while treating them as homogeneous signals inflates correlations and introduces noise.
+Ignoring the shared zeros in cell expression profiles, as cosine similarity does, is not an effective strategy and results in the worst performance.
+Conversely, methods that consider all the shared zeros in the expression profiles as an indication of cell homogeneity —such as Pearson’s, Spearman’s, and Kendall’s Tau—neglect the fact that many shared zeros are merely a result of the high dimensionality of scRNA-seq data.
+When sequencing results from multiple cells are merged, many zero entries are introduced since different cells often express different sets of genes.
+Consequently, cell correlations are inflated by these methods when data is sparse.
+The clustering outcomes from these methods are suboptimal when compared with approaches that do not require correlation computation, such as CoMem, DIPHW, and PCA, potentially due to misleading signals in the inflated correlation networks S1 4 5 6.
+Our results suggest that alternative approaches, such as hypergraph-based methods, provide a better solution by circumventing the need for unipartite network projections and avoiding the intricacy associated with handling zeros in correlation computations.
+.4 Embedding Visualizations
 We present the visualizations and clustering performance of embeddings generated by DIPHW, CoMem, and CoMem-DIPHW in Fig.
@@
 Additional Results for Human Brain Dataset
-Figure S7: CoMem-DIPHW Clustering Performance Assessment on the Human Brain Dataset [57] and Cell Type Annotation using DEGs and the PanglaoDB Marker Database.
+Figure S7: Clustering Performance of CoMem-DIPHW on the Human Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers.
 (a) Top 10 DEGs identified for each cluster by CoMem-DIPHW.
@@
 Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
-Figure S8: PCA Clustering Performance Assessment on the Human Brain Dataset [57] and Cell Type Annotation using DEGs and the PanglaoDB Marker Database.
+Figure S8: Clustering Performance of PCA on the Human Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers.
 (a) Top 10 DEGs identified for each cluster by PCA.
@@
 Additional Results for Mouse Pancreas and Mouse Brain Datasets
-Figure S9: CoMem-DIPHW Clustering Performance Assessment on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and the PanglaoDB Marker Database.
+Figure S9: Clustering Performance of CoMem-DIPHW on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and Canonical Markers.
 (a) Top 10 DEGs identified for each cluster by CoMem-DIPHW.
@@
 Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
-Figure S10: PCA Clustering Performance Assessment on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and the PanglaoDB Marker Database.
+Figure S10: Clustering Performance of PCA on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and Canonical Markers.
 (a) Top 10 DEGs identified for each cluster by PCA.
@@
 Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
-Figure S11: CoMem-DIPHW Clustering Performance Assessment on the Mouse Brain Dataset [59] and Cell Type Annotation using DEGs and the PanglaoDB Marker Database.
+Figure S11: Clustering Performance of CoMem-DIPHW on the Mouse Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers.
 (a) Top 10 DEGs identified for each cluster by CoMem-DIPHW.
@@
 Violin plots show the distribution of average expression levels of these DEGs across all clusters.
+Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
 (c) Cell type annotation using the PanglaoDB marker database.
-Figure S12: PCA Clustering Performance Assessment on the Mouse Brain Dataset [59] and Cell Type Annotation using DEGs and the PanglaoDB Marker Database.
+Figure S12: Clustering Performance of PCA on the Mouse Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers.
 (a) Top 10 DEGs identified for each cluster by PCA.

### v4

@@
-Hypergraph Representations of scRNA-seq Data for Improved Clustering with Random Walks
+Hypergraph Representations of Single-Cell RNA Sequencing Data for Improved Cell Clustering
 Abstract
-Analysis of single-cell RNA sequencing data is often conducted through considering their network projections such as co-expression networks, primarily due to the abundant availability of network analysis tools for downstream tasks.
-However, this approach has several limitations.
-First, projecting scRNA-seq data into networks can result in the loss of higher-order information since it summarizes the gene expression profiles across cells into pairwise cell or gene similarities.
-Secondly, scRNA-seq data is often very sparse, but the constructed co-expression network projections are fully connected without sparsification, resulting in dense and inefficient data representation.
-Further, the sparsity of scRNA-seq data leads to co-occurrence of zero expression in a large proportion of the genes.
-Such zero-expression co-occurence will be accounted as similarity in gene expression among cells by correlation methods such as Spearman’s rank and Pearson’s correlation coefficient and lead to an overestimation of co-expression.
-To address these limitations, we propose conceptualizing scRNA-seq expression data as hypergraphs, which are generalized graphs in which the hyperedges can connect more than two vertices.
-This flexibility makes hypergraphs particularly well-suited for representing scRNA-seq data without requiring further transformation such as co-expression network generation.
-In the context of scRNA-seq data, the hypergraph nodes represent cells and the edges represent genes.
-Specifically, each hyperedge connects to all the cells where the gene is actively expressed and records the expression of the gene across different cells.
-The hypergraph conceptualization enables us to explore multi-way relationships beyond the pairwise interactions in co-expression networks for scRNA-seq data analysis while utilizing network analysis tools.
-We propose two novel hypergraph random walk-based cell clustering methods: (1) the Dual-Importance Preference Hypergraph Walk (DIPHW) and (2) the co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk (CoMem-DIPHW).
-These methods outperform widely used and established methods in terms of clustering accuracy on both simulated and real scRNA-seq datasets.
-The improvement is especially significant when data modularity is weak.
-By conceptualizing the scRNA-seq data as hypergraphs, our methods provide a more nuanced representation of gene expression data, allowing the incorporation of network perspectives without the information loss caused by enforcing a unipartite co-expression network projection.
-Furthermore, CoMem-DIPHW incorporates the gene co-expression network, cell co-expression network, and the cell-gene expression hypergraph from the single-cell abundance counts data altogether for embedding computation.
-This approach accounts for both the local information from single-cell level gene expression and the global information from pairwise similarities in the two co-expression networks.
+Abstract
+Motivation:
+Single-cell RNA sequencing (scRNA-seq) data analysis is often performed using network projections that produce co-expression networks.
+These network-based algorithms are attractive because regulatory interactions are fundamentally network-based and there are many tools available for downstream analysis.
+However, most network-based approaches have two major limitations.
+First, they are typically unipartite and therefore fail to capture higher-order information.
+Second, scRNA-seq data are often sparse, so most algorithms for constructing unipartite network projections are inefficient and may overestimate co-expression relationships, or may under-utilize the sparsity when clustering (e.g., with cosine distance).
+To address these limitations, we propose representing scRNA-seq expression data as hypergraphs, which are generalized graphs where a hyperedge can connect more than two nodes.
+In this context, hypergraph nodes represent cells, and hyperedges represent genes.
+Each hyperedge connects all cells in which its corresponding gene is actively expressed, indicating the expression of that gene across different cells.
+The resulting hypergraph can capture higher-order information and appropriately handle varying levels of data sparsity.
+This representation enables clustering algorithms to leverage higher-order relationships for improved cell-type differentiation.
+Results: To distinguish cell types using hypergraph representations of scRNA-seq data, we introduce two novel clustering algorithms: (1) Dual-Importance Preference Hypergraph Walk (DIPHW) and (2) Co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk (CoMem-DIPHW).
+DIPHW is a new hypergraph-based random walk algorithm that computes cell embeddings by considering the relative importance of genes to cells and cells to genes, incorporating a preference exponent to facilitate clustering.
+CoMem-DIPHW integrates two unipartite projections, the gene co-expression and cell co-expression networks, along with the cell-gene expression hypergraph derived from single-cell abundance count data into the random walk model.
+The advantage of CoMem-DIPHW is that it accounts for both local information from single-cell gene expression and global information from pairwise similarity in the two co-expression networks.
+We benchmark the performance of our algorithms against established and state-of-the-art deep learning approaches using both real-world and simulated scRNA-seq data.
+Real-world datasets include cells from the human pancreas, mouse pancreas, human brain, and mouse brain tissues.
+We also use a ground-truth labeled cell-type annotation dataset based on human lung adenocarcinoma cell lines.
+Quantitative evaluation shows that CoMem-DIPHW consistently outperforms established algorithms and state-of-the-art deep learning algorithms for cell-type clustering.
+Our proposed algorithms show the greatest improvement on scRNA-seq data with weak modularity.
+Moreover, CoMem-DIPHW successfully annotates clusters with biologically relevant cell types.
+Our results highlight the utility of hypergraph representations in the analysis of scRNA-seq data.
+Availability and Implementation: Our methods are implemented in Python and are available at https://github.com/wanhe13/CoMem-DIPHW.
 I Introduction
-Network analysis has become a popular tool for studying complex systems in biology, as a reflection of the combinatorial interaction mechanism of biomolecules.
-Its applications include functional analysis  [1, 2, 3, 4] and interaction prediction in protein-protein interaction (PPI) networks [5, 6, 7], regulator identification and pathway analysis in gene regulatory networks [8, 9, 10], disease modeling, prediction, and intervention in epidemiology [11, 12, 13], and cell type identification in co-expression networks [14, 15, 16, 17].
-However, traditional methods represent the biology systems using unipartite networks, which only captures relationships (edges) among one type of entity (nodes), are often not the most natural or information-preserving representation of the complex systems since many involve interactions between multiple types of entities.
-Instead of modelling these systems using network models that could capture higher order interactions, most often, a unipartite network projection is chosen such that the system could be analyzed with the abundant network analysis methods developed based on unipartite networks [23, 14].
-For example, in protein-protein interaction (PPI) networks, unipartite representations capture relationships between proteins but fail to capture the biological functions associated with groups of proteins.
+Network analysis has become a popular tool for studying complex systems in biology, reflecting the combinatorial interactions of biomolecules.
+Its applications include functional analysis [szklarczyk2015string, veres2015comppi, vazquez2003global, ashtiani2018systematic] and prediction of interactions in protein-protein interaction (PPI) networks [kovacs2019network, lei2013novel, murakami2017network], identification of regulators and analysis of pathways in gene regulatory networks [coleman2023gene, muzio2021biological, zickenrott2016prediction], disease modeling, prediction, and intervention in epidemiology [kraemer2020effect, chinazzi2020effect, davis2021cryptic], and identification of cell types in co-expression networks [langfelder2008wgcna, satija2015spatial, butler2018integrating, wolf2018scanpy].
+However, traditional methods represent biological systems using unipartite networks, which capture relationships only between a single type of entity.
+Such representations are often not the most natural or information-preserving choice, since many complex systems involve interactions between multiple types of entities.
+Most often, a unipartite network projection is chosen such that the system could be analyzed with the abundant network analysis methods developed based on unipartite networks [horvat2012one, langfelder2008wgcna].
+However, this focus on unipartite networks restricts our attention to biological interactions just between pairs of genes, proteins, or other kinds of nodes.
+[^1] We use the terms node and vertex interchangeably.
+In reality, biological processes often entail multiway interactions that cannot be represented by unipartite networks.
+For instance, three-gene epistatic interactions vastly outnumber pairwise epistasis in experimental analyses of yeast genetic networks [kuzmin2018systematic].
+In protein-protein interaction (PPI) networks, unipartite representations capture relationships between proteins but fail to capture the biological functions associated with groups of proteins.
 Similarly, unipartite projections of gene regulatory networks connect genes based on regulatory relationships but do not capture relationships among different entities such as transcription factors or their binding sites.
-Single-cell RNA sequencing (scRNA-seq) has enabled the profiling of gene expression at the individual cell level [24, 25, 26, 27], whereas conventional bulk tissue RNA sequencing measures expression at the tissue level, averaging gene expression over an ensemble of cells.
-Cell type identification, as one of the most important downstream tasks in single-cell RNA-seq (scRNA-seq) data analysis  [28], has applications across biology and medicine, including tracing the trajectories of distinct cell lineages in development for cell differentiation studies  [29], tissue heterogeneity analysis for cancer research  [30, 31, 32], immune cell profiling for therapy development  [33, 34, 35, 36], and biomarker discovery for diagnosis.
+Therefore, network analysis methods are needed to more accurately capture the higher-order interactions that are apparently ubiquitous in biological systems.
+Single-cell RNA sequencing (scRNA-seq) has enabled the profiling of gene expression at the individual cell level [yanai2019cel, macosko2015highly, zheng2017massively, picelli2014full], whereas conventional bulk tissue RNA sequencing measures expression at the tissue level, averaging gene expression over an ensemble of cells.
+Cell type identification, as one of the most important downstream tasks in single-cell RNA-seq (scRNA-seq) data analysis [hwang2018single], has applications in biology and medicine, including tracing the trajectories of different cell lineages in the development of cell differentiation studies [marioni2017single], tissue heterogeneity analysis for cancer research [wong2022single, laviron2022tumor, karaayvaz2018unravelling], immune cell profiling for therapy development [landhuis2018single, lyons2017immune, ermann2015immune, chung2017single], and biomarker discovery for diagnosis.
 […]
-Given the issues with unipartite projections, it is crucial to explore alternative network representations that can accurately reflect the scRNA-seq data.
-Hypergraphs offer a neat solution by representing multiway relationships in scRNA-seq directly, without the need for further data projection.
-Figure 1: Network Representations of the scRNA-seq Data.
-The scRNA-seq cell-gene expression data can be represented by various networks, including unipartite network projections, such as cell and gene co-expression networks.
-In a cell co-expression network, the nodes are cells and the edge weights represent similarity in cell-pair expression.
-Similarly, in gene co-expression networks, nodes represent genes.
-Additionally, hypergraph and bipartite network representations preserve the exact expression level of a gene in a cell.
+Given the issues with unipartite projections, it is crucial to explore alternative network representations that can accurately reflect the scRNA-seq data and the underlying biological reality of higher-order interactions among genes.
+Hypergraphs [bianconi2021higher, battiston2020networks, battiston2021physics] offer a nice solution by directly representing multiway relationships in scRNA-seq data, without requiring further data projection.
+Existing work on hypergraph representations has demonstrated the improved performance they can achieve in modeling complex biological systems by incorporating higher-order interactions.
+For example, Liu et al.
+[liu2022multi] developed a hypergraph-based neural network to predict synergistic drug combinations for cancer treatment.
+Wang et al.
+[wang2024hypertmo] integrated multi-omics data with hypergraph convolutional neural networks to classify patients with diseases such as breast cancer and Alzheimer’s.
+Gaudelet et al.
+[gaudelet2018higher] used hypergraphs to model interactions across different levels of protein organization (such as protein-protein interactions, complexes, and pathways) to better capture the complexity of biological systems and predict biological functions.
+Ma and Ma [ma2022hypergraph] developed a hypergraph-based logistic matrix factorization method that predicts possible metabolite–disease interactions and uncovers novel disease-related metabolites.
+These successes highlight the potential of hypergraphs to model multiway relationships inherent in biological systems, motivating their application to scRNA-seq analysis.
+Figure 1: Network Representations of scRNA-seq Data.
+Expression data can be represented by various networks, including unipartite network projections such as cell and gene co-expression networks.
+In a cell co-expression network, nodes are cells and edge weights represent similarity in cell-pair expression.
+In gene co-expression networks, nodes represent genes.
+Hypergraph and bipartite network representations preserve the exact expression level of a gene in a cell.
 In the hypergraph representation, nodes represent cells and hyperedges represent genes.
-In the bipartite representation, one set of nodes represents cells, and the other set represents genes.
-Figure 2: CoMem Random Walk Illustration.
-In the hypergraph representation of the scRNA-seq data, nodes represent cells and edges represent genes.
+In the bipartite representation, one set of nodes represents cells and the other set represents genes.
+Figure 2: Hypergraph Clustering Algorithm Illustration.
+In the hypergraph representation of scRNA-seq data, nodes represent cells and edges represent genes.
 The expression level of a gene in a cell is captured by the weight of a node in an edge.
-Random walks are performed on the hypergraph by alternately considering the probability of the walker choosing a node given an edge, and choosing an edge given the current node.
+Random walks are performed on the hypergraph by alternately considering the probability of the walker choosing a node given an edge and choosing an edge given the current node.
 For DIPHW, the input is the hypergraph representation of the scRNA-seq data.
-For CoMem-DIPHW, the input includes the hypergraph, cell co-expression network, and gene co-expression network, which serve as node and edge similarity matrices to aid memory incorporation during the walk.
+For CoMem-DIPHW, the input includes the hypergraph, cell co-expression network, and gene co-expression network, which serve as node and edge similarity matrices to incorporate memory during the walk.
 A small neural network (word2vec) is applied to the sampled node paths from the hypergraph random walk process to compute cell embeddings.
-Finally, cell clustering is achieved by applying K-means to the obtained cell embeddings.
+Finally, cell clustering is performed by applying K-means to the resulting cell embeddings.
 II Methods
-We conceptualize scRNA-seq data as a hypergraph $\mathcal{H}$ and design two hypergraph random walk algorithms that capture cell-cell relationships by sampling the cell-to-gene and gene-to-cell interactions in the conceptualized hypergraph.
-We emphasize that this hypergraph representation of the scRNA-seq data does not require further transformation of the data and consequently would not result in any information loss or additional computational cost.
+We conceptualize scRNA-seq data as a hypergraph $\mathcal{H}$ and design two hypergraph random walk algorithms that capture cell-cell relationships by sampling cell-to-gene and gene-to-cell interactions in the hypergraph.
+This hypergraph representation of scRNA-seq data does not require additional data transformation, does not cause information loss, or increase computational cost.
 II.1 Hypergraph representation
-The single-cell RNA sequencing (scRNA-seq) data can be conceptualized as a hypergraph $\mathcal{H}=(V,E)$, where:
-  • $V$ represents the vertex set corresponding to the cells profiled in the experiment.
-  • $E$ denotes the hyperedge set representing genes, with each hyperedge $e\in E$ connecting to all cells $\in V$ in which the gene $e$ is actively expressed.
+scRNA-seq data can be conceptualized as a hypergraph $\mathcal{H}=(V,E)$, where:
+  • $V$ represents the set of nodes corresponding to the cells profiled in the experiment.
+  • $E$ denotes the set of hyperedges representing genes, with each hyperedge $e\in E$ connecting to all cells $\in V$ in which the gene $e$ is actively expressed.
 The weight of connection of a hyperedge (gene) to each node (cell) is determined by the gene’s expression level in that cell.
@@
 $\mathbf{I}_{\mathcal{H}}$ Incidence matrix of the hypergraph $\mathcal{H}$
-$V$ Set of all nodes (cells)
-$E$ Set of all hyperedges (genes) in the hypergraph
-$G_{V}$ Cell co-expression network/ Node similarity matrix
-$G_{E}$ Gene co-expression network/ Edge similarity matrix
-$\Gamma_{e}(v)$ Weight of node $v$ in hyperedge $e$, i.e., expression of gene $e$ in cell $v$
+$V$ Set of nodes (cells)
+$E$ Set of hyperedges (genes) in the hypergraph
+$G_{V}$ Cell co-expression network (node similarity matrix)
+$G_{E}$ Gene co-expression network (edge similarity matrix)
+$\Gamma_{e}(v)$ Weight of node $v$ in hyperedge $e$ (i.e., expression of gene $e$ in cell $v$)
 $P(V=v\mid E=e)$ Probability of selecting node $v$ given hyperedge $e$
@@
 $P(E_{t+1}=e^{\prime}\mid E_{t}=e,V_{t}=v)$ Probability of transitioning to hyperedge $e^{\prime}$ at time $t+1$, given current hyperedge $e$ and node $v$ at time $t$
-$P(V_{t+1}=w\mid E_{t+1}=e,V_{t}=v)$ Probability of selecting node $w$ at time $t+1$, given hyperedge $e$ and previous node $v$
+$P(V_{t+1}=w\mid E_{t+1}=e,V_{t}=v)$ Probability of selecting node $w$ at time $t+1$, given hyperedge $e$ at time $t+1$ and previous node $v$
 $P(V_{t+1}=w\mid V_{t}=v)$ Unipartite node-to-node transition probability
-Table 1: Hypergraph Random Walk Notation Table.
+Table 1: Notation Table for Our Hypergraph Random Walk Representation.
 II.2 Method 1: Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW)
-Building on this hypergraph representation, we introduced a novel hypergraph random walk approach for cell embedding computation, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW), which accounts for both the relative importance of edges to nodes and nodes to edges, or importance of genes relative to cells and cell to genes, along with a preference exponent for faster clustering.
+Building on this hypergraph representation, we introduce a novel hypergraph random walk approach for cell embedding computation, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW), which accounts for both the relative importance of edges to nodes and nodes to edges, or the importance of genes relative to cells and cells to genes, along with a preference exponent for faster clustering.
 In a hypergraph random walk process, the walker alternates between transitioning from a hyperedge to a node and from a node to a hyperedge, following the node-to-edge transition probabilities $P_{E|V}(e,v)$ and edge-to-node transition probabilities $P_{V|E}(v,e)$.
-In the hypergraph random walk with edge-dependent vertex weight (EDVW) by Chitra and Raphael (2019) [41], the edge-to-node transition probability is edge-dependent, but the node-to-edge transition probability does not consider the importance of the edge to the node.
-As a natural extension to EDVW, we defined the node-to-edge transition probability to be vertex-dependent in DIPHW, and used a preference exponent to accelerate the clustering process.
-The modified random walk process is described as follows:
+In the hypergraph random walk with edge-dependent vertex weight (EDVW) by Chitra and Raphael (2019) [Chitra2019], the edge-to-node transition probability is edge-dependent, but the node-to-edge transition probability does not consider the importance of the edge to the node.
+As a natural extension to EDVW, we define the node-to-edge transition probability to be vertex-dependent in DIPHW and use a preference exponent to accelerate the clustering process.
+The modified random walk process is as follows:
   1. 1.
-    Node-to-Hyperedge Transition Probability:
+    Node-to-Hyperedge Transition Probability
     The probability of selecting hyperedge $e_{t+1}$ given that the walker is currently at vertex $v_{t}$ is defined as:
@@
   2. 2.
-    Hyperedge-to-Node Transition Probability:
+    Hyperedge-to-Node Transition Probability
     Upon the walker’s arrival at hyperedge $e_{t+1}$, the probability of moving to vertex $v_{t+1}\in e_{t+1}$ is given by:
@@
   3. 3.
-    Node-to-Node Transition Probability:
-    The unipartite node-to-node transition probability of DIPHW is obtained by summing the product of the node-to-edge and edge-to-node transition probabilities over all hyperedges.
-This probability is used to sample random walk node paths, which serve as inputs to a neural network for cell embedding computation.
+    Node-to-Node Transition Probability
+    The unipartite node-to-node transition probability of DIPHW is calculated by summing the products of the node-to-edge and edge-to-node transition probabilities across all hyperedges.
+This probability is used to sample random walk node paths, which are then input into a neural network to compute cell embeddings.
     $P(v_{t}\rightarrow v_{t+1})=\sum_{e\in E}P_{E|V}(e|v_{t})P_{V|E}(v_{t+1}|e).$
@@
   4. 4.
-    Node-to-Node Transition Probability in Vectorized Format:
-    The $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ could be written in the vectorized form as,
+    Node-to-Node Transition Probability in Vectorized Format
+    The $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ could be written in the matrix form as
     $P=D_{E|V}^{-1}W_{E|V}D_{V|E}^{-1}W_{V|E}$
-    Vectorization eliminates explicit iteration and reduces the computation time.
-Here, $W_{E|V}=\mathbf{I}_{\mathcal{H}}^{T}W_{E}$ is the $|V|\times|E|$ node-to-edge transition weight matrix where column $e$ of $W_{E|V}$ contains the vertex weights in hyperedge $e$ multiplied by the weight of hyperedge $e$.
+    The matrix formulation eliminates explicit iteration and reduces computation time.
+Here, $W_{E|V}=\mathbf{I}_{\mathcal{H}}^{T}W_{E}$ is the $|V|\times|E|$ node-to-edge transition weight matrix, where column $e$ of $W_{E|V}$ contains the vertex weights in hyperedge $e$ multiplied by the weight of hyperedge $e$.
 Similarly, $W_{V|E}$ represents the edge-to-node transition weight matrix before normalization with dimensions $|E|\times|V|$, given by $W_{V|E}=\mathbf{I}_{\mathcal{H}}^{\alpha},\quad w_{ve}=\left[\mathbf{I}_{\mathcal{H}}(e,v)\right]^{\alpha}$.
@@
 II.3 Method 2: co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk (CoMem-DIPHW)
-We now introduce CoMem-DIPHW, an extension of DIPHW that incorporates co-expression networks and single-cell transcriptomic profiles.
+CoMem-DIPHW is an extension of DIPHW.
+It incorporates co-expression networks and single-cell transcriptomic profiles.
 The memory mechanism of CoMem-DIPHW leverages cell and gene co-expression networks to account for previously visited nodes and edges during transitions to capture both local expression relationships and global co-expression patterns.
+  1. 1.
     Hypergraph Representation of the scRNA-seq data
-We remark that a matrix can be directly interpreted as the incidence matrix of a hypergraph to represent the relationship between two variables, without additional computation.
+    We note that a matrix can be directly interpreted as the incidence matrix of a hypergraph to represent the relationship between two variables, without additional computation.
+  2. 2.
     The Node and Edge Similarity Networks: Cell and Gene co-expression Networks and their construction
-Using scRNA-seq data, we construct two co-expression networks: the cell-co-expression network (CCN) $G_{V}$ and the gene-co-expression network (GCN) $G_{E}$.
-In the CCN, cells are nodes, and edges correspond to correlation coefficients between cell pairs, indicating the similarity between their expression profiles.
-Similarly, in the GCN, genes are nodes and edges correspond to correlation coefficients between gene pairs.
-Spearman’s rank coefficient was chosen to measure pairwise similarity of expression profiles as it captures non-linearity better than Pearson’s correlation.
+    Using scRNA-seq data, we construct two co-expression networks: the cell co-expression network ($G_{V}$) and the gene co-expression network ($G_{E}$).
+In the cell co-expression network, cells are nodes, and edges represent correlation coefficients between cell pairs, indicating the similarity of their expression profiles.
+Similarly, in the gene co-expression network, genes are nodes, and edges represent correlation coefficients between gene pairs.
+We use Spearman’s rank coefficient to measure pairwise similarity of expression profiles, as it captures non-linearity better than Pearson’s correlation.
 These unipartite networks serve as node and edge similarity matrices, embedding global-level information that summarizes associations between cell and gene pairs.
+  3. 3.
     Memory-Integrated Random Walk
-CoMem-DIPHW integrates a memory component by considering the similarity between consecutively visited nodes and edges.
-This memory mechanism adjusts the transition probabilities such that the node-to-edge transition probability depends on the previous edge visited, and the edge-to-node transition probability depends on the previous node visited.
-The node-node and edge-edge dependencies are incorporated through the node and edge similarity networks, which are the two correlation networks derived from the transcriptomic expression data.
-The memory mechanism constrains the hypergraph walker to focus on cell-gene interactions among similar nodes(cells) and edges(genes) to better uncover modularity within the data.
-While the co-expression networks provide global perspectives through unipartite projections, specific expression levels of genes in individual cells (local information) are preserved in the hypergraph structure.
-CoMem-DIPHW integrates both global and local information for more accurate cell type identification.
+    CoMem-DIPHW incorporates a memory component by considering the similarity between consecutively visited nodes and edges.
+This memory mechanism adjusts the transition probabilities so that the node-to-edge transition probability depends on the previously visited edge, and the edge-to-node transition probability depends on the previously visited node.
+Node-node and edge-edge dependencies are incorporated through the node and edge similarity networks, which are two correlation networks derived from transcriptomic expression data.
+The memory mechanism constrains the hypergraph walker to focus on cell-gene interactions among similar nodes (cells) and edges (genes), thus better revealing modularity within the data.
+    While the co-expression networks provide global perspectives through unipartite projections, the specific expression levels of genes in individual cells (local information) are preserved in the hypergraph structure.
+CoMem-DIPHW integrates global and local information for more accurate cell type identification.
+  4. 4.
     Formulation
-The random walk process in CoMem-DIPHW is formulated as follows:
-Let $G_{V}$ be the vertex similarity matrix, where $G_{V}(v,w)$ quantifies the similarity between vertices $v$ and $w$.
+The random walk process in CoMem-DIPHW is formulated as follows.
+Let $G_{V}$ be the vertex similarity matrix, where $G_{V}(v,w)$ quantifies the similarity between nodes $v$ and $w$.
 Similarly, let $G_{E}$ be the edge similarity matrix, where each element $G_{E}(e,f)$ represents the similarity between edge $e$ and $f$.
 Method 2.0: CoMem
+In CoMem, we calculate the unipartite node-to-node transition probability of the hypergraph random walker. To do this, we first calculate the following four probabilities.
   1. 1.
@@
 To encourage exploration and ensure non-lazy walks, we modify the node and edge similarity graphs such that the diagonal elements are set to zero.
-This ensures that during the random walk, the random walker does not remain at the same edge or node in consecutive steps.
+This ensures that during the random walk, the random walker does not remain at the same node or edge in consecutive steps.
 Method 2.1: CoMem based on DIPHW
@@
 Specifically, the vertex selection probability $P_{V|E}(e\rightarrow v)$ and the edge selection probability $P_{E|V}(v\rightarrow e)$ are both defined as in DIPHW (section II.2), while all other probability distributions remain as specified in section II.3.
+II.4 Cell Embedding and Clustering
+We perform random walks using the defined hypergraph node-to-edge and edge-to-node transition probabilities to sample sequences of visited cells.
+These sampled cell paths, which capture the geometry of the hypergraph representing the scRNA-seq data, are then used as input to word2vec [mikolov2013distributed] to learn low-dimensional vector representations of cells.
+Cells with similar expression profiles frequently co-occur in the sampled random walk paths and are closer in the embedded space.
+Finally, K-means clustering is applied to cell embeddings to obtain cluster assignments.
+II.5 scRNA-seq Data Simulation
+In our study, we develop an algorithm to simulate data that mimics the characteristics of scRNA-seq expression data (see Algorithm 1).
+In addition to the underlying assumption of within-cell-type homogeneous expression that governs most cell clustering methods, we incorporate between-cell-type crosstalk [armingol2021deciphering, zhang2023defining, zepp2019cellular, bayik2021cancer, lehuen2010immune] by integrating configurable intermodular covariance, density, and signal strength into the simulation model.
+The algorithm generates a sparse matrix representing the simulated scRNA-seq expression data, with numerous user-configurable parameters, including number of modules, module density, shape, background signal strength, modular signal strength, intermodular signal strength, intermodular covariance, and noise level (see Table S4).
+This method is implemented as a function in our codebase and is available for use by other researchers [wan2025comem].
+Our simulation model is highly flexible to accommodate various types of scRNA-seq data.
+For example, the model can simulate both sequenced datasets with high density (up to 50%) and sequenced datasets with low density (as low as 1%) [andrews2021tutorial].
+The proportion of non-housekeeping genes can be adjusted by changing the number of differentially expressed genes (DEGs) in each module.
+In addition, the variance in module sizes and the number of co-expressed genes in each cell type can be configured by setting different variances for the Poisson distribution used in the embedded module simulation.
+The modules embedded in the simulated data can be configured with higher density and average signal strength to model the expression profiles of biomarker genes that are highly expressed in certain cell types.
+This enables us to mimic the sparsity and modularity of real scRNA-seq data, where a small number of genes are highly expressed in specific cell types.
+Alternatively, the embedding modules can be user-specified without using the embedded module simulation.
+Algorithm 1 Simulation Algorithm for scRNA-seq Data
+1:Input: Number of genes $E$, number of cells $V$, number of cell types $K$; average genes per cell type $\bar{g}$ and average cells per cell type $\bar{c}$; average expression levels for within-cell-type $\lambda_{\text{ct}}$, cross-cell-type $\lambda_{\text{cross}}$, and background $\lambda_{\text{bg}}$; densities within-cell-type $\rho_{\text{ct}}$, cross-cell-type $\rho_{\text{cross}}$, and background $\rho_{\text{bg}}$; and crosstalk probability $p_{\text{cross}}$.
+2:Output: Sparse expression matrix $\mathbf{X}\in\mathbb{R}^{E\times V}$ with ground-truth labels.
+3:Initialize $\mathbf{X}\leftarrow\mathbf{0}_{E\times V}$.
+4:Sample cell type sizes: $g_{k}\sim\text{Poisson}(\bar{g})$, $c_{k}\sim\text{Poisson}(\bar{c})$ for $k=1,\ldots,K$.
+5:Phase 1: Within-Cell-Type Signal
+6:for each cell type $k=1,\ldots,K$ do
+7:  For each gene $e\in E_{k}$ and cell $v\in V_{k}$: with probability $\rho_{\text{ct}}$, set $\mathbf{X}[e,v]\leftarrow\text{Poisson}(\lambda_{\text{ct}})$.
+8:end for
+9:Phase 2: Background Noise
+10:Randomly fill $\rho_{\text{bg}}$ fraction of remaining zero entries with $\text{Poisson}(\lambda_{\text{bg}})$ values.
+11:Phase 3: Cross-Cell-Type Crosstalk
+12:for each cell type pair $(k,l)$ where $k\neq l$ do
+13:  With probability $p_{\text{cross}}$: randomly fill $\rho_{\text{cross}}$ fraction of $\mathbf{X}[E_{k},V_{l}]$ with $\text{Poisson}(\lambda_{\text{cross}})$ values.
+14:end for
+15:return $\mathbf{X}$, ground-truth labels.
 III Experiments and Results
-III.1 Baseline Methods
-To evaluate the performance of DIPHW and CoMem-DIPHW, we compared them to several widely used and established cell clustering methods, including five embedding-based methods (PCA [43], Node2Vec [44], EDVW [41], UMAP, and t-SNE) and four unipartite-graph-based community detection methods (Louvain, Greedy Modularity, Infomap, and Leiden).
-These methods are adopted by popular scRNA-seq analysis packages such as Seurat [15], WGCNA [14], and Scanpy [17].
-We provide a summary of these methods in Table 2.
-For a fair comparison under consistent conditions, we focus on the core clustering algorithms rather than specific implementations in packages such as Seurat and Scanpy.
-This allows us to evaluate the clustering capabilities of each method independently of any additional preprocessing or optimization steps.
-The hyperparameters used for each experiment are detailed in Table S1 in the supplementary materials Supplementary Materials.
-Method Type Data Input Packages
-Greedy
-Modularity [45]
-Community Detection Unipartite network
-(Cell co-expression network)
-igraph [46]
-Louvain [47] Community Detection Unipartite network
-(Cell co-expression network)
-Seurat [15], Scanpy [17], WGCNA [14]
-Infomap [48] Community Detection Unipartite network
-(Cell co-expression network)
-igraph [46]
-Leiden [49] Community Detection Unipartite network
-(Cell co-expression network)
-igraph [46]
-PCA [43] Embedding Matrix (Genes by cells) Seurat [15], Scanpy [17], SC3 [50], Cell Ranger [26]
-t-SNE [51] Embedding Matrix (Genes by cells) Seurat [15], Scanpy [17]
-Node2Vec [44] Embedding Unipartite network
-(Cell co-expression network)
-Multiple implementations available [52, 53] .
-UMAP [54] Embedding Matrix (Genes by cells) Seurat [15], Scanpy [17]
-EDVW [41] +
-word2vec [55]
-Embedding Hypergraph incidence matrix
-(Genes by cells)
-Implemented based on Chitra and Raphael [41]; available in our GitHub repository [56]. Not in standard scRNA-seq packages.
-DIPHW +
-word2vec [55]
-Embedding Hypergraph incidence matrix
-(Genes by cells)
-Our proposed method and implementation is available on GitHub [56]
-CoMem +
-word2vec [55]
-Embedding Hypergraph incidence matrix
-(Genes by cells);
-Gene co-expression network;
-Cell co-expression network
-Our proposed method and implementation is available on GitHub [56]
-Table 2: Comparison of Clustering Methods Evaluated in Experiments. For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
+III.1 Baseline and Competing Methods
+To evaluate the performance of DIPHW and CoMem-DIPHW, we compared them with 13 cell-clustering methods from various categories, including community detection, embedding-based methods, and recent deep learning methods that cluster scRNA-seq data.
+Table 2 summarizes these methods.
+For a fair comparison under consistent conditions, we focus on core clustering algorithms rather than specific implementations in packages such as Seurat [satija2015spatial], Scanpy [wolf2018scanpy], and WGCNA [langfelder2008wgcna].
+This approach allows us to evaluate the clustering capabilities of each method independently of any additional preprocessing or optimization steps.
+The hyperparameters used for each experiment are detailed in Table S4 in the supplementary materials.
+Method Type Input Packages
+Greedy Modularity [newman2004fast] Community detection Cell co-expression network igraph [csardi2006igraph]
+Louvain [blondel2008fast] Community detection Cell co-expression network Seurat [satija2015spatial], Scanpy [wolf2018scanpy], WGCNA [langfelder2008wgcna]
+Infomap [rosvall2008maps] Community detection Cell co-expression network igraph [csardi2006igraph]
+Leiden [traag2019louvain] Community detection Cell co-expression network igraph [csardi2006igraph]
+PCA [pearson1901liii] Linear embedding Gene $\times$ cell matrix Seurat [satija2015spatial], Scanpy [wolf2018scanpy], SC3 [kiselev2017sc3], Cell Ranger [zheng2017massively]
+t-SNE [van2008visualizing] Manifold embedding Gene $\times$ cell matrix Seurat [satija2015spatial], Scanpy [wolf2018scanpy]
+UMAP [mcinnes2018umap] Manifold embedding Gene $\times$ cell matrix Seurat [satija2015spatial], Scanpy [wolf2018scanpy]
+Node2Vec [grover2016Node2Vec] Graph embedding Cell co-expression network Multiple implementations available [grover_Node2Vec_2016, snap_Node2Vec_nodate]
+graph-sc [ciortan2022gnn] Deep graph embedding Cell co-expression network GitHub [graphsc_github]
+tsImpute [zheng2023tsimpute] Imputation Gene $\times$ cell matrix GitHub [tsimpute_github]
+CAKE [liu2024cake] Deep contrastive clustering Gene $\times$ cell matrix GitHub [cake_github]
+scASDC [min2024scasdc] Deep clustering Gene $\times$ cell matrix GitHub [scasdc_github]
+EDVW [Chitra2019] + word2vec [mikolov2013efficient] Hypergraph embedding Gene $\times$ cell matrix Implemented based on [Chitra2019]; GitHub [wan2025comem]
+DIPHW + word2vec [mikolov2013efficient] Hypergraph embedding Gene $\times$ cell matrix Our proposed method; GitHub [wan2025comem]
+CoMem + word2vec [mikolov2013efficient] Hypergraph embedding Gene $\times$ cell matrix, plus gene and cell co-expression networks Our proposed method; GitHub [wan2025comem]
+Table 2: List of Clustering Methods Evaluated in Our Study.
+K-means was used to cluster the output of all embedding-based methods that do not include a clustering algorithm.
+Since tsImpute is an imputation method, we evaluated its impact on clustering performance by applying PCA followed by K-means to its imputed data.
 III.2 Data and Preprocessing
-The data used in this study consists of both simulated and publicly available scRNA-seq datasets.
-The simulated data were generated using our proposed simulation algorithm described in detail in the Methods section, with code available on GitHub [56].
-For preprocessing, we first normalized the raw single-cell RNA-seq data to account for differences in sequencing depth.
-Genes and cells with zero expression were then removed, and only the top n most variable genes were retained for analysis.
-For empirical analysis, we used four publicly available scRNA-seq datasets: human brain [57], human pancreas [58], mouse brain [59], and mouse pancreas [60], and a curated benchmarking dataset with ground-truth cell line labels [61].
-III.3 Sparsity-Induced Correlation Inflation in scRNA-seq Co-expression Networks
-The zero inflation problem in single-cell RNA sequencing (scRNA-seq) data analysis refers to the excessive number of zero counts (or "dropouts") observed in the data.
-These zeros could either be a result of technical limitation in detecting low abundance transcripts (false negatives), or true absence of biological zeros (true negatives) [37, 39, 40].
-The process of integrating individual cell expression profiles into a unified scRNA-seq dataset requires constructing a matrix that accounts for every gene detected in any cell, which artificially introduces zero entries wherever genes are not expressed in those cells.
-Since cells typically express different sets of genes, this inclusion introduces significant sparsity into the merged dataset.
-In co-expression network construction, edge weights are inflated due to these artificially introduced common zeros since the common zeros will be considered as evidence of their expression homogeneity.
-This issue is particularly pronounced at the single-cell level compared to bulk tissue RNAseq data, as the proportion of expressed genes at the single-cell level is much smaller, resulting in higher induced data sparsity.
-In the experiment shown in Fig.
-3, we investigate how induced sparsity affects correlation coefficients and the interpretation of gene and cell connections in co-expression networks.
-To simulate this effect, we generated gene expression data with a fixed base size and introduced varying proportions of zeros to mimic different sparsity levels.
-Our results show that as the level of induced sparsity increases, correlation coefficients between expression profiles become artificially inflated, leading to an overestimation of similarity between cell expression profiles.
-These findings highlight the importance of addressing zero inflation in scRNA-seq data analysis and the need for accurate representation of the underlying biological relationships.
-While induced sparsity inflates correlations, our supplementary experiments (.3) show that ignoring shared zeros using cosine similarity leads to worse clustering performance, as these inactive genes still hold biological significance.
-[…]
-III.4 Simulation Analysis Results
-In our study, we developed an algorithm to simulate data that mimics the characteristics of scRNA-seq expression data.
-In addition to the underlying assumption of within-cell-type homogeneous expression that governs most, if not all, cell clustering methods, we also incorporated between-cell-type crosstalk [62, 63, 64, 65, 66] by integrating configurable intermodular covariance, density, and signal strength into the simulation model.
-The algorithm generates a sparse matrix representing the simulated scRNA-seq expression data, with numerous user-configurable parameters, including number of modules, module density, shape, background signal strength, modular signal strength, intermodular signal strength, intermodular covariance, and noise level.
-This method is implemented as a function in our codebase and is available for use by other researchers.
-Our simulation model is designed with high flexibility to accommodate different types of scRNA-seq data.
-For instance, the model can simulate both deeply sequenced datasets with high density (up to 50%) and shallowly sequenced datasets with low density (as low as 1%) [67].
-The proportion of non-housekeeping genes can be adjusted by changing the number of differentially expressed genes in each module.
-Moreover, the variance in module sizes and the number of co-expressed genes in each cell type can be configured by setting different variances for the Poisson distribution used in the embedded module simulation.
-The modules embedded in the simulated data can be configured with higher density and average signal strength to model the expression profiles of biomarker genes that are highly expressed in certain cell types.
-This allows us to mimic the sparsity and modularity of real scRNA-seq data, where a small number of genes are highly expressed in specific cell types.
-Alternatively, the embedding modules can also be user-specified without using the embedded module simulation.
-III.4.1 Evaluation Metrics
-To evaluate the performance of the clustering methods on datasets with known ground-truth, we employed two widely used metrics for measuring coherence between the identified clusters and the ground-truth clusters: Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI).
-Both metrics are invariant to permutations, with a score of 0 representing random labeling and a score of 1 indicating identical clusters.
-For embedding-based methods that do not directly assign cluster membership, we used K-means to partition the embedded data into clusters.
-[…]
-III.4.2 Clustering Performance vs. Bipartite Modularity and Method Comparison
-We evaluated the performance of these clustering methods under varying data modularity conditions.
-The modularity variation was controlled by two key parameters: the average number of co-expressed genes per cell type and the number of embedded cell modules.
-The number of modules here represents the number of cell types in the simulated scRNA-seq data.
-Weak modularity corresponds to a small number of co-expressed genes per module or a large number of embedded modules, making data structures more challenging to detect.
+We use simulated and publicly available scRNA-seq datasets for our study.
+The simulated data is generated using our simulation algorithm described in Section II.5, with pseudocode in Algorithm 1 and code at [wan2025comem].
+For both simulated and real scRNA-seq data, preprocessing includes Counts Per Million normalization [vallejos2017normalizing] to account for differences in sequencing depth, log transformation to reduce skewness, and removal of genes and cells with zero total expression.
+For real scRNA-seq datasets, we additionally retain the top $n$ highly variable genes using Scanpy’s implementation of the Seurat [satija2015spatial] method.
+The values of $n$ used for each dataset are provided in Table S4.
+We use six publicly available scRNA-seq datasets: (1) human brain [camp2015human], (2) human pancreas [muraro2016single], (3) mouse brain [zeisel2015cell], (4) mouse pancreas [baron2016single], and (5) scMixology benchmark with 3 classes and (6) scMixology benchmark with 5 classes.
+scMixology benchmark which is a curated benchmark dataset with ground-truth cell line labels [tian2019benchmarking].
+Table 2 provides a summary of these datasets.
+Datasets  # of Cells  # of Genes  Sparsity Results in
+scMixology Benchmark (3-class) [tian2019benchmarking] 902 16,468 45.02% Table 4
+scMixology Benchmark (5-class) [tian2019benchmarking] 3,918 11,786 63.01% Table 4
+Human Pancreas  [muraro2016single] 3,072 18,348 78.74% Section III.6
+Human Brain  [camp2015human] 735 18,929 80.11% Section S7
+Mouse Brain  [zeisel2015cell] 3,006 19,973 81.21% Section S8
+Mouse Pancreas  [baron2016single] 1,065 14,881 87.79% Section S8
+Table 3: List of Publicly Available Datasets Used in Our Study. Sparsity measures the percentage of zero entries in a cell $\times$ gene matrix.
+III.3 Sparsity-Induced Inflated Correlation in scRNA-seq Co-expression Networks
+The zero inflation problem in scRNA-seq data analysis refers to the excessive number of zero counts (or "dropouts") observed in the data.
+These zeros may arise from technical limitations in detecting low-abundance transcripts (false negatives) or from the true absence of transcripts (true negatives) [jiang2022statistics, qiu2020embracing, kim2020demystifying].
+Integrating individual cell expression profiles into a unified scRNA-seq dataset requires constructing a matrix that includes every gene detected in any cell, which introduces zero entries wherever genes are not expressed in those cells.
+Because cells typically express different sets of genes, this process creates significant sparsity in the merged dataset.
+In co-expression network construction, edge weights are inflated due to these artificially introduced common zeros, as the common zeros are interpreted as evidence of expression homogeneity.
+This issue is especially pronounced at the single-cell level compared to bulk tissue RNA-seq data, since the proportion of expressed genes per cell is much smaller, resulting in greater data sparsity.
+In Fig.
+3, we examine how induced sparsity affects correlation coefficients and the interpretation of gene and cell connections in co-expression networks.
+To simulate this effect, we generate gene expression data with a fixed base size and introduce varying proportions of zeros to represent different sparsity levels.
+Our results indicate that as induced sparsity increases, correlation coefficients between expression profiles become artificially inflated, leading to an overestimation of similarity between cell expression profiles.
+These findings highlight the importance of addressing zero inflation in scRNA-seq data analysis and the need for accurate representation of underlying biological relationships.
+Table S1 (in the Supplementary Section S2) shows that ignoring shared zeros using cosine similarity results in poorer clustering performance, as these inactive genes still have biological significance.
+Figure 3: Impact of Increased Sparsity on Correlation Coefficients.
+This figure shows the relationship between the proportion of zeros added to the expression profile (representing increased sparsity) and the resulting correlation coefficients.
+This process was repeated 100 times.
+Shaded areas indicate the 95% confidence intervals around the average correlation values.
+The dashed lines represent the average initial correlation without added sparsity.
+Induced sparsity in scRNA-seq data can inflate correlations in co-expression networks, highlighting the need for alternative scRNA-seq data representations.
+III.4 Results on the Simulated Data
+To evaluate performance of the clustering methods on datasets with known ground-truth, we employ five widely used measures to compute coherence between the identified clusters and the ground-truth clusters: Adjusted Rand Index (ARI), Normalized Mutual Information (NMI), Adjusted Mutual Information (AMI), Clustering Accuracy (ACC), and F1 Score.
+All evaluation measures are invariant to permutations, with a score of 0 representing random labeling and a score of 1 indicating identical clusters.
+K-means is used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
+Figure 4: Comparing Performances of Clustering Methods on Simulated scRNA-seq Data Across Modularity.
+The performance of clustering methods is measured by ARI, NMI, ACC, AMI, and F1 across different levels of data modularity.
+The x-axis represents parameters determining data modularity, including the average number of co-expressed genes per cell type and the number of embedded modules.
+Our proposed methods, DIPHW, CoMem, and CoMem-DIPHW, are highlighted with dashed lines.
+CoMem-DIPHW, in particular, maintains strong performance and is least affected by decreases in modularity, performing significantly better even under conditions of extremely weak modularity.
+This behavior is observed consistently across all evaluation measures.
+Experiments are repeated 10 times for each parameter setting, and the 95% confidence intervals (CIs) are shown.
+The bar plots represent Barber’s bipartite modularity, normalized so that the most modular graphs have modularity = 1.
+K-means is used to cluster the output of all embedding-based methods that do not include an inherent clustering mechanism.
+We evaluate the performance of clustering methods under varying data modularity conditions.
+Modularity is controlled by two key parameters: the average number of co-expressed genes per cell type and the number of embedded cell modules.
+The number of modules corresponds to the number of cell types in the simulated scRNA-seq data.
+Weak modularity refers to a small number of co-expressed genes per module or a large number of embedded modules, which makes data structures more difficult to detect.
 The modularity bar plots in Fig.
-4 and heatmap visualization in Fig.
-S6 demonstrate that the simulated scRNA-seq data displays stronger modularity when the average number of co-expressed genes per cell type is greater or when the number of modules is smaller.
-We implemented Barber’s bipartite modularity [42], as shown by the barplots in Fig.
-4, to help quantify modularity of the underlying data for clustering and analyze the association between modularity and clustering performance.
-Clustering performance generally improves with increased data modularity.
-Our proposed methods—DIPHW, CoMem, and CoMem-DIPHW—consistently outperformed all other clustering methods that use unipartite network projections, including Node2Vec, and community detection-based methods such as Louvain and Greedy Modularity.
-Notably, CoMem-DIPHW consistently demonstrated superior performance across almost all parameter settings, with its advantage being particularly pronounced under weak data modularity conditions.
-For clarity of presentation, we present the most contrasting cases from our analysis: (i) an average of 10 and 100 co-expressed genes per module in Fig.
+4 and the heatmap visualization in Fig.
+S11 show that the simulated scRNA-seq data exhibit stronger modularity when the average number of co-expressed genes per cell type is higher or when the number of modules is smaller.
+We implemented Barber’s bipartite modularity [barber2007modularity], as shown in the bar plots in Fig.
+4, to quantify the modularity of the underlying data for clustering and to analyze the association between modularity and clustering performance.
+Clustering performance generally improves as data modularity increases.
+Our proposed methods, DIPHW, CoMem, and CoMem-DIPHW, demonstrate strong and highly competitive performance across varying modularity.
+Specifically, CoMem-DIPHW consistently remains competitive with the best-performing methods evaluated, with its advantage particularly pronounced under weak data modularity conditions.
+For clarity, we present the most contrasting cases from our analysis: (i) an average of 10 and 100 co-expressed genes per module in Fig.
 5 and (ii) 10 and 100 embedded modules in Fig.
 6.
-[…]
-[…]
-In Fig.
-5 and Fig.
-6, we compare the performance of our proposed hypergraph-based clustering methods (highlighted in red and yellow) against five embedding-based methods and four community detection methods under both strong and weak modularity conditions.
-All embedding-based methods are coupled with K-means to produce clustering results.
-In weak modularity regimes (where the average number of co-expressed genes = 10 or when number of modules = 100), CoMem-DIPHW and DIPHW exhibited a significant performance advantage—for example, when the average number of co-expressed genes = 10, CoMem-DIPHW had an ARI of 0.29 compared to 0.05 for the best-performing baseline method (PCA), representing a performance improvement of approximately 450%.
-Under strong modularity conditions (when the average number of genes per module = 100 or when number of modules = 10), several methods—including PCA, CoMem-DIPHW, CoMem, Node2Vec, and DIPHW—performed well, achieving average ARI scores above 0.9.
-Additional results from intermediate parameter settings (average number of co-expressed genes per module: 20, 30, 40, 70, 80, 90; number of embedded modules: 20, 30, 40, 70, 80, 90) and by NMI support the same observation.
-Out of the 16 tested scenarios, CoMem-DIPHW or DIPHW consistently ranked first by ARI and in 10 scenarios by NMI.
-Our hypergraph-based methods showed significant advantages in weak modularity regimes, where data structures are harder to detect.
-Under comparatively strong modularity conditions, several methods (PCA, CoMem-DIPHW, CoMem, Node2Vec, DIPHW, and Louvain) performed well, with average NMI scores above 0.9.
-Notably, PCA slightly outperformed CoMem-DIPHW by NMI under strong modularity, with differences in NMI scores ranging from 0.01 to 0.02.
-These additional results are presented in Supplementary Figs.
-S2, Fig.
-S3, Fig.
-S4, Fig.
-S5.
-III.5 scMixology Benchmark Results
-A major challenge in scRNA-seq analysis is the lack of reliable ground-truth for cell type annotation, which complicates the evaluation of clustering algorithms.
-In most scRNA-seq datasets, cell type labels are typically assigned using a combination of clustering results, differential expression (DE) analysis, and canonical marker genes.
-First, cells are grouped into clusters based on transcriptomic similarity using clustering algorithms (e.g., Leiden, Louvain, or k-means).
-Then, DE analysis is performed to identify genes that are differentially expressed across clusters.
+Figure 5: Impact of Module Size on Clustering Performance by ARI.
+The bar plots show the mean ARI (top) and average ranking (bottom) for each method, ordered by their average ARI.
+Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval.
+Red dashed lines indicate the highest ARI values or best ARI rankings.
+Our proposed hypergraph-based methods, highlighted in red and yellow, show significant advantages when data modularity is weak, that is, when the average number of co-expressed genes per cell type module is low (10 genes per module).
+Under strong modularity conditions, where the average number of genes per cell type module is high (100 genes per module), many methods (e.g., PCA, CoMem-DIPHW, CoMem, and Node2Vec) perform well and achieve average ARI scores above 0.9.
+These observations are consistent across NMI, ACC, AMI, and F1, as shown in Supplementary Section S6.
+K-means is used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
+Figure 6: Impact of Module Count on Clustering Performance by ARI.
+The bar plots show the mean ARI (top) and average ranking (bottom) of each method, ordered by their average ARI.
+Our proposed hypergraph-based methods, highlighted in red and yellow, demonstrate significant advantages in weak modularity regimes, where the number of embedded modules (representing the number of cell types) in the simulated data is high (100 modules).
+Each experiment is repeated 10 times per parameter setting, with error bars representing the 95% confidence interval.
+Red dashed lines indicate the highest ARI values or best ARI rankings.
+These observations are consistent across NMI, ACC, AMI, and F1, as shown in Supplementary Section S6.
+K-means is used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
+In Figures 5 and 6, we compare the performance of our proposed hypergraph-based clustering methods (highlighted in red and yellow) with five embedding-based approaches, four community detection algorithms, and four state-of-the-art scRNA-seq clustering methods.
+K-means is used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
+In weak modularity regimes (average number of co-expressed genes per module = 10 or number of modules = 100), CoMem-DIPHW and DIPHW show a clear performance advantage.
+For example, when the average number of co-expressed genes is 10, CoMem-DIPHW achieves an ARI of 0.30 and DIPHW achieves an ARI of 0.25, compared to 0.05 for the best baseline method (PCA), representing a performance improvement of nearly 450%.
+Under strong modularity conditions (average number of genes per module = 100 or number of modules = 10), multiple methods (PCA, CoMem-DIPHW, CoMem, and Node2Vec) performed well, achieving average ARI scores above 0.9.
+Additional results from intermediate parameter settings (average number of co-expressed genes per module: 20, 30, 40, 70, 80, 90; number of embedded modules: 20, 30, 40, 70, 80, 90), as measured by NMI, ACC, AMI, and F1, support the same observation.
+Across the 16 parameter settings under different modularity conditions, CoMem-DIPHW ranks first in all 16 settings by ACC and F1, 15 by ARI, and 10 each by NMI and AMI.
+DIPHW and CoMem also perform strongly.
+Either DIPHW or CoMem rank among the top three methods in at least 10 of 16 settings across all evaluation measures.
+Our hypergraph-based methods show significant advantages in weak modularity regimes, where data structures are harder to detect.
+Under comparatively strong modularity conditions, several methods (PCA, CoMem-DIPHW, CoMem, Node2Vec, DIPHW, and Leiden) perform well, with average NMI scores above 0.9.
+Notably, PCA slightly outperforms CoMem-DIPHW by ARI, NMI, and AMI under strong modularity, with differences in scores ranging from 0.01 to 0.02.
+These additional results are in Supplementary Section S6.
+III.5 Results on the scMixology Benchmark Datasets
+A major challenge in scRNA-seq analysis is the lack of reliable ground truth for cell type annotation, which complicates the evaluation of clustering algorithms.
+In most scRNA-seq datasets, cell type labels are assigned using a combination of clustering results, differential expression (DE) analysis, and canonical marker genes.
+First, cells are grouped into clusters based on transcriptomic similarity using clustering algorithms such as K-means.
+Next, DE analysis identifies genes that are differentially expressed across clusters.
 Finally, these clusters are annotated based on known canonical marker genes associated with specific cell types.
-Since this annotation process relies on clustering results, using these labels to evaluate clustering performance introduces circular logic and bias toward the method originally used for annotation.
-To address this issue, we utilized the mixture control benchmarking dataset [61], where ground truth is established based on cell line identity.
-We evaluated the performance of our proposed method, CoMem-DIPHW, using two benchmarking datasets from scMixology [61].
-These datasets provide ground-truth labels for cell type annotation based on human lung adenocarcinoma cell lines.
-As shown in Table 3, CoMem-DIPHW outperformed other methods, particularly on the 5-class dataset, achieving the highest NMI and ARI scores.
-In comparison, Leiden showed strong performance on the 3-class dataset but underperformed significantly on the 5-class dataset.
-Among the methods evaluated, PCA, CoMem, Node2Vec and CoMem-DIPHW consistently demonstrated robust performance across both datasets.
-Dataset Greedy Modularity Louvain Infomap Leiden PCA t-SNE Node2Vec UMAP EDVW DIPHW CoMem CoMem-DIPHW
-scMixology ARI 0.5836 0.1976 0.2327 0.9933 0.9900 0.9898 0.9900 0.9898 0.6619 0.9868 0.9933 0.9900
-scMixology NMI 0.7193 0.3818 0.3843 0.9864 0.9797 0.9810 0.9797 0.9810 0.6609 0.9743 0.9864 0.9797
-scMixology 5 class ARI 0.6645 0.3374 0.2998 0.6746 0.9873 0.9867 0.9461 0.9838 0.6873 0.9373 0.9880 0.9923
-scMixology 5 class NMI 0.7852 0.4498 0.4355 0.7917 0.9747 0.9748 0.9148 0.9707 0.7740 0.9001 0.9770 0.9849
-Table 3: Comparison of ARI and NMI across Clustering Methods for the ScMixology [61] Benchmark Datasets.
-The table is divided into 5-class (top) and 3-class (bottom) results.
-Best values are in bold, second-best are underlined.
-While PCA also performs well on these simple clustering tasks, CoMem and CoMem-DIPHW show robust performance across both datasets and are expected to outperform on more complex, weakly modular data, based on the previous simulation results.
-III.6 Human Pancreas Results
-For datasets without ground-truth annotations—human pancreas, mouse pancreas, human brain, and mouse brain—we conducted differential expression (DE) analysis based on clustering results.
-We present our differential expression analysis results from the human pancreas dataset by CoMem-DIPHW and PCA in detail and include a brief overview of results from the other datasets in the Supplementary Materials Section  Supplementary Materials.
-We report the top 10 differentially expressed genes (DEGs) from each human pancreas cell cluster identified by CoMem-DIPHW and PCA in Fig.
-7a and Fig.
-8a respectively.
-To evaluate clustering performance, we examined the expression of these DEGs across clusters (Fig.
-7b) and PCA (Fig.
-8b).
-Violin plots show that the DEGs identified by CoMem-DIPHW exhibit more cluster-specific expression patterns than PCA-derived DEGs, which tend to show higher expression in multiple clusters.
+Because this annotation process relies on clustering results, using these labels to evaluate clustering performance introduces circular logic and bias toward the method originally used for annotation.
+To address this issue, we use the mixture control benchmarking dataset scMixology [tian2019benchmarking], where ground truth is established based on cell line identity.
+We evaluate the performance of our proposed method, CoMem-DIPHW, using two benchmarking datasets from scMixology [tian2019benchmarking].
+These datasets provide ground truth labels for cell type annotation based on human lung adenocarcinoma cell lines.
+As shown in Table 4, tsImpute achieved the highest scores across all five evaluation measures (ARI, NMI, AMI, ACC, and F1) on the 3-class dataset, with CoMem-DIPHW, Louvain, and Leiden tied for second.
+On the 5-class dataset, CoMem-DIPHW achieved the best performance across all evaluation measures, followed by tsImpute and graph-sc.
+This result is consistent with the pattern observed in our simulation analysis (see Section III.4), where CoMem-DIPHW maintains strong performance under low modularity conditions.
+Traditional Embedding and Community Detection Methods
+Dataset / Evaluation Measures GreedyModularity Louvain Infomap Leiden PCA t-SNE Node2Vec UMAP
+scMixology 3 class
+ARI $0.5836$ $\pm$ 0.000 $0.9933$ $\pm$ 0.000 $0.2327$ $\pm$ 0.000 $0.9933$ $\pm$ 0.000 $0.9900$ $\pm$ 0.000 $0.9898$ $\pm$ 0.000 $0.9913$ $\pm$ $0.0037$ $0.9898$ $\pm$ 0.000
+NMI $0.7193$ $\pm$ 0.000 $0.9864$ $\pm$ 0.000 $0.3843$ $\pm$ 0.000 $0.9864$ $\pm$ 0.000 $0.9797$ $\pm$ 0.000 $0.9810$ $\pm$ 0.000 $0.9829$ $\pm$ $0.0075$ $0.9810$ $\pm$ 0.000
+ACC $0.6929$ $\pm$ 0.000 $0.9978$ $\pm$ 0.000 $0.3370$ $\pm$ 0.000 $0.9978$ $\pm$ 0.000 $0.996$ $\pm$ 0.000 $0.9967$ $\pm$ 0.000 $0.9971$ $\pm$ $0.0013$ $0.9967$ $\pm$ 0.000
+AMI $0.7189$ $\pm$ 0.000 $0.9864$ $\pm$ 0.000 $0.2335$ $\pm$ 0.000 $0.9864$ $\pm$ 0.000 $0.9796$ $\pm$ 0.000 $0.9810$ $\pm$ 0.000 $0.9829$ $\pm$ $0.0075$ $0.9810$ $\pm$ 0.000
+F1 $0.5868$ $\pm$ 0.000 $0.9978$ $\pm$ 0.000 $0.4588$ $\pm$ 0.000 $0.9978$ $\pm$ 0.000 $0.9967$ $\pm$ 0.000 $0.9967$ $\pm$ 0.000 $0.9971$ $\pm$ $0.0013$ $0.9967$ $\pm$ 0.000
+scMixology 5 class
+ARI $0.6645$ $\pm$ 0.000 $0.6746$ $\pm$ 0.000 $0.3002$ $\pm$ 0.0008 $0.6746$ $\pm$ 0.000 $0.9873$ $\pm$ 0.000 $0.9861$ $\pm$ 0.000 $0.9451$ $\pm$ $0.0073$ $0.9838$ $\pm$ 0.000
+NMI $0.7852$ $\pm$ 0.000 $0.7953$ $\pm$ 0.000 $0.4355$ $\pm$ $0.0003$ $0.7917$ $\pm$ 0.000 $0.9747$ $\pm$ 0.000 $0.9739$ $\pm$ 0.000 $0.9150$ $\pm$ $0.0086$ $0.9707$ $\pm$ 0.000
+ACC $0.7325$ $\pm$ 0.000 $0.7361$ $\pm$ 0.000 $0.4150$ $\pm$ $0.0018$ $0.7358$ $\pm$ 0.000 $0.9941$ $\pm$ 0.000 $0.9939$ $\pm$ 0.000 $0.9621$ $\pm$ $0.0059$ $0.9929$ $\pm$ 0.000
+AMI $0.7851$ $\pm$ 0.000 $0.7951$ $\pm$ 0.000 $0.2735$ $\pm$ $0.0003$ $0.7916$ $\pm$ 0.000 $0.9747$ $\pm$ 0.000 $0.9738$ $\pm$ 0.000 $0.9149$ $\pm$ $0.0086$ $0.9707$ $\pm$ 0.000
+F1 $0.6556$ $\pm$ 0.000 $0.6576$ $\pm$ 0.000 $0.5250$ $\pm$ $0.0028$ $0.6571$ $\pm$ 0.000 $0.9941$ $\pm$ 0.000 $0.9939$ $\pm$ 0.000 $0.9618$ $\pm$ $0.0062$ $0.9929$ $\pm$ 0.000
+Advanced ScRNA-seq and Hypergraph-Based Methods
+Dataset / Evaluation Measure EDVW graph-sc tsImpute CAKE scASDC DIPHW CoMem CoMem-DIPHW
+scMixology 3 class
+ARI $0.5960$ $\pm$ $0.0454$ $0.7020$ $\pm$ $0.2650$ 0.997 $\pm$ 0.000 $0.8318$ $\pm$ $0.0196$ $0.4227$ $\pm$ $0.4342$ $0.9900$ $\pm$ $0.0052$ $0.9912$ $\pm$ $0.0019$ $0.9933$ $\pm$ 0.000
+NMI $0.6140$ $\pm$ $0.0266$ $0.7654$ $\pm$ $0.1883$ 0.993 $\pm$ 0.000 $0.8609$ $\pm$ $0.0205$ $0.4500$ $\pm$ $0.4397$ $0.9803$ $\pm$ $0.0097$ $0.9829$ $\pm$ $0.0033$ $0.9864$ $\pm$ 0.000
+ACC $0.7996$ $\pm$ $0.0557$ $0.8612$ $\pm$ $0.1275$ 0.999 $\pm$ 0.000 $0.9958$ $\pm$ $0.0011$ $0.6435$ $\pm$ $0.2956$ $0.9967$ $\pm$ $0.0018$ $0.9971$ $\pm$ $0.0006$ $0.9978$ $\pm$ 0.000
+AMI $0.6132$ $\pm$ $0.0266$ $0.7649$ $\pm$ $0.1887$ 0.993 $\pm$ 0.000 $0.8605$ $\pm$ $0.0206$ $0.4479$ $\pm$ $0.4418$ $0.9803$ $\pm$ $0.0097$ $0.9829$ $\pm$ $0.0033$ $0.9864$ $\pm$ 0.000
+F1 $0.7991$ $\pm$ $0.0561$ $0.8513$ $\pm$ $0.1380$ 0.999 $\pm$ 0.000 $0.9957$ $\pm$ $0.0011$ $0.5477$ $\pm$ $0.3810$ $0.9967$ $\pm$ $0.0018$ $0.9971$ $\pm$ $0.0006$ $0.9978$ $\pm$ 0.000
+scMixology 5 class
+ARI $0.6572$ $\pm$ $0.0221$ $0.9884$ $\pm$ $0.0015$ $0.9892$ $\pm$ 0.000 $0.7333$ $\pm$ $0.0746$ $0.3919$ $\pm$ $0.3602$ $0.9440$ $\pm$ $0.0031$ $0.9879$ $\pm$ $0.0013$ 0.990 $\pm$ 0.002
+NMI $0.7421$ $\pm$ $0.0197$ $0.9782$ $\pm$ $0.0019$ $0.9796$ $\pm$ 0.000 $0.8435$ $\pm$ $0.0217$ $0.4575$ $\pm$ $0.4160$ $0.9092$ $\pm$ $0.0043$ $0.9771$ $\pm$ $0.0026$ 0.980 $\pm$ 0.003
+ACC $0.6985$ $\pm$ $0.0161$ $0.9950$ $\pm$ $0.0006$ $0.9954$ $\pm$ 0.000 $0.8568$ $\pm$ $0.0192$ $0.5553$ $\pm$ $0.2156$ $0.9628$ $\pm$ $0.0020$ $0.9946$ $\pm$ $0.0007$ 0.996 $\pm$ 0.001
+AMI $0.7418$ $\pm$ $0.0197$ $0.9782$ $\pm$ $0.0019$ $0.9796$ $\pm$ 0.000 $0.8431$ $\pm$ $0.0218$ $0.4565$ $\pm$ $0.4169$ $0.9091$ $\pm$ $0.0043$ $0.9771$ $\pm$ $0.0026$ 0.980 $\pm$ 0.003
+F1 $0.7113$ $\pm$ $0.0143$ $0.9950$ $\pm$ $0.0006$ $0.9949$ $\pm$ $0.000$ $0.8579$ $\pm$ $0.0175$ $0.3443$ $\pm$ $0.2246$ $0.9628$ $\pm$ $0.0018$ $0.9946$ $\pm$ $0.0007$ 0.996 $\pm$ 0.001
+Table 4: Performance of Clustering Methods on the scMixology Benchmark Datasets [tian2019benchmarking] across Five Evaluation Measures (ARI, NMI, ACC, AMI, and F1).
+The scMixology benchmark datasets include a 3-class and a 5-class dataset.
+Results for each are shown in the top and bottom sections, respectively.
+Best values are in bold, and second best are underlined.
+Results are reported as mean $\pm$ standard deviation over five independent runs.
+On the 3-class dataset, tsImpute achieved the highest scores across all evaluation measures, with CoMem-DIPHW, Louvain, and Leiden tied for second.
+On the more challenging 5-class dataset, CoMem-DIPHW achieved the best performance across all evaluation measures, outperforming tsImpute, which ranked second.
+III.6 Results on the Human Pancreas Dataset
+For datasets without ground-truth annotations (specifically, human pancreas, mouse pancreas, human brain, and mouse brain), we qualitatively evaluate clustering performance using differential expression (DE) analysis.
+We present detailed DE analysis results for the human pancreas dataset using CoMem-DIPHW, PCA, graph-sc, tsImpute, CAKE, and scASDC in this section and provide results for the human brain, mouse brain, and mouse pancreas in the Supplementary Sections S7 and S8.
+We report the top 10 differentially expressed genes (DEGs) from each human pancreas cell cluster identified by CoMem-DIPHW (Fig.
+7a), PCA (Fig.
+8a), graph-sc (Fig.
+9a), tsImpute (Fig.
+10a), CAKE (Fig.
+11a), and scASDC (Fig.
+12a).
+To evaluate cluster purity, we examine the expression of these DEGs across clusters (Fig.
+7b, Fig.
+8b, Fig.
+9b, Fig.
+10b, Fig.
+11b, Fig.
+12b).
+An effective clustering method should yield DEGs that uniquely characterize each cluster while showing minimal expression in other clusters.
+Violin plots show that the DEGs identified by CoMem-DIPHW exhibit more cluster-specific expression patterns than those derived from PCA, which tend to show higher expression in multiple clusters.
 For example, the average expression of the DEGs from PCA Cluster 1 is higher in Clusters 2, 5, and 6.
 Similarly, the DEGs from Cluster 7 show higher average expression in Cluster 6.
-In general, the average expression of the DEGs from the PCA clusters is less cluster-specific than the DEGs from the CoMem-DIPHW clusters.
-To tentatively label the cell clusters, we used canonical gene markers reported in the original human pancreas scRNA-seq study [58].
-We assigned clusters based on the expression of these canonical markers including GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells).
-Each marker is distinctly expressed in a well-defined cluster in the embedding generated by CoMem-DIPHW, as shown in Fig.
+Among the comparison methods, CAKE and scASDC produced highly distinctive DEGs, with scASDC yielding clusters that showed the strongest intra-cluster cohesion and inter-cluster separability.
+Graph-sc and PCA show less cluster-specific profiles, with DEGs frequently expressed across multiple clusters.
+To assess the biological relevance of the clusters, we cross-reference the cluster-specific DEGs (see Fig.
+7a) with markers from the PanglaoDB cell type marker database (version 27 March 2020)[franzen2019panglaodb].
+Cell-type assignment is based on the overlap between the DEGs of each cluster and the cell type-specific markers in the PanglaoDB marker database, with a match score calculated as the proportion of matched markers (Table 7c).
+CoMem-DIPHW (Fig.
+7c) successfully annotates all nine clusters with biologically relevant pancreatic cell types.
+CAKE produces similarly meaningful annotations, while graph-sc has several clusters with no annotation or biologically implausible labels.
+To further validate the clustering results, we examine the expression of canonical pancreatic marker genes [muraro2016single], including GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells) in the embeddings (Fig.
+7d, Fig.
+8d, Fig.
+9d, Fig.
+10d, Fig.
+11d, Fig.
+12d).
+Each marker is distinctly expressed in the embedding generated by CoMem-DIPHW, as shown in Fig.
 7d, confirming that CoMem-DIPHW embeddings effectively distinguish different cell types.
-In contrast, PCA embeddings lacked clear separation between different cell types (Fig.
-8d).
-For example, in Fig.
-7d, cluster 0 was identified as acinar cells by the high expression of PRSS1.
-Clusters 2 and 7 were identified as alpha cells by the high expression of GCG.
-Cluster 3 was classified as ductal cells by the high expression of KRT19.
-Both COL1A1 and ESAM are highly expressed in cluster 4, suggesting mesenchymal or endothelial cells.
-Cluster 5 was classified as delta cells by the high expression of SST, while cluster 6 was classified as beta cells by the high expression of INS.
-The specific DEGs supporting each classification are detailed in Fig.
-7a.
-To further validate the cell type assignments and clustering results, we cross-referenced the cluster-specific DEGs (see Fig.7a) with markers from the PanglaoDB cell type marker database (version 27 Mar 2020)[74].
-The assignment of cell types was determined based on the overlap between the DEGs of each cluster and the cell type-specific markers in the PanglaoDB marker database, with a match score computed based on the proportion of matched markers (Table 7c).
-These results align well with our previous analysis based on marker gene distribution (Figure 7d).
-The only discrepancy lies in cluster 4.
+CAKE and scASDC also produce well-separated embeddings.
+In contrast, PCA embeddings show poor separation.
+The results from CoMem-DIPHW align well with the cell-type annotation by the PanglaoDB database.
+The only discrepancy is in Cluster 4.
 Matching the DEGs with PanglaoDB cell type markers suggests Pancreatic stellate cells (PSCs) or Granulosa cells, while the COL1A1 and ESAM expression plots indicate mesenchymal or endothelial cells.
-However, this apparent discrepancy can be explained by the fact that PSCs can differentiate into cells with endothelial-like properties [75], and granulosa cells can exhibit mesenchymal-like behaviors [76].
-In contrast, PCA clustering results (Fig.8) showed limited cluster separation.
+However, this apparent discrepancy can be explained by the fact that PSCs can differentiate into cells with endothelial-like properties [cheng2021generation], and granulosa cells can exhibit mesenchymal-like behaviors [jozkowiak2020stemness].
+In contrast, PCA clustering results (Fig.
+8) showed limited cluster separation.
 While some PCA clusters could be annotated consistently using both PanglaoDB markers and canonical markers, significant overlap remained among clusters.
@@
 (a) Top 10 DEGs per cluster identified by CoMem-DIPHW.
-For example, Cluster 0 (acinar cells) includes genes associated with digestive enzyme production, a primary function of pancreatic acinar cells, including CPA1, CPA2 [68], PLA2G1B [69], CTRB1 [70], and PRSS1 [71].
-Clusters 2 and 7 (alpha cells) include markers GLS, CRYBA2, FAP, LOXL4, MAFB, and RGS4 (cluster 2) and LOXL4, PAX6, RGS4, and CRYBA2 (cluster 7).
+For example, Cluster 0 (acinar cells) includes genes associated with digestive enzyme production, a primary function of pancreatic acinar cells, including CPA1, CPA2 [tamura2018mutations], PLA2G1B [hui2019group], CTRB1 [pinho2011adult], and PRSS1 [masson2013conservative].
+Clusters 2 and 7 (alpha cells) include markers GLS, CRYBA2, FAP, LOXL4, MAFB, and RGS4 (cluster 2), and LOXL4, PAX6, RGS4, and CRYBA2 (cluster 7).
 Cluster 3 (ductal cells) includes SLC4A4, ANXA4, KRT19, CFTR, CLDN1, and ONECUT2.
@@
 Each subplot displays the expression of a canonical marker gene for a specific human pancreatic cell type.
-The cell-type-specific markers considered include: GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells).
+The cell-type-specific markers considered include GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells).
 Color intensity represents gene expression levels, with red hues indicating higher expression.
-Clusters show high expression of cell-type-specific canonical markers are assigned to the corresponding cell types.
-Overall, clustering quality is demonstrated through (b) high within-cluster DEG expression, (c) consistency between PanglaoDB-based annotation and clustering, and (d) distinct separation of cell types.
+Clusters showing high expression of cell-type-specific canonical markers are assigned to the corresponding cell types.
+Overall, clustering quality is demonstrated by (b) high within-cluster DEG expression, (c) consistency between PanglaoDB-based annotation and clustering, and (d) distinct separation of cell types.
 Figure 8: Clustering Performance of PCA on the Human Pancreas Dataset and Cell Type Annotation Using Differentially Expressed Genes (DEGs) and Canonical Markers.
 (a) Top 10 DEGs per cluster identified by PCA.
-(b) Across-cluster average expression of cluster-specific DEGs.
+(b) Average expression of cluster-specific DEGs across clusters.
 Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
@@
 (d) Expression of human pancreatic canonical markers in PCA embeddings.
-The cell-type-specific markers considered include: GCG (alpha cells), INS (beta cells), SST (delta cells), PPY (PP cells), PRSS1 (acinar cells), KRT19 (ductal cells), COL1A1 (mesenchymal cells), and ESAM (endothelial cells).
 UMAP visualizations show PCA cell embeddings, with cluster IDs assigned by K-means clustering.
-The overall clustering of cell types is limited compared to CoMem-DIPHW clustering, as evidenced by the overlapping beta cells, delta cells and PP cells in (d) and the lack of distinct high expression patterns of the DEGs in (b).
-In addition to the human pancreas results presented here, we also analyzed mouse pancreas, human brain, and mouse brain datasets.
-For these datasets, we have included visualizations in the appendix that show the average expression of the top DEGs across clusters identified by different methods, which shows our proposed algorithm’s ability in effectively distinguishing cell clusters by their expression profiles.
-The supplements also contain tables listing the top DEGs for clusters identified by CoMem-DIPHW and PCA in each dataset for reference.
+Overall, the clustering of cell types is limited compared to CoMem-DIPHW clustering, as shown by the overlapping beta cells, delta cells, and $P$ cells in (d) and the absence of distinct high expression patterns of the DEGs in (b).
+Figure 9: Clustering Performance of graph-sc on the Human Pancreas Dataset and Cell Type Annotation Using Differentially Expressed Genes (DEGs) and Canonical Markers.
+(a) Top 10 DEGs per cluster identified by graph-sc.
+(b) Across expression of cluster-specific DEGs across clusters.
+Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
+Graph-sc clusters generally show moderate specificity with substantial cross-cluster overlap.
+(c) Cell type annotation using the PanglaoDB marker database.
+Four clusters (1, 4, 7, and 8) yield no annotation, indicating that their DEGs did not match known cell type markers.
+(d) Expression of human pancreatic canonical markers in graph-sc embeddings.
+UMAP visualizations show graph-sc cell embeddings, with cluster IDs assigned by K-means clustering.
+Overall, cell cluster separation is limited.
+Figure 10: Clustering Performance of tsImpute on the Human Pancreas Dataset and Cell Type Annotation Using Differentially Expressed Genes (DEGs) and Canonical Markers.
+(a) Top 10 DEGs per cluster identified by tsImpute combined with PCA embeddings and K-means.
+(b) Average expression of cluster-specific DEGs across clusters.
+Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
+tsImpute DEGs show high specificity in Clusters 2, 4, 5, 6, 7, 8, and 9.
+(c) Cell type annotation using the PanglaoDB marker database.
+Several annotations are biologically implausible for pancreas tissue, including parietal cells (Cluster 1), salivary mucous cells (Cluster 3), gamma delta T cells (Cluster 5), and pulmonary vascular smooth muscle cells (Cluster 8).
+These unexpected annotations suggest that the imputation process may introduce artifacts that confound cell type identification.
+(d) Expression of human pancreatic canonical markers in tsImpute combined with PCA embeddings.
+UMAP visualizations show tsImpute plus PCA cell embeddings, with cluster IDs assigned by K-means clustering.
+tsImpute improves cluster separation compared to PCA without imputation.
+Figure 11: Clustering Performance of CAKE on the Human Pancreas Dataset and Cell Type Annotation Using Differentially Expressed Genes (DEGs) and Canonical Markers.
+(a) Top 10 DEGs per cluster identified by CAKE.
+(b) Average expression of cluster-specific DEGs across clusters.
+Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
+CAKE DEGs show good cluster specificity.
+(c) Cell type annotation using the PanglaoDB marker database.
+Most CAKE clusters correspond to biologically relevant cell types: gamma/$P$ cells (Clusters 0 and 3), ductal cells (Clusters 1 and 5), pancreatic stellate cells (Cluster 2), acinar cells (Cluster 4), alpha cells (Cluster 7), and beta cells (Cluster 8).
+(d) Expression of human pancreatic canonical markers in CAKE embeddings.
+UMAP visualizations show CAKE cell embeddings, with cluster IDs assigned by K-means clustering.
+Marker genes show distinct concentration in the well-separated CAKE embeddings: GCG in Cluster 7, INS in Cluster 8, SST in Clusters 0 and 3, PRSS1 in Cluster 4, and KRT19 in Cluster 1.
+Figure 12: Clustering Performance of scASDC on the Human Pancreas Dataset and Cell Type Annotation Using Differentially Expressed Genes (DEGs) and Canonical Markers.
+(a) Top 10 DEGs per cluster identified by scASDC.
+(b) Average expression of cluster-specific DEGs across clusters.
+Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
+scASDC DEGs show good cluster specificity.
+(c) Cell type annotation using the PanglaoDB marker database.
+While most clusters are matched to relevant pancreatic cell types, some are annotated with unexpected cell types, including retinal progenitor cells (Cluster 0) and nuocytes (Cluster 7).
+Clusters 1 and 3 received no annotation.
+(d) Expression of human pancreatic canonical markers in scASDC embeddings.
+scASDC embeddings show the best internal cohesion and separability.
+The marker genes show distinct concentration in the well-separated scASDC embeddings.
+However, the cell type annotations suggest that the embeddings may not necessarily correspond to biologically meaningful clusters.
+In addition to the human pancreas results presented here, we also analyze mouse pancreas, human brain, and mouse brain datasets.
+For these datasets, we include visualizations in the supplementary material in Figures S12b through S29b that show the average expression of the top DEGs across clusters identified by different methods.
+These results show our proposed algorithm’s ability in effectively distinguishing cell clusters by their expression profiles.
+The supplementary material (Figures S12a through S29a) also contain tables listing the top DEGs for clusters identified by CoMem-DIPHW, PCA, graph-sc, tsImpute, CAKE, and scASDC in each dataset for reference.
+To complement the qualitative biological validation presented above, we further assess clustering performance using quantitative cluster quality measures, including the Silhouette score, Calinski-Harabasz index, Davies-Bouldin index, and Coefficient of Variation.
+These evaluations are conducted in two settings: (1) directly in the learned embedding space and (2) by evaluating the same cluster assignments on the preprocessed expression data used as input to the embedding algorithms.
+We refer the reader to Supplementary Tables S2 and S3.
+Evaluation in the embedding space captures the geometric separability and cohesion of clusters, whereas evaluation on the preprocessed data assesses whether the resulting cluster assignments capture meaningful structure in the input gene expression data.
+As demonstrated by differential expression and cell-type annotation analyzes in the human pancreas data set, strong embedding-space cohesion and separation alone do not necessarily guarantee biological interpretability.
 IV Conclusion
-In this study, we developed and analyzed two hypergraph-based clustering algorithms, DIPHW and CoMem-DIPHW for improved cell clustering performance on scRNA-seq data.
-Firstly, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW) we developed extends the edge-dependent vertex weight (EDVW) hypergraph random walk proposed by [41], and consistently outperformed it in our experiments.
-DIPHW’s key idea lies in its symmetrized account of importance: it considers both the relative importance of edges to nodes and nodes to edges, or importance of genes relative to cells and cell to genes, along with a preference exponent for faster clustering.
-Secondly, the CoMem and CoMem-DIPHW clustering methods we developed incorporate a memory mechanism that considers both the abundance counts data (conceptualized as hypergraphs) and the two co-expression networks to integrate both local and global information when computing embeddings for clustering.
-Thirdly, we identified and showed with experiments 3 the issue of inflated signals in co-expression networks brought by sparsity of scRNA-seq data, which provides additional motivation for the use of hypergraph representations.
-Fourthly, we compared our proposed methods with well-established clustering methods commonly used in scRNA-seq analysis packages.
-Our clustering performance evaluation spanned a broad range of modularity conditions, where CoMem-DIPHW and DIPHW consistently exhibited superior performance, particularly in weak modularity regimes.
-When clusters were harder to detect, baseline methods experienced a much more pronounced decay in performance compared to CoMem-DIPHW.
-For empirical analysis, we tested our algorithms on 4 real scRNA-seq datasets and 2 benchmark datasets.
-In these experiments, CoMem-DIPHW consistently demonstrated superior performance in differentiating and classifying distinct cell types compared to the baseline methods.
-Additionally, we designed a versatile simulation algorithm for scRNA-seq data that is highly customizable through its many user-specifiable parameters.
-This algorithm is designed to model a wide range of expression scenarios and modular structures.
+In this study, we develop and analyze two hypergraph-based clustering algorithms, DIPHW and CoMem-DIPHW, to improve cell clustering performance on scRNA-seq data.
+First, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW) extends the edge-dependent vertex weight (EDVW) hypergraph random walk [Chitra2019] and consistently outperforms it in our experiments.
+The key idea of DIPHW is its symmetrized treatment of importance, considering both the relative importance of edges to nodes and nodes to edges, or the importance of genes relative to cells and cells to genes, along with a preference exponent for faster clustering.
+Second, the CoMem and CoMem-DIPHW clustering methods incorporate a memory mechanism that considers both the abundance counts data (conceptualized as hypergraphs) and the two co-expression networks to integrate local and global information when computing embeddings for clustering.
+Third, we identify and demonstrate through experiments (see Fig.
+3) the issue of inflated signals in co-expression networks caused by the sparsity of scRNA-seq data, which further motivates the use of hypergraph representations.
+Fourth, we compare our proposed methods with 13 cell-clustering methods from various categories, including community detection, embedding-based methods, and recent deep learning methods used in clustering scRNA-seq data.
+Our evaluation covers a broad range of modularity conditions in the data, where CoMem-DIPHW and DIPHW consistently show superior performance, particularly in weak modularity regimes.
+When clusters are harder to detect, other methods experience a much more pronounced decline in performance compared to CoMem-DIPHW.
+Fifth, we test our proposed methods on four real scRNA-seq datasets and two benchmark datasets.
+In these experiments, CoMem-DIPHW consistently demonstrates superior performance in differentiating and classifying distinct cell types compared to the other methods.
+Sixth, we design a versatile simulation algorithm for scRNA-seq data that is highly customizable through many user-specified parameters.
+This algorithm models a wide range of expression scenarios and modular structures.
 Notably, it incorporates intermodular co-expressions, which mimic cell-type cross-talk observed in real biological systems.
-For future work, we plan to optimize the implementation of CoMem-DIPHW to reduce its memory complexity, enhancing its scalability for larger datasets.
-We are also interested in exploring alternative ways to incorporate the memory mechanism into the hypergraph random walk.
-Currently, CoMem employs a simple product approach, moving beyond the simple product used in CoMem could potentially uncover even more effective clustering strategies.
+IV.1 Future Work
+We plan to optimize the implementation of CoMem-DIPHW to reduce its memory complexity, enabling it to scale to larger datasets.
+We will also explore alternative methods for incorporating the memory mechanism into the hypergraph random walk.
+Currently, CoMem uses a simple product approach; moving beyond this could reveal more effective clustering strategies.
+Although our current evaluation focuses on the biological relevance of clusters through cell-type annotation using PanglaoDB markers, future application-focused studies could include Gene Ontology enrichment analysis to characterize the functions of identified clusters.
+With the increasing availability of data and studies in spatial transcriptomics [li2024spadit, rao2021exploring, longo2021integrating, kleshchevnikov2022cell2location], extending the CoMem-DIPHW formulation for spatial transcriptomics analysis is a promising direction.
+While the current CoMem-DIPHW formulation focuses on cell clustering based on expression profile similarity, it could be extended to incorporate spatial proximity by replacing the cell expression similarity matrix $G_{V}$ with functions that characterize joint cell similarity based on both expression profiles and spatial information.
 Supplementary Materials
-Figure/Table Description Value
-Clustering results on simulated scRNA-seq data using embedding-based and community detection Methods (Figure S1) Number of cells 3000
-Number of genes 1000
-Number of embedded modules 30
-Average number of genes per module 30
-Average number of cells per module 38
-Target density of the gene expression matrix 0.03
-Density within modules 0.3
-Density between modules 0.1
-Probability of inter-module connections 0.6
-Average background expression 10
-Average within-module expression 20
-Average inter-module expression 10
-Workers 30
-Walk length 30
-Embedding dimensions 30
-Preference exponent 50
-Performance comparison of clustering methods on simulated scRNA-seq data modularity (Figure 4 for average genes per cell type plots) Average number of genes per module {10, 20, 30, 40, 70, 80, 90, 100}
-Other parameters same as Figure S1
-Performance comparison of clustering methods on simulated scRNA-seq data modularity (Figure 4 for number of modules plots) Number of embedded modules {10, 20, 30, 40, 70, 80, 90, 100}
-Average number of genes per module 50
-Average number of cells per module ( # of genes / # of modules) +3
-Other parameters same as Figure S1
-Clustering performance comparison by ARI across varying module sizes (Figure 5) (Same as in Figure 4 for average genes per cell type plots)
-Clustering performance comparison by ARI across varying module counts (Figure 6) (Same as in Figure 4 for number of modules)
-Comparison of ARI and NMI across clustering methods for the ScMixology benchmark datasets (Table 3) Number of highly expressed genes used 500
-Preference exponent 10
-Workers 50
-Walk length 10
-Embedding dimensions 10
-Human pancreas (Section III.6) Number of highly expressed genes used 1000
-Number of clusters 9
-Preference exponent 10
-Workers 100
-Walk length 50
-Embedding dimensions 10
-Mouse pancreas (Section Additional Results for Mouse Pancreas and Mouse Brain Datasets) Number of clusters 13
-Other parameters same as for human pancreas analysis (Section III.6)
-Human brain (Section Additional Results for Human Brain Dataset) Number of clusters 7
-Other parameters same as for human pancreas analysis (Section III.6)
-Mouse brain (Section Additional Results for Mouse Pancreas and Mouse Brain Datasets) Number of clusters 9
-Other parameters same as for human pancreas analysis (Section III.6)
-Table S1: Hyperparameters Used in Experiments.
-.1 Complexity Analysis
-DIPHW
-The computation cost of each element $P(u\rightarrow v)$ in the $|V|\times|V|$ node-to-node transition probability matrix is $O(|E|)$ when the intermediate matrices for the node-to-edge and edge-to-node transition probabilities $P_{E|V}(e|u)$ and $P_{V|E}(v|e)$ are precomputed.
-Constructing the entire $|V|\times|V|$ node-to-node transition probability matrix thus involves computing these probabilities for all node pairs, resulting in an overall time complexity of $O(|V|^{2}\cdot|E|)$.
-For efficient computation, the $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ in vectorized form is given by:
-$P_{V|V}=D_{E|V}^{-1}W_{E|V}D_{V|E}^{-1}W_{V|E}$
-While the theoretical complexity remains $O(|E|^{2}\cdot|V|+|E|\cdot|V|^{2})$ for the vectorized approach, it utilizes parallel computation and is significantly faster and hence more suitable for large datasets.
-The memory complexity is $O(|E|\cdot|V|)$ for the node-to-edge and edge-to-node transition probability matrices.
-CoMem-DIPHW
-The time complexity of CoMem-DIPHW is $O(|E|^{2}\cdot|V|^{2})$, as for each node pair, the memory-incorporated transition traverses all possible edges that could have arrived at the first node and all possible edges that could then connect the fisrt node with the second node.
-The memory complexity is $O(|E|^{2}\cdot|V|+|E|\cdot|V|^{2})$ for the memory incorporated node-to-edge and edge-to-node transition probability matrices.
-.2 Bipartite representation
-We chose the hypergraph conceptualization over bipartite representation for consistency since the DIPHW algorithm we introduced is an extension of the edge-dependent vertex weight hypergraph random walk (EDVW) [41].
-DIPHW extended EDVW hypergraph random walk [41] with a vertex-dependent edge selection probability, as well as a preference exponent for clustering acceleration.
-A bipartite conceptualization could be adopted for our proposed methods DIPHW, CoMem, and CoMem-DIPHW by considering the incidence matrix $\mathbf{I}_{\mathcal{H}}$ of the hypergraph $\mathcal{H}$ as a bipartite graph.
-The two sets of nodes in the bipartite representation are the set of cells V and the set of genes E.
-Similarly, each gene is connected to all cells in which it is actively expressed, with weights corresponding to the expression level.
-This alternative representation allows for the use of bipartite graph algorithms and conceptualization for analyses.
-For example, we used Barber’s bipartite modularity  [42] to assess the how the clustering performance of different algorithms vary with modularity of the underlying scRNA-seq dataset.
-.3 Evaluating the Impact of Correlation Methods on Network-Based Clustering Performance
+S2 Evaluating the Impact of Correlation Methods on Network-Based Clustering Performance
 There are two primary strategies for handling shared zeros in cell expression profiles of scRNA-seq data during correlation computation.
 The first strategy, exemplified by cosine similarity, is to ignore the shared zeros.
-This means that inactive expression (zeros) does not contribute to the similarity between cells.
-However, our experiments indicate that this strategy results in significantly worse clustering performance.
-Understandably, the absence of expression in the same genes between cell pairs provides information about cellular similarity and should not be disregarded entirely.
-In Table S2, we evaluated how co-expression networks generated by different correlation computation methods lead to varying performance across various network-based clustering methods.
-The correlation methods considered were Pearson’s, Spearman’s, Kendall’s Tau, and cosine similarity.
-Our results show that cosine similarity often led to the worst clustering results among the correlation computation methods tested.
+In this approach, inactive expression (zeros) does not contribute to the similarity between cells.
+However, our experiments indicate that this strategy leads to significantly worse clustering performance.
+The absence of expression in the same genes between cell pairs provides information about cellular similarity and should not be disregarded.
+We evaluate how co-expression networks generated by different correlation computation methods affect the performance of various network-based clustering methods.
+See Table S1.
+Our results show that cosine similarity often produces the worst clustering results among Pearson, Spearman, and Kendall Tau.
 Method ARI NMI
@@
 Infomap 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
-Leiden -0.001 -0.001 -$0$0.000 0.062 0.040 0.082 0.073
+Leiden -0.001 -0.001 -0.000 -0.000 0.062 0.040 0.082 0.073
 Multilevel 0.032 0.034 0.034 0.035 0.305 0.289 0.325 0.319
@@
 Table S1: Impact of Correlation Methods on Network-Based Clustering Performance.
-The table evaluates the impact of correlation computation methods—Pearson’s, Spearman’s, Kendall’s Tau, and cosine similarity—on the performance of various network-based clustering methods.
-The correlation computation method resulting in the worst clustering performance is marked in bold.
-Cosine similarity, which ignores shared zeros in cell expression profiles, often resulted in the worst clustering performance.
-These results show that, although using correlation computation methods that treat zeros the same as non-zero signals could inflate correlations, ignoring these zeros, as in cosine similarity, can lead to worse performance.
-The primary implication of this experiment is that both strategies have inherent drawbacks when dealing with the sparsity of scRNA-seq data.
-Ignoring zeros misses out on significant information, while treating them as homogeneous signals inflates correlations and introduces noise.
-Ignoring the shared zeros in cell expression profiles, as cosine similarity does, is not an effective strategy and results in the worst performance.
-Conversely, methods that consider all the shared zeros in the expression profiles as an indication of cell homogeneity —such as Pearson’s, Spearman’s, and Kendall’s Tau—neglect the fact that many shared zeros are merely a result of the high dimensionality of scRNA-seq data.
-When sequencing results from multiple cells are merged, many zero entries are introduced since different cells often express different sets of genes.
-Consequently, cell correlations are inflated by these methods when data is sparse.
-The clustering outcomes from these methods are suboptimal when compared with approaches that do not require correlation computation, such as CoMem, DIPHW, and PCA, potentially due to misleading signals in the inflated correlation networks S1 4 5 6.
-Our results suggest that alternative approaches, such as hypergraph-based methods, provide a better solution by circumventing the need for unipartite network projections and avoiding the intricacy associated with handling zeros in correlation computations.
-.4 Embedding Visualizations
-We present the visualizations and clustering performance of embeddings generated by DIPHW, CoMem, and CoMem-DIPHW in Fig.
-S1.
-We compare these against five other embedding-based methods (PCA, Node2Vec, EDVW [41], UMAP, and t-SNE) and four community detection methods (Louvain, Greedy Modularity, Infomap, and Leiden) applied to simulated single-cell RNA sequencing (scRNA-seq) data.
-Embedding-based methods compute representations of the cells in a lower-dimensional space while preserving their global geometry [77, 78, 79].
-For these methods, we used UMAP for embedding visualization and K-means for clustering.
-In contrast, community detection methods identify clusters (communities) based on the structures of the unipartite networks projected from the scRNA-seq data.
-Although t-SNE and UMAP are typically not used as a standalone embedding algorithms, we have included them in the comparison for reference, as they are adopted by popular packages in single-cell analysis, such as Seurat [15, 16], Scanpy [17], SC3 [50], and Cell Ranger [26].
-There has been ongoing discussions about using them as the main clustering methods in scRNA-seq analysis [80, 81].
-We note that the strength of t-SNE and UMAP lies in preserving local structures, such as the nearest neighbors [80, 54], while UMAP can also preserve global distance [54].
-However, they are not designed to group data into clusters with high internal cohesion.
-Therefore, they are generally used in conjunction with other clustering methods as visualizing tools that map computed embeddings to a 2D space for evaluation.
-The results are compared with the ground-truth cell-type membership in the embedding visualization and in the NMI and ARI bar plots.
-For visualization, UMAP was applied to the embeddings produced by the dimensionality reduction algorithms.
-Each dot represents a cell, with its color indicating the ground-truth cell cluster membership and its position representing the embedding.
-As illustrated, DIPHW and CoMem-DIPHW demonstrated the highest degree of internal cohesion within clusters and the best cluster separation, as evidenced by both the embedding visualizations and the ARI and NMI scores.
-We also evaluate CoMem as a baseline method, which shares the same memory mechanism as CoMem-DIPHW but without the dual importance (edge-dependent vertex selection probability and vertex-dependent edge selection probability) or preference exponent.
-This comparison shows the improvements brought by integrating the hypergraph-based random walk DIPHW with the memory mechanism in CoMem.
-The parameters used to simulate the scRNA-seq data in this experiment were chosen to capture the main characteristics and underlying assumptions of scRNA-seq analysis.
-Although these parameters may not perfectly match the values observed in real scRNA-seq data, they provide a challenging enough clustering task to differentiate the performance of the evaluated clustering algorithms.
-The inter-module density is set at 0.1, and inter-module connection probability at 0.6 to represent cell-type cross-talk [63, 66, 65, 64, 82], resulting in a weaker expression signal density than within cell-type expression.
-To distinguish cell-type-specific signals from background expression, we set the baseline gene expression level at 10 and the within-module average expression at 20.
-Figure S1: Clustering Results on Simulated ScRNA-seq Data Using Embedding-based and Community Detection Methods.
-(a) UMAP visualization of the embedding-based clustering results, with colors indicating ground-truth cell cluster membership.
-(b) NMI and ARI bar plots comparing the performance of the embedding and community detection methods.
-Higher NMI and ARI values suggest greater consistency with ground-truth clusters, indicating better clustering performance.
-DIPHW and CoMem-DIPHW demonstrated the highest degree of internal cohesion within clusters and the best cluster separation, as evidenced by both the embedding visualizations and the ARI and NMI scores.
-For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
-Additional Results for Clustering Performance Across Modularity
+The table shows how different correlation measures (specifically, Pearson, Spearman, Kendall Tau, and cosine similarity) affect the performance of various network-based clustering methods.
+The correlation computation method that results in the worst clustering performance is marked in bold.
+Cosine similarity, which ignores shared zeros in cell expression profiles, often produces the worst clustering performance.
+These results indicate that while correlation computation methods that treat zeros the same as non-zero signals may inflate correlations, ignoring these zeros, as in cosine similarity, can lead to poorer performance.
+The main implication of this experiment is that both strategies have limitations when handling sparse scRNA-seq data.
+Cosine similarity’s omission of shared zeros results in the poorest performance.
+In contrast, methods that treat all shared zeros in expression profiles as indicators of cell homogeneity (such as Pearson, Spearman, and Kendall Tau) overlook the fact that many shared zeros arise simply from the high dimensionality of scRNA-seq data.
+When sequencing results from multiple cells are combined, many zero entries appear because different cells often express different sets of genes.
+Our results indicate that alternative approaches, such as hypergraph-based methods, offer a better solution by avoiding the need for unipartite network projections and the complexity of handling zeros in correlation calculations.
+S3 Computational and Memory Complexities of DIPHW and CoMem-DIPHW
+DIPHW.
+The computational cost of each element $P(u\rightarrow v)$ in the $|V|\times|V|$ node-to-node transition probability matrix is $O(|E|)$ when the intermediate matrices for the node-to-edge and edge-to-node transition probabilities $P_{E|V}(e|u)$ and $P_{V|E}(v|e)$ are precomputed.
+The construction of the entire $|V|\times|V|$ node-to-node transition probability matrix involves computing these probabilities for all node pairs, resulting in an overall time complexity of $O(|V|^{2}\cdot|E|)$.
+For efficient computation, the $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ in martix form is given by:
+$P_{V|V}=D_{E|V}^{-1}W_{E|V}D_{V|E}^{-1}W_{V|E}$
+The theoretical complexity remains $O(|E|^{2}\cdot|V|+|E|\cdot|V|^{2})$.
+However, the matrix representation utilizes parallel computation and is significantly faster, making it more suitable for large datasets.
+DIPHW’s memory complexity is $O(|E|\cdot|V|)$ for the node-to-edge and edge-to-node transition probability matrices.
+CoMem-DIPHW.
+The computational complexity of CoMem-DIPHW is $O(|E|^{2}\cdot|V|^{2})$.
+For each pair of nodes, the memory-incorporated transition traverses all possible edges that could have reached the first node and all possible edges that could then connect the first node with the second node.
+CoMem-DIPHW’s memory complexity is $O(|E|^{2}\cdot|V|+|E|\cdot|V|^{2})$ for the memory-incorporated node-to-edge and edge-to-node transition probability matrices.
+S4 Bipartite representation
+We choose the hypergraph conceptualization over the bipartite representation for consistency, as the DIPHW algorithm we introduce is an extension of the edge-dependent vertex weight hypergraph random walk (EDVW) [Chitra2019].
+DIPHW extends the EDVW hypergraph random walk by incorporating a vertex-dependent edge selection probability and a preference exponent to accelerate clustering.
+A bipartite conceptualization could also be applied to our proposed methods by interpreting the incidence matrix $\mathbf{I}_{\mathcal{H}}$ of the hypergraph $\mathcal{H}$ as a bipartite graph.
+In this representation, the two sets of nodes are the set of cells $V$ and the set of genes $E$.
+Each gene is connected to all cells in which it is actively expressed, with weights corresponding to the expression level.
+This alternative representation enables the use of bipartite graph algorithms.
+For example, we use Barber’s bipartite modularity [barber2007modularity] to assess how the clustering performance of different algorithms varies with the modularity of the underlying scRNA-seq dataset.
+S5 Quantitative Evaluation of Clustering Quality on Tissue Datasets
+The quantitative evaluations in Tables S2 and S3 offer complementary perspectives on clustering performance for tissue scRNA-seq datasets.
+The evaluation in the learned embedding space (Table S2) reflects the geometric separability and compactness of the clusters produced by each method, while the evaluation of the preprocessed expression data (Table S3) assesses whether the same cluster assignments capture the structure of the preprocessed gene expression data used as input to the embedding methods.
+In the embedding space, graph-sc and scASDC achieve strong geometric separation and cohesion across datasets.
+However, they show reduced clustering performance when evaluated on the preprocessed data.
+This aligns with the results in Fig.
+12, where scASDC produces embeddings with strong cluster separation and cohesion, but some identified DEGs correspond to biologically implausible cell types.
+In particular, PCA performs best in the preprocessed data evaluation.
+This is consistent with the findings of Ciortan et al.
+[ciortan2022gnn], where PCA outperformed many state-of-the-art methods, including graph-sc, when evaluated by the Silhouette score and the Calinski-Harabasz index on real and simulated datasets.
+Human Brain Human Pancreas Mouse Brain Mouse Pancreas
+Method SC CH DB CV SC CH DB CV SC CH DB CV SC CH DB CV
+scASDC 0.421 244.710 0.975 2.373 0.763 7005.583 0.549 0.821 0.816 5230.004 0.640 1.845 0.874 2843.512 0.462 0.563
+graph-sc 0.352 705.099 0.844 0.560 0.529 4164.981 0.584 0.816 0.418 2431.617 0.982 0.905 0.394 976.784 0.768 0.495
+CoMem_DIPHW 0.285 224.443 1.207 1.287 0.384 1048.921 1.119 1.058 0.396 1222.851 0.936 1.403 0.274 313.957 1.183 1.677
+CoMem 0.265 199.820 1.285 1.922 0.474 1447.208 0.852 3.575 0.455 1432.727 0.817 1.086 0.291 287.810 1.179 4.768
+tsImpute 0.263 703.410 1.163 8.200 0.457 715.814 0.789 5.051 0.409 867.564 1.095 8.499 0.424 666.677 0.902 10.673
+PCA 0.258 236.768 1.412 6.802 0.294 542.832 1.270 89.515 0.229 512.127 1.652 83.396 0.107 120.386 1.804 125.705
+CAKE 0.156 164.736 1.679 2.849 0.417 3999.076 0.794 1.944 0.461 12194.111 0.800 2.684 0.237 373.750 1.391 1.825
+node2vec 0.133 172.629 2.118 2.307 0.409 1405.125 1.034 1.359 0.332 1108.268 1.188 27.464 0.198 265.955 1.467 1.014
+DIPHW 0.122 110.724 2.138 0.903 0.204 480.360 1.328 1.060 0.127 264.797 2.444 5.145 0.055 35.369 3.380 3.147
+EDVW 0.109 97.634 2.315 1.720 0.152 402.944 2.039 6.156 0.126 246.265 2.616 5.840 0.060 40.552 3.482 5.804
+Table S2: Clustering Quality Evaluated in the Learned Embedding Space on Tissue scRNA-seq Datasets.
+Clustering performance is assessed using the Silhouette score (SC), Calinski–Harabasz index (CH), Davies–Bouldin index (DB), and Coefficient of Variation (CV) on the learned embeddings produced by each method.
+Higher values indicate better performance for the Silhouette score and Calinski–Harabasz index, while lower values indicate better performance for the Davies–Bouldin index and Coefficient of Variation.
+This evaluation reflects the geometric separability and cohesion of clusters in the embedding space.
+The best values within each dataset/evaluation measure column are highlighted in bold, and the second-best values are underlined.
+Human Brain Human Pancreas Mouse Brain Mouse Pancreas
+Method SC CH DB CV SC CH DB CV SC CH DB CV SC CH DB CV
+CoMem 0.027 17.529 4.654 3.951 0.071 230.969 2.048 4.559 0.041 173.331 2.838 7.603 -0.002 37.954 4.182 5.064
+CoMem_DIPHW 0.027 17.529 4.575 3.951 0.021 235.958 2.364 4.834 0.018 169.640 3.363 7.795 0.002 40.049 3.891 5.008
+PCA 0.025 19.080 4.991 4.012 0.276 395.673 1.461 4.012 0.087 183.428 2.829 6.995 0.016 46.001 3.051 4.699
+tsImpute 0.009 12.901 7.122 3.958 0.114 171.384 2.715 3.734 -0.005 67.857 6.082 6.721 -0.025 22.239 6.907 5.127
+graph-sc 0.008 10.692 6.286 4.043 0.042 201.025 2.218 4.137 0.047 127.852 5.446 6.905 0.052 28.514 3.696 4.749
+node2vec 0.007 15.731 8.386 4.010 0.060 259.873 2.019 4.609 0.040 174.822 3.240 7.726 -0.001 37.939 4.459 5.027
+DIPHW 0.006 14.937 7.794 4.018 0.015 195.839 3.152 4.328 0.012 153.568 4.081 7.782 -0.006 36.516 5.353 5.115
+EDVW 0.002 13.243 8.843 4.042 -0.075 177.732 7.734 4.956 0.034 172.331 3.403 7.805 -0.001 39.774 4.790 4.991
+scASDC -0.009 7.670 7.899 3.947 -0.128 139.358 3.605 4.655 0.070 166.681 3.032 7.177 -0.027 26.217 6.188 5.148
+CAKE -0.012 7.713 8.723 4.002 -0.025 184.071 3.063 4.444 0.068 75.122 6.077 5.624 0.006 33.657 6.005 5.042
+Table S3: Clustering Quality Evaluated on Preprocessed Expression Data for Tissue scRNA-seq Datasets.
+Using the same cluster assignments as in Supplementary Table S2, we evaluate cluster quality on the preprocessed gene expression data used as input to each embedding method.
+We measure cluster quality using the Silhouette score (SC), Calinski–Harabasz index (CH), Davies–Bouldin index (DB), and Coefficient of Variation (CV).
+Higher values indicate better performance for SC and CH, while lower values indicate better performance for DB and CV.
+This evaluation assesses whether the cluster assignments by each method capture meaningful structure in the input gene expression data.
+The best values within each dataset/evaluation measure column are highlighted in bold, and the second-best values are underlined.
+S6 Additional Results for Clustering Performance Across Modularity Levels
+Figures S1 and S2 show additional results on the impact of module size and module count on ARI clustering performance.
+Figures S3 and S4 show the impact of module size and module count on NMI clustering performance.
+Figures S5 and S6 show the impact of module size and module count on ACC clustering performance.
+Figures S7 and S8 show the impact of module size and module count on AMI clustering performance.
+Figures S9 and S10 show the impact of module size and module count on F1 clustering performance.
+Finally, Fig.
+S11 shows a visualization of modularity in our imulated scRNA-seq Data.
 Figure S1: Clustering Performance Comparison by ARI Across Varying Module Sizes.
@@
 Red dashed lines indicate the highest ARI values or best ARI rankings.
-For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
+K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
 Figure S2: Clustering Performance Comparison by ARI Across Varying Module Counts.
@@
 Red dashed lines indicate the highest ARI values or best ARI rankings.
-For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
+K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
 Figure S3: Clustering Performance Comparison by NMI Across Varying Module Sizes.
@@
 Red dashed lines indicate the highest NMI values or best NMI rankings.
-For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
+K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
 Figure S4: Clustering Performance Comparison by NMI Across Varying Module Counts.
@@
 Red dashed lines indicate the highest NMI values or best NMI rankings.
-For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
+K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
+Figure S5: Clustering Performance Comparison by ACC Across Varying Module Sizes.
+Simulated scRNA-seq data were used for this evaluation.
+The ACC results support the same conclusion: when modularity is weak (i.e., when the average number of co-expressed genes per module is small), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced.
+Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval.
+Red dashed lines indicate the highest ACC values or best ACC rankings.
+K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
+Figure S6: Clustering Performance Comparison by ACC Across Varying Module Counts.
+Simulated scRNA-seq data were used for this evaluation.
+The results by ACC support the same conclusion: when modularity is weak (i.e., when the number of modules is greater), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced.
+Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval.
+Red dashed lines indicate the highest ACC values or best ACC rankings.
+K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
+Figure S7: Clustering Performance Comparison by AMI Across Varying Module Sizes.
+Simulated scRNA-seq data were used for this evaluation.
+The AMI results support the same conclusion: when modularity is weak (i.e., when the average number of co-expressed genes per module is small), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced.
+Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval.
+Red dashed lines indicate the highest AMI values or best AMI rankings.
+K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
+Figure S8: Clustering Performance Comparison by AMI Across Varying Module Counts.
+Simulated scRNA-seq data were used for this evaluation.
+The results by AMI support the same conclusion: when modularity is weak (i.e., when the number of modules is greater), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced.
+Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval.
+Red dashed lines indicate the highest AMI values or best AMI rankings.
+K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
+Figure S9: Clustering Performance Comparison by F1 Across Varying Module Sizes.
+Simulated scRNA-seq data were used for this evaluation.
+The F1 results support the same conclusion: when modularity is weak (i.e., when the average number of co-expressed genes per module is small), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced.
+Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval.
+Red dashed lines indicate the highest F1 values or best F1 rankings.
+K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
+Figure S10: Clustering Performance Comparison by F1 Across Varying Module Counts.
+Simulated scRNA-seq data were used for this evaluation.
+The results by F1 support the same conclusion: when modularity is weak (i.e., when the number of modules is greater), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced.
+Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval.
+Red dashed lines indicate the highest F1 values or best F1 rankings.
+K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
 Figure S11: Visualization of Modularity in Simulated scRNA-seq Data.
@@
 The number of modules models the number of cell types in the simulated data.
-Additional Results for Human Brain Dataset
+S7 Results for the Human Brain Dataset
+Figures S12S13S14S15S16, and S17 show the results on the human brain dataset for CoMem-DIPHW, PCA, graph-sc, tsImpute, CAKE, and scASDC, respectively.
 Figure S12: Clustering Performance of CoMem-DIPHW on the Human Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers.
@@
 Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
-Additional Results for Mouse Pancreas and Mouse Brain Datasets
+Figure S14: Clustering Performance of graph-sc on the Human Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers.
+(a) Top 10 DEGs identified for each cluster by graph-sc.
+(b) Across-cluster average expression of cluster-specific DEGs.
+Violin plots show the distribution of average expression levels of these DEGs across all clusters.
+Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
+(c) Cell type annotation using the PanglaoDB marker database.
+Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
+Two of the graph-sc clusters had less than 5 cells, one was excluded from the expression plots in (b), and both failed to map to known cell types in (c).
+Figure S15: Clustering Performance of tsImpute on the Human Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers.
+(a) Top 10 DEGs identified for each cluster by tsImpute.
+(b) Across-cluster average expression of cluster-specific DEGs.
+Violin plots show the distribution of average expression levels of these DEGs across all clusters.
+Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
+(c) Cell type annotation using the PanglaoDB marker database.
+Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
+Figure S16: Clustering Performance of CAKE on the Human Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers.
+(a) Top 10 DEGs identified for each cluster by CAKE.
+(b) Across-cluster average expression of cluster-specific DEGs.
+Violin plots show the distribution of average expression levels of these DEGs across all clusters.
+Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
+(c) Cell type annotation using the PanglaoDB marker database.
+Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
+Figure S17: Clustering Performance of scASDC on the Human Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers.
+(a) Top 10 DEGs identified for each cluster by scASDC.
+(b) Across-cluster average expression of cluster-specific DEGs.
+Violin plots show the distribution of average expression levels of these DEGs across all clusters.
+Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
+(c) Cell type annotation using the PanglaoDB marker database.
+Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
+S8 Results for the Mouse Pancreas and the Mouse Brain Datasets
+Figures S18S19S20S21S22, and S23 show the results on the mouse pancreas dataset for CoMem-DIPHW, PCA, graph-sc, tsImpute, CAKE, and scASDC, respectively.
+Figures S24S25S26S27S28, and S29 show the results on the mouse brain dataset for CoMem-DIPHW, PCA, graph-sc, tsImpute, CAKE, and scASDC, respectively.
 Figure S18: Clustering Performance of CoMem-DIPHW on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and Canonical Markers.
@@
 (a) Top 10 DEGs identified for each cluster by PCA.
+(b) Across-cluster average expression of cluster-specific DEGs.
+Violin plots show the distribution of average expression levels of these DEGs across all clusters.
+Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
+(c) Cell type annotation using the PanglaoDB marker database.
+Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
+Figure S20: Clustering Performance of graph-sc on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and Canonical Markers.
+(a) Top 10 DEGs identified for each cluster by graph-sc.
+(b) Across-cluster average expression of cluster-specific DEGs.
+Violin plots show the distribution of average expression levels of these DEGs across all clusters.
+Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
+(c) Cell type annotation using the PanglaoDB marker database.
+Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
+Figure S21: Clustering Performance of tsImpute on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and Canonical Markers.
+(a) Top 10 DEGs identified for each cluster by tsImpute.
+(b) Across-cluster average expression of cluster-specific DEGs.
+Violin plots show the distribution of average expression levels of these DEGs across all clusters.
+Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
+(c) Cell type annotation using the PanglaoDB marker database.
+Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
+Figure S22: Clustering Performance of CAKE on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and Canonical Markers.
+(a) Top 10 DEGs identified for each cluster by CAKE.
+(b) Across-cluster average expression of cluster-specific DEGs.
+Violin plots show the distribution of average expression levels of these DEGs across all clusters.
+Strong clustering performance is indicated by high expression of cluster-specific DEGs within their respective clusters and low expression in other clusters.
+(c) Cell type annotation using the PanglaoDB marker database.
+Cell types are determined by the overlap between each cluster’s DEGs and cell type-specific markers in the PanglaoDB database, with match scores computed based on the proportion of matched markers.
+Figure S23: Clustering Performance of scASDC on the Mouse Pancreas Dataset and Cell Type Annotation using DEGs and Canonical Markers.
+(a) Top 10 DEGs identified for each cluster by scASDC.
 (b) Across-cluster average expression of cluster-specific DEGs.
@@
 (c) Cell type annotation using the PanglaoDB marker database.
+Figure S26: Clustering Performance of graph-sc on the Mouse Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers.
+(a) Top 10 DEGs identified for each cluster by graph-sc.
+(b) Across-cluster average expression of cluster-specific DEGs.
+Violin plots show the distribution of average expression levels of these DEGs across all clusters.
+(c) Cell type annotation using the PanglaoDB marker database.
+Figure S27: Clustering Performance of tsImpute on the Mouse Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers.
+(a) Top 10 DEGs identified for each cluster by tsImpute.
+(b) Across-cluster average expression of cluster-specific DEGs.
+Violin plots show the distribution of average expression levels of these DEGs across all clusters.
+(c) Cell type annotation using the PanglaoDB marker database.
+Figure S28: Clustering Performance of CAKE on the Mouse Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers.
+(a) Top 10 DEGs identified for each cluster by CAKE.
+(b) Across-cluster average expression of cluster-specific DEGs.
+Violin plots show the distribution of average expression levels of these DEGs across all clusters.
+(c) Cell type annotation using the PanglaoDB marker database.
+Figure S29: Clustering Performance of scASDC on the Mouse Brain Dataset and Cell Type Annotation using DEGs and Canonical Markers.
+(a) Top 10 DEGs identified for each cluster by scASDC.
+(b) Across-cluster average expression of cluster-specific DEGs.
+Violin plots show the distribution of average expression levels of these DEGs across all clusters.
+(c) Cell type annotation using the PanglaoDB marker database.
+S9 Hyperparameters
+Table S4 lists the hyperparameters used in each experiment.
+Figure/Table/Section Description and Value
+Experiments on simulated scRNA-seq data across modularity levels (Fig. 4 for average genes per cell type plots) Number of cells = 3000
+Number of genes = 1000
+Number of embedded modules = 30
+Average number of cells per module = 38
+Target density of the gene expression matrix = 0.03
+Density within modules = 0.3
+Density between modules = 0.1
+Probability of inter-module connections = 0.6
+Average background expression = 10
+Average within-module expression = 20
+Average inter-module expression = 10
+Workers = 30
+Walk length = 30
+Embedding dimensions = 30
+Preference exponent = 50
+Average number of genes per module $\in$ {10, 20, 30, 40, 70, 80, 90, 100}
+Experiments on simulated scRNA-seq data across modularity levels (Fig. 4 for number of modules plots) Number of embedded modules $\in$ {10, 20, 30, 40, 70, 80, 90, 100}
+Average number of genes per module = 50
+Average number of cells per module = (# genes / # modules) +3
+Other parameters are the same as Fig. 4 for average genes per cell type plots.
+Experiments on the impact of module sizes on ARI clustering performance (Fig. 5) Same as in Fig. 4 for average genes per cell type plots.
+Experiments on the impact of module counts on ARI clustering performance (Fig. 6) Same as in Fig. 4 for number of modules.
+Comparison of ARI and NMI across clustering methods for the ScMixology benchmark datasets (Table 4) Number of highly expressed genes used = 500
+Preference exponent = 10
+Workers = 50
+Walk length = 10
+Embedding dimensions = 10
+Experiments on human pancreas (Section III.6) Number of highly expressed genes used = 1000
+Number of clusters = 9
+Preference exponent = 10
+Workers = 100
+Walk length = 50
+Embedding dimensions = 10
+Experiments on mouse pancreas (Section S8) Number of clusters = 13
+Other parameters are the same as for human pancreas analysis (Section III.6).
+Experiments on human brain (Section S7) Number of clusters = 7
+Other parameters are the same as for human pancreas analysis (Section III.6).
+Experiments on mouse brain (Section S8) Number of clusters = 9
+Other parameters are the same as for human pancreas analysis (Section III.6).
+Table S4: Hyperparameters Used in Each Experiment.